A Complete Bibliography of Publications in Marine Policy (1980–1989)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

14 October 2023 Version 1.02

Title word cross-reference

$100 [Bro83a, Van87, Vuk85]. $107.00 [Mar89c]. $11.75 [Ran82]. $14.95 [Roe81]. $15.00 [Lia82]. $17.50 [Smi82]. $18.50 [Mun86]. $189.00 [Gla89c]. $22 [Wat82a, Mor81a]. $22.00 [Sad81]. $27.50 [Pea81b]. $27.95 [BC83b]. $28.50 [Ale89]. $29.50 [Kim86a]. $3.90 [Wat81a]. $30.00 [Bro80, Mun88]. $34.00 [Roe81]. $38.50 [Gla89h]. $40.00 [Low81a]. $45.00 [For80b, Pla86]. $49 [Wat85a]. $49.00 [Wat86b]. $5.75 [Ran82]. $50.00 [Hol88b]. $6.50 [Nai82]. $60.00 [Hat88]. $61.00 [Gil86]. $64.00 [Mar89a]. $72.00 [Mar89a]. $82.00 [Mar89b]. $90.00 [Mar89b]. $99.00 [Mar89c]. i [Kwi84]. x [Wat87f].

/International [Kim86b].

00.00 [Kim88a].

1 [Van87]. 10.00 [Wat80a]. 10.85 [Arc81a]. 10.95 [Gul88a]. 112.25 [Kwi87]. 12.50 [Law86, Pon86]. 120 [McR86]. 124pp [Wat81a]. 14.95 [hP81].

142-236 [Ano83g]. 14th [Edm81]. 15.00 [Alf80, Pri80]. 15.95 [Loc84]. 152pp [Abr84a]. 158-268 [Ano80f]. 15th [Mar82b]. 16.25 [Ran81]. 16.50 [Bro85]. 16.95 [Wat87i]. 162-255 [Ano82f]. 163-256 [Ano87l]. 16th [CF81, Din83]. 17.00 [Spa87]. 17.50 [Len87, Pap84]. 170-255 [Ano85g]. 171-240 [Ano86l]. 171-292 [Ano81h]. 178-272 [Ano89h]. 179-323 [Ano88h]. 17th [Wat84a]. 18 [App85]. 18.50 [Man81]. 18.75 [Cop85]. 19.50 [Wat81f, Wat85b]. 19.85 [Law85]. 19.95 [Wat87e]. 1978 [Gon80]. 1979 [Bir80b, Bir80d, Sha80b, Sha82a]. 1980 [Arc81b, CF81, Edm81, Gre80, Kim80a, Lon81, Low81b, Man80, Pow80, Pul80a, Sha82b, Pon86]. 1980s [CF81, Lon81, Aub86]. 1981 [All82, Bir81c, Bir82a, CS82, Edm82, For81a, Geo82, Kim82b, McE82a, Mer83, Pau82, Sco82]. 1982 [Ano84n, Bir83a, De 83, Din83, Gon83, Kim82a, Mar82b, McE82b, Ode84, Pul82, Sim83]. 1983 [Gul83, Kil83, McB84, Met84, Mit83, Pul84, Wat84a, Kwi85, Stu86]. 1984 [Bea85, Kim84, Kim85a, Lei85, Mad85b, Stu85b]. 1985 [Bir86a, Bir86b, Fab85, Gla89h, Kim85b, Kim86b, PM85, PT86, You86, Gla89c, Kim88a]. 1986 [Bir87c, Hum87b, Kim87, Mat86, Phi87, Wec87, tH87, Gla89g]. 1987 [Ale88, Ano87f, Ano88g, Ano88o, Hum87c, Kim88b, Kwi88b, Kwi88a, Mor89, Ram88]. 1988 [Bir89, BC89, Doe89, Gul89a, Gul89b, Kim89, SK88, Wec89, Yat88]. 1989 [Ano89a, Ree89b]. 199pp [McR86]. 1st [Kwi88a].

2 [Wat82b]. 2.95 [Law84]. 20 [Spe83]. 200-mile [Cop80]. 203-296 [Ano84j]. 21.50 [Ash82]. 21st [Gul87b]. 22 [Doe89]. 230pp [Heg81]. 238-324 [Ano83h]. 24.40 [Gul87d]. 243-324 [Ano86m]. 25 [Bir80b, Law83]. 25.00 [Gol87b, Mor88b, Wat87b, Wat87f]. 256pp [Wat87i]. 258-351 [Ano85i]. 258-352 [Ano82h]. 259-334 [Ano87m]. 268pp [Wat80a]. 27.05 [AC82]. 27.50 [Bea88]. 270-340 [Ano80h]. 274-368 [Ano89j]. 27th [Bar89]. 288.50 [Mar88c]. 290pp [Alf80]. 294-353 [Ano81j]. 299-376 [Ano84k].

30 [Ano88o, Bir87b]. 30.00 [Wal89, Wat87a]. 309pp [Pri80]. 31.95 [You87a]. 32.00 [Len80]. 32.50 [Jls89b]. 326-420 [Ano88j]. 33rd [Bir82a]. 34th [Bir83a]. 35 [Wat85c]. 35.00 [And89a, Wal86, Stu88]. 354pp [Gul86]. 36.50/US [Mar89a]. 38.00/US [Gil86]. 39.50 [Ros86]. 394pp [Wat81e].

4 [Gin86, Wat85a]. 40 [Gul86, Low85]. 40th [Bir89]. 41.00/US [Mar89a]. 45.50 [Kwi89b]. 46.50/US [Mar89b]. 47.50 [Hoy85]. 48.50 [Bla86].

5 [Wat86b]. 5.50 [Wat81c]. 51.00/US [Mar89b]. 55.00 [Geo89, Wat87d]. 56.50/US [Mar89c]. 574 [Moo86]. 576pp [Kim86a].

6 [Ree89b]. 61.00/US [Mar89c]. 685pp [Gla89c].

7.50 [Wat81b]. 7.95 [Jls89a]. 713pp [Wat82b]. 73.50 [Mar88b]. 74-140 [Ano83j]. 75.00 [Gol87a]. 7AF [Law84].

8.50 [Bir85]. 8.75 [Fri80]. 80 [Len80]. 82-176 [Ano88k]. ’86 [tH87]. 87-176 [Ano89k]. 8PH [For85a].

9.50 [Nai81, Wat81e]. 90-156 [Ano80i]. 90-158 [Ano87o]. 90-160 [Ano82i]. 90-166 [Ano86o]. 90-168 [Ano85j]. 98-168 [Ano81k]. 9JN [Wat82c].

aan [Bro80]. Abecassis [Gol87a]. Abrahamson [Pea81b]. Abt [Til85]. academia [Sim83]. Academic [Bar85]. accès [Moo86]. Access [Tei87, GGM86, Moo86]. Accord [Low80]. accuracy [McM85]. ACMRR [Bir80c]. ACOPS [Seb88]. Act [Ode84, Nag80, Fin85, Tur85, CS85]. active [PK85]. activities [Bir83b]. adaptation [Arc83]. Adaptive [HS88a]. Adede [Kwi89b]. Ademuni [Gil85]. Ademuni-Odeke [Gil85]. adjustment [Böh89, Pon88]. Administering [Stu84]. Administrative [Wal82]. Admiral [Wat87e]. adoption [Tre83]. Advanced [PT86]. Advances [Wat87a]. advent [Wat87h]. Affairs [Ano88g, Hum87a, Kwi88b, Low84, Mor88b, Nai80a, Nan89, Wat81a, Gla89c, Moo81, Wat87d]. affecting [McM85]. Africa [Moo87, LR83, RL86, WGG82]. African [Kac89]. After [TL84, Bir81a, Lav87, SK89, Jin89]. Aftermath [Nag80]. against [Dre88, Low82c, Sal85, Wat84b]. age [Gul84a]. AGel [Fab85]. agendas [Mer83]. agreement [Bil86, Chu83, Coh82]. agreements [Bla83, Tei87, Wat81a]. Agriculture [Gla89e, Gon80, Loc84, Mad85b]. ahead [Cou89, For81a, Gla89g, Nai80b]. aid [App83a, RL86]. airmail [Mar89a, Mar89b, Mar89c]. Airy [Gla86]. Alaska [BL89]. Alastair [Ale85]. Aldershot [Abr84a, Abr84b, Bro85, Wat85b]. Alexander [Aub86, Cou89, Gon80, Hat88]. Alford [Nai81]. Alfred [Hum83b]. Ali [Wat81f]. Allanheld [Nai81]. Allen [Wat81c, Wat87c]. Alliance [Ran81]. allocation [Bin89, Fri85, Wes82]. Allsopp [Law86]. Alphen [Bro80]. Alternative [Cru85]. American [Abr83, Gla89a, Lun80, Mor81b]. Amoco [KS85, Nag80]. among [DT86, TD86]. Amos [Stu85a]. Amsterdam [Mor88a, Sha82b]. Analysis [BG87, Nai82, Bim86, CL82, Cop82b, Hee88, Kwi87, MT82, RBA87, Sto80b, Smi82]. anchor [MSV85]. anchoveta [Gul80b]. Anderson [Gul87d]. Andronico [Kwi89b]. Angeles [Wat81e]. Anglo [Pla86, But88b, Ree89b]. Anglo-Canadian [Ree89b]. Ann [For85a, Law83, Law83]. Annapolis [Wat84c]. Annual [Bir86a, Mar82b, Wat84a, Wec87, CF81, Din83, Edm81, Gla89h, Gla89g, Gul89a, Ram88, Wec89]. Antarctic [Gla89d, Hum83a, Ber85, McE84, Aub86]. Antarctica [Ano84f, SK89, Wes82, Gla89b, Gla89a, Sut81a]. Antilles [Kwi89a]. Antonio [Moo86]. Antony [Wat87i]. appendices [Ros86]. Appendix [Ano86a, Sut81a]. appendixes [Fri80]. application [PM85]. applications [Mun88, Cop85]. Applied [Bol89]. appraisal [Cru86, Fra87]. approach [Jen85, Lan86, Ran82, You82]. approaches [Alh82, BC84, Hay84, Men84]. April [Ano80i, Ano81k, Ano82i, Ano83j, Ano85j, Ano86o, Ano87f, Ano87o, Ano88g, Ano88k, Ano88o, Ano89k, For81a, Kim80a, Kim82a, Kim84, Kim85b, Kim87, Kim88b, Kim89, McB84, McE82b, PT86, Pul82, Yat88, Ano88n]. Aquaculture [Pul80a, Bjø88, DOS87, Pul82, Pul84, PN84, Wil82, Pen84, Smi82, Smi83]. Arabian [Mar88b, Mar88c, Gra88]. Arbor [For85a, Law83]. Archer [Mor88b]. archipelagic [Ham84]. Arctic [McR86, Har87, Kil83, Lam87, Øst86, You81]. area [Chu85, Mar89b]. areas [Ano84n, Bro81a, Gil86, Mar89a, SG83, Kis88, Bir85]. Argentine [Hu87]. argument [vH81]. arising [Low86]. arms [LT82, Nai80a, Wat81a]. arrangement [vH81]. arrangements [Mer83, Øst86]. Arthur [Mun86, Van85]. articles [Ano80m, Ano80n, Ano81n, Ano81o, Ano82p, Ano82n, Ano82o, Ano83o, Ano83p, Ano83q, Ano84m]. Artificial [Wah83]. Artisanal [LR83, Law88, Dah88]. ASEAN [Mor82]. Asia [Ano88g, hP81, Mor84, Tie84, Gla89f]. Asian [Ano88g, Spa87]. Aspects [Bir80d, Pow80, Bir87c, Chu88, Gol86, KG88, Sal83, SK82, Wal82]. assembly [Ano84f]. assessment [Kil83, Loc84, For81b]. assessments [PM85]. Assisting [Lof81]. Associated [Pow80]. Association [Fab85, Geo83, Pul80a, McR86, Stu85b, Nan89]. ASW [LT82]. Athens [Geo89, Ale88]. Atlantic [PT86, Ran81, DT86, Sul89, TD86, Mun86]. Atlas [Wat87i, Ale85]. Auburn [Abr83]. Auckland [Bir89]. August [Doe89, Gul89a, Gul89b, Kim82b, Kim85a, Kim86b, Kim87, Kim88b, Kim89, Kwi85, Kwi88a, Ram88]. Australia [Hum83a, Hum87b, Hum87c, Pau82, Pow80, Ber85, Ber87, Hum87c, Sut82a, Hum87a]. Australian [Boo81, Hum87a, Pow80, Ber86a, Haw89]. Author [Ano81a]. Authority [Kim86b, Ric81b]. authorizing [Bil86]. authors [Ano81g, Ano83f, Ano84g, Ano84h, Ano85f, Ano86i, Ano86j, Ano86k, Ano87h, Ano87i, Ano87j, Ano87k]. automobiles [Böh89]. autonomous [Sul80]. available [CS85, For83a].

B [Aub84, Bir85, Hum83a, Jls89b, Kwi84, Nai80a]. Bali [Ano84n]. Baltimore [Gul87d]. Bandung [Kwi88a]. Bangkok [Ano88g]. Bank [Heg81]. Baradi [Mer83]. Barbara [Sut81a, Sut81b]. bargaining [Bir81b]. Barston [Wat81c]. base [PN84]. based [PM85, Sal85, WC88]. basic [Loc84, Smi82]. basin [Mar82b]. basis [Bir80c, BC82, Dah88, PT86]. Bateman [Hum87a]. Bates [Wal86]. battle [Sid80]. Baux [Wat82c]. be [Wat82c]. Beckenham [And89a]. bed [Gil86, Mar89a, Mar89b, Mar89c]. Beddington [Gul86]. behaviour [Hee88]. Bekiashev [Bir83b]. Belgium [Vol86]. Belhaven [BC89]. Belize [Shu84]. Bellver [Abe83]. Benefit [Bla83]. benefits [CG81]. Bennathan [Heg81]. Berkeley [Wat81e]. Berlin [Kwi89a, Low82b]. Bernhard [Pea81b]. better [Böh86]. between [BC89]. Beverton [Gul86]. beyond [Gul84a, Mar89b, SK89]. bibliography [Gla89c, You85, Stu85a, Mar89c]. bilateralism [Mar87, Stu86]. billfish [Gul89b]. Bioeconomic [GGM86]. Biological [Pul80a, FJ85, Moo86]. biologiques [Moo86]. Birnie [Hol88b, Wat81c]. bitter [Bro81a]. Blake [Bea88]. bleue [Mor88a]. blow [Abe83]. Blue [Gul80a]. bluffs [Boo81]. Board [Ano80e, Ano81e, Ano82e, Ano83e, Ano84e, Ano85e, Ano86h, Ano87e, Ano88e, Ano89g]. body [Sul80]. Boesch [Bol89]. Book [Abr83, Abr84a, Abr84b, Ale85, Ale89, Alf80, And89a, App85, Arc81a, Arc83, Ash82, Aub84, Aub86, AC82, Bar82, Bar85, Bea88, Bim86, Bir82b, Bir83b, Bir85, Bir87b, Bla86, BC83b, Bol89, Boo81, Bro80, Bro83a, Bro85, Chu84, CS85, Cop85, Cou89, For83a, For85a, For80b, Fri80, Geo89, Gil86, Gla86, Gla89d, Gla89a, Gla89h, Gla89b, Gla89f, Gla89e, Gla89g, Gla89c, Gol87b, Gol87a, Gon80, Gul85, Gul86, Gul87d, Gul88a, Hat88, Heg81, Hol88b, Hoy85, Hum83a, Hum83b, Hum87a, Jls89a, Jls89b, Kim86a, Kim88a, Kwi84, Kwi85, Kwi87, Kwi89a, Kwi89b, Law83, Law85, Law86, Len80, Len87, Lia82, Loc84, Low81a, Low82b, Low82a, Low84, Low85, Man81, Mar88b, Mar88c, Mar89a, Mar89b, Mar89c, McC80c, McR86, Mer83, Moo87, Mor81a]. Book [Mor89, Mor88a, Mor88b, Mun86, Mun88, Nai80a, Nai81, Nai82, Ogl84, Pap84, hP81, Pea81b, Pen84, Pla86, Pon86, Pri80, Ran81, Ran82, Roe81, Ros86, Sad81, Sal85, Sco89, Sha80b, Sha82a, Sha82b, Smi82, Smi83, Spa87, Spe83, Stu85a, Stu86, Stu88, Sut81a, Sut81b, Til85, Van87, Van85, Vuk85, Wal86, Wal89, War85, Wat80a, Wat81a, Wat81f, Wat81c, Wat81e, Wat81b, Wat82a, Wat82b, Wat82c, Wat84b, Wat84c, Wat85b, Wat85c, Wat85a, Wat86a, Wat86b, Wat87a, Wat87b, Wat87c, Wat87e, Wat87d, Wat87h, Wat87g, Wat87f, Wat87i, Wei84, Wil82, You85, You87a]. Books [Ale85, Ash82, Gul85, Jls89a, Law86, Man81, Til85, Wat81a, Wat87g]. Bookwatch [Bai82, Bai83a, Bai83b, Bai84]. Booth [Wat87c]. Bordeaux [Man80]. Borgese [Wat82b, Wat85a, Wat86b]. Borough [For85a]. Boston [Abr83, For85a, Kwi87, Roe81]. bottom [Böh83]. Boulder [Ale89, Cop85, Mor81a, Pea81b, Ran81, Smi82, Smi83, Wat82a, Wil82]. boundaries [Ale88, Bea88, Bir87b]. Boundary [Wat85b, And89a, Gla89f, MSV85, McM85, Bla86]. Bournemouth [Bir86b]. Bowen [Wil82]. Bowett [Bro80]. Bowker [Pap84]. Brandt [Gul85]. Brazil [Wat82a]. Brian [McC80c, Wal89]. Brighton [Bir82a, Bir83a]. Britain [Bar85, Bir82b, Edm82, For85a, Ano89a, Spe83]. British [Jls89b, Sad81, Wat86a, And88, Cab87a, Cop80, You87b]. Broder [Bar82]. Brown [Gil86, Gla89h, Mar89a, Mar89b, Mar89c]. Bruce [Lia82, Pen84, Wat84c]. Brussels [Mit83]. Buikema [Van85]. burden [Bla83]. business [Ree89a]. Butler [Pla86, Vuk85]. Butterworth [For85a].

C [Aub86, Bro85, Gla89d, Lia82, Ros86, Sha82a, Smi82, Wat80a, You85]. Cable [Spe83, Wat87b]. Cadiz [KS85, Nag80]. CAIMAR [Nan89]. Cairns [Van85]. Calendar [Ano80a, Ano80b, Ano80c, Ano80d, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86b, Ano86c, Ano86d, Ano86e, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano89e]. California [Wat81e, Wil82]. Cambridge [Arc81a, Bim86, Kil83, Nai82, Til85]. can [CC81]. Canada [Ano88g, Din83, Lon81, Cop82a, Doe89, Gil84, Hag84].

Canadian [Wat85b, Cru85, DT86, For81b, Gil88, HDG81, McD83, McD86, Ree89b, Sil85, TD86]. Canal [Sin88]. Canberra [Boo81, Hum87a, Hum87b]. capital [Sin86]. Captain [Wat84b, Wat81b]. Capture [Yah81]. Cardiff [Bir86a, Yat88, You85]. Cargo [Mar87, FAR81, Wij80]. Cargo-sharing [Mar87]. cargoes [Böh85]. Caribbean [Kwi89a]. carriage [Gol87b]. Carroz [Gla89e]. Carson [Bir82b]. case [Arc83, Bro83b, Gil88, Mee83a, Mun86, Vol86, Wat82a, Wil82]. Cases [Øst86]. Castillo [Abe83]. Casual [GS80]. catch [MT82]. catching [Gul85]. Catholic [Law84]. causes [Sad81]. CECAF [Law82]. CENSA [Stu85a]. Center [Aub86, CS85, Ram88, Stu86, Van87, Ano80o]. Central [Wat82c]. Centre [Boo81, Hum87a, You85, Ano88g, BC83b, Sad81]. century [Gul87b, Ran82]. cetaceans [Bim81]. Challenge [Ran81, Nai82]. Challenger [Yat88]. challenges [Gol89, Nai81]. Chamber [Stu85b]. change [Böh89, McC80c]. changing [App83b, Böh84]. Chattertons [LC82]. Cheap [Sla84]. Chemical [De 83]. Chennat [For85a]. Chicago [Wat82b, Wat86b, Wat85a]. Chichester [AC82, Hoy85, Loc84]. Child [Gla89a]. Chilean [RBA87]. China [Wat84c, You84, LP87]. Chinese [Nan89]. Chleo [Mer83]. choices [Aub86]. choke [Nai82]. Chopra [Gla89d]. Chouraqui [Sha80b]. Chris [Len87]. Christian [Gla89c]. Christopher [Gla89d]. Christos [Mor89]. Chulalongkorn [Ano88g]. Churchill [Gla89h, Wei84]. civilized [Sla84]. claims [Gri88, RBA87, Kwi87, Ros86]. Clarendon [Jls89b, Low85, Wal89]. Clark [Jls89b]. Clay [Bar82]. Cleopatra [Gol87b]. Clingan [Gla89g]. Clive [Mor88b]. closed [CC81]. Closing [Cab87a, You87b]. cloth [Ran82]. Clyde [Wat87g]. CNDP [Wat82c]. CO [Mor81a, Pea81b, Ran81, Smi82, Smi83, Wat82a, Wil82, Ale89, Jen89]. co-management [Jen89]. coalition [Til85]. coast [For81b, Lon81]. Coastal [CS85, Fri80, Mun85, Pul80a, Ano84n, AC82, Ber86b, CC81, Cop80, Edm82, Hoy85, Kac89, Kwi89a, Mun86, Wil82]. coastwise [Asp88]. Cochin [Pul80a]. cod [Sul89]. Code [Ode84, Stu85a, Geo88, Geo89, Moo84, Pap84, Stu84, Stu86, Wij80]. Coleman [Law84]. collaboration [GGM86, Ree89b]. College [Ree89b]. colloquium [Arc81b]. Colombo [Kwi88b]. Colorado [Cop85, Wat81a]. Columbia [Sad81, Cop80, Wat87h]. Combating [De 83]. comment [Dow84, You87b]. commentary [Van87]. comments [Spe82]. Commerce [Stu85b]. Commission [Bir82a, Bir83a, Bir86b, Gon80, Kim84, Kim85a, Kim85b, Kim86b, Kim87, Kim88b, Kim89, Phi87, Bir89]. commissioning [Ben87]. Committee [Bar89, For81a, Kwi88b, Val89]. Common [Kee83, Chu84, Fri85, Lei85, Cun80]. commons [Ber86b, Pon88]. Commonwealth [Kum82]. communautaire [Mor88a]. communications [Lun84, Mer80]. communities [Mun86]. community [Lei85, Nag80]. Company [Abr83]. Comparative [Kwi84, Ber86b, Car81, Kin80b, Mor81b]. comparison [Fri85, Tom84]. compensation [KS85]. competitive [Abr84a]. complex [Shu82]. complexities [Tsa88]. comprehensive [Hoy85]. compromise [Bir81b, CC81]. concept [Har87, Low82a]. concepts [Smi82]. Concern [tH87]. concerned [Bim86]. concerning [Fra87]. conditions [Pan80]. Conduct [Stu85a, Geo88, Geo89, Stu84, Wij80, Ode84]. Conference [Bea85, Bir86a, CS82, Fab85, Geo82, Kwi88b, Mad85b, Mar82b, Nai80a, Wat84a, Wec87, All82, Ano88g, Ano88n, Ano88o, Bro83a, CF81, Cia85, Din83, Edm81, Geo88, Gla89h, Gla89g, KS85, Kim80a, Kim82a, Mad85a, Mar82a, McB84, McE82b, Met84, Pee84, Ram88, Ree89b, Seb88, SK88, Stu85b, Wec89, Dad85, Kim82b, Kwi88b, Mit83, Geo89]. Conferences [Ode84, Stu85a, Ell83, Geo88, Wij80, Stu85a]. confiscated [Sha80b]. confisquée [Sha80b]. Conflict [Cab85, Sul89, BB83, Faw82, hP81, Wah83, You84]. conflicting [Jls89b]. conflicts [Bir80b, Sha82b]. confrontation [Kim86a]. confrontations [Sut82b]. congress [Ano84n]. Conner [CS85]. conquest [Wat81b]. Consciousness’ [Geo83]. Consensus [Kim86a]. consent [Muk81]. Consequences [McC80b]. Conservation [Bir80d, CS85, Pow80, Ano84f, Bim86, Goo80, Hol82, Hol88b]. considerations [GO89, Lon81, Mun85, hP81]. Constituting [Sul80]. constitutional [Haw89]. constraints [Cop82a]. constructive [Gil88]. consultants [Hat88]. Consultation [Gul83]. container [Abr84a, Abr84b]. containerized [Mar87]. contaminated [GO89]. contemporary [Cop85]. context [Gon83, Hat88]. continental [Bro83b, Gil84, Hol88a, McM85, RBA87, Val89]. Contribution [Kwi84]. Control [Bir82b, Eva86, HC86, Alh82, BC82, BC83a, BC84, LT82, Nai80a, Wat81a, Wat87f]. Controlling [MT82]. controls [Sad81]. Convenience [War85, Met81]. Convention [Bir86a, Kim84, Kim85a, Kim85b, Stu85a, Ano88g, Bor88, But88a, Cos87, Egi88, Gla89h, Gri88, Kim81b, Kim87, KAK88, Kwi89b, McD85, McR84, Ogl81, SG83, Tre83, Van87, McD83, Mil83, Tre81, Kim86a]. conventions [Low88, Tom84]. Conway [Pon86]. Cooperation [Kwi88b, Mar88b, Mar88c, BC89, Gil89, Gon80, Gon83, Sto80b]. cooperative [Jen85]. coordination [You80]. Copies [CS85]. Coping [Cru85]. Corazón [hP81]. corner [Kim80b]. cosponsored [Din83, Edm81, Ram88, Wat84a]. cost [Bjø88]. Costa [LS85]. costs [Sad81]. Council [Gre80, Mar88b, Mar88c, Pow80, Ran81, Sco82, Val89, God80]. countries [Bla83, Lev88, Smi83, Heg81]. country [ST86]. Couper [Ale85]. course [Cre83, MSV85]. Court [Gil84, Shi88]. cover [Wat81c]. Creating [GFGJ83]. Crescent [Wat82c]. crime [Wat84b]. crippled [Kas87]. Critical [Aub86, Car88]. Criticism [Ols82]. critique [Cun80]. Cronulla [Pau82]. Croom [And89a, Bea88, Chu84, Mor88b, Smi82, Stu88, Wat85c, Wat87f]. Croome [Wat87e]. crossroads [Pap84]. crowded [Mer83, Mer83]. Crutchfield [Mun88]. CSIRO [Pau82]. Cultivating [For85b]. Cunningham [Law85]. Current [Red89]. Custom [But88a, Car88]. cuts [Spe82]. Cuyvers [You87a]. Cynthia [Mun86]. Cyprus [Dem85].

D [And89a, Bar82, Gil86, Gla89h, Gul86, Low84, Low85, Mar89a, Mar89b, Mar89c, Roe81, Ros86, Sad81, Sha82a, Wal89, You85]. Dagenhardt [Wat87d]. Dahmani [Sco89]. Dakar [SK88]. Dale [Hat88]. Dalhousie [Din83, Moo87, Mun86]. damage [KS85, Sil85]. Dana [AC82, BC83b]. dangerous [Gol86, Gol87b]. Daniel [CS85]. data [Fri85, FJ85, Mil83]. David [Gol87a, Gul88a, Law85, Mor88b, Wat85b]. Davis [Hum83a]. Day [Gul88a, Stu85b, Wat81b]. DC [CF81, Geo82, Nai82, Ran82, Sim83]. de-commissioning [Ben87]. death [Lav87]. debris [McC80b]. decade [CF81, Kil83]. December [Bir80d, Bir80c, Gon80, Lon81, Low81b, Moo86]. Deciding [Ale88]. decimation [Tom80]. Decision [Ric81b, Fra82, Mun86, Wal89]. decision-making [Mun86]. decisions [Nai80b]. Deep [Car81, Abr84b, Bir81c, Kum82, Lan81, Low81a, Low81b, Shu82, Gre80]. deep-sea [Abr84b]. Deepsea [Arc81a]. Defence [Boo81, Jam82, Mee83a]. defense [Til85]. Defining [Tom84]. definitions [Tom84]. del [PM85]. Delegating [Jen89]. Delhi [For83a]. Delimitation [Bro81a, Øst86, And88, Chu85, Gla89f, Lan86]. demand [Rob80]. democratic [Ste84]. Denman [Low84]. Denmark [Arc83, Gul89a]. Denver [Nai80a, Wat81a]. Department [Arc83, Bar82, Hum83a, Pow80]. depths [SGG85]. Derek [Bro80]. determining [Bir80c]. developed [Smi83]. developing [Bla83, Heg81, HS88a, Lev88, ST86]. Development [Aub84, Bar82, Bla86, For81b, Kwi89b, Mad85b, Mer81a, Sut81a, Ano88g, App83a, BL89, Bur88, Fra89, Gon83, Goo80, Jen85, JM87, Kas89, Law86, Lon81, Lun80, Man81, Mar82a, Mar82b, Met80, Moo87, RS80, Sco89, SK88, Shi88, Shu84, Smi83, Sto80a, Sto80b, Vuk85, Lar80, App85, Bol89]. developments [Ano81p, Ber85, Boo80, Cop85, Moo81, Red89, And89a]. Deventer [Hum83b]. Dfl [McR86]. dimension [Wat81c]. dimensions [Roe81]. diplomacy [Wat87c, Nai80a, Wat87b]. diplomatic [KS85]. Directed [Wat82c]. directions [MB81]. director [Ols82]. disadvantaged [Ale81]. disarmament [Sut83]. Discussions [Val89]. disposal [BC86, PT86, Van88]. Dispute [Bro81b, Gil84, Wat85b]. disputes [Kwi89b]. Distant [QJ89, Dou87, Mun85, Tei87]. distant-water [Dou87, Mun85]. distress [Kas87]. Distributed [For85a, Lia82]. distribution [DT86, TD86]. Division [Pow80]. Dobbs [Bro80, Bro83a, For80b, Low81a, Pla86]. Doctorate [Moo86]. documentary [Kim88a]. documents [Kwi87, Mar89c, Bro83a]. dollars [DT86, TD86]. Donald [Bol89]. Donat [McR86]. Dordrecht [Bla86, Gla89d, Gla89c, Kim88a, Kwi87, Kwi89b, McR86, Mor89, Roe81, Sco89, Van87, You87a]. Dordrecht-Boston-Lancaster [Kwi87]. dory [Cla85]. Douglas [Low82b]. Dover [You87a, Lav87]. Dowanhill [Wat82c]. Dr [Gil85]. Dr. [Kin80a]. draft [Kim81b, Ogl81, Rei81]. Drafting [Tre81]. dragger [Sin86]. dragon [Wat84c]. drawing [Ale88]. drilling [For80b]. drive [Bin89]. Dropping [MSV85]. Dublin [Bir85, McB84, Sal85]. Dumping [BC83a, BG87, Bur88]. Dunn [Law85]. Düsseldorf [CS82]. Dutchmen [Coh82]. Dyke [Cou89, Kim86a]. dynamics [Abr84a, Mor80].

Earthscan [Sut81a]. Earthscan/International [Sut81a]. East [Gla89f, Ram88, Spa87, For81b, Lon81, hP81, Mor84]. East-West [Ram88]. Eastern [Vuk85, Wat81f, Jos80]. ecology [Hoy85]. Ecomomic [Kwi88a]. Economic [Bar82, FJ85, Low84, Moo86, Pan80, Pow80, Sco89, Bir87c, CL82, Cru86, Fra89, Hol88a, Kwi87, MT82, Ogl81, Øst86, hP81, Sil85, Smi83, Sto80b, Wat87a, WY85, GO89]. Economics [Gon80, Gul89a, Roe81, Gul87d, Joh84, Met80, Mun88, Oma83, Smi82, Law85, App85]. économique [Moo86]. economists [Met84]. economy [For85a]. ecotoxicology [WC88]. Ecuadorian [RBA87]. ed [Gol87a]. Edgar [Kwi89a, Ash82]. Edinburgh [Bar85]. editad [Ale85]. Edited [Arc81a, Kwi84, Ale89, And89a, Aub86, AC82, Bea88, BC83b, Bol89, CS85, Cou89, For85a, Gla86, Gla89d, Gla89h, Gla89b, Gla89g, Gul86, Hoy85, Hum83a, Hum87a, Kim86a, Kwi85, Kwi89a, Law83, Len80, Lia82, Low85, Mar88b, Mar88c, Mor88b, Mun86, Mun88, Nai81, Ogl84, Pla86, Pri80, Ran82, Sha82a, Sha82b, Smi83, Van87, Van85, Wat80a, Wat81c, Wat82b, Wat85c, Wat85a, Wat86a, Wat86b, Wat87d, Wat87h]. edition [Gul85, Gul87d]. Editions [Sha80b]. Editor [Kwi89b, Ano87g, CM81, Gin86]. Editore [Gla89b]. Editorial [Ano86h, Ano89g, Ano80e, Ano81e, Ano82e, Ano83e, Ano84e, Ano85e, Ano87e, Ano88e]. education [Ano89a, Gol89]. Edward [Mun88]. EEC [Chu83, Cun80, Ste84]. EEZ [Ano87g, Fra87, Kac89, Lof81, Pla85b]. EEZs [Gul84b]. effective [Wat80a]. Effects [CDW80, Wij80, Bol89, Cop80, Bin89, Mar87]. Efficacy [Mer83]. efficient [Fra85]. Effort [Gul83, CDW80, CW80]. EFJ [Mun85]. Eighth [Geo82, Phi87, Wat84c]. El-Hakim [Wat81f]. eleventh [Kim82a]. Elisabeth [Wat82b, Wat85a, Wat86b]. Ellon [Arc81b]. Elmira [Gol87b]. Elsevier [Bol89]. emergency [You86]. Emerging [Fri80, For85a]. Emond [Pen84]. Emond-Montgomery [Pen84]. empire [Sla84]. employment [GS80]. enclosed [Edm82]. enclosure [Pon88]. end [BC86]. Energy [Geo82, Gil86, Mar89a, Mar89b, Mar89c]. Enforcement [Dre88, Gil88, LS85, Ros86]. Engineering [For81a, Wat87a]. England [HDG81]. enigma [McE82c]. Enterprise [Ran82, Bin89, Kir81]. Entry [Pow80, Ann83, Lia82]. Environment [Aub84, Bar89, Mit83, Sal85, Sut81a, Ano84f, Bir80b, HC86, Lon81, Seb88, SK88, Sha82b, Sil85, Wal89, tH87, Wat87h]. Environmental [BC89, Geo83, Hay84, Men84, AC82, Bol89, Lam87, Lon81, Low82b, Mun86, WC88]. environs [You84]. Epping [Pap84]. equiratio [Lan86]. era [Lof81]. Erich [Low82b]. Ernst [Abr83, Stu88]. Esbjerg [Gul89a]. ESC [Stu85a]. Esra [Heg81]. Essays [Gla89e, Mun88, Bir83b]. Essex [Pap84]. estuarine [Kwi89a]. étrangers [Moo86]. Eugene [CS85]. Europe [Vuk85, BC84, Men88b]. European [Mit83, Bir87c, Lei85, Nag80, Chu84]. evaluated [Har87]. evaluation [Law80]. evalued [Low82a]. evolution [Ash82, Bir85, Kin80b, Ell83]. examination [Kwi89a]. exchange [Mil83].

Exclusive [Sco89, Hol88a, Moo86, Sco85, Wat87a, Moo86, Kwi87]. exhibition [Man80]. Expanding [Kin80b]. Expansion [LP87]. expendable [Kir81]. experience [Cre83, Yah81]. experiences [Law86]. Expert [Gul83]. Explaining [Sco85]. exploitation [Eva86, Man80, Moo86]. exploration [Eva86]. export [Kac84]. Extended [Cru86, QJ89, Ram88]. extension [Tie84]. extent [McM85].

F [Bol89]. Factors [McM85]. Falkland [BB83, McE83c]. Falklands [Chu88, Faw82, Gul87e, McE83b]. fantasies [Boo81]. FAO [Bir80c, Gon80, Loc84, Mad85b, Lof81, Mad85a]. FAO/Wiley [Loc84]. Farnborough [Ash82, McC80c, Nai81]. Farnham [Gul85, Law86]. favouritism [JM87]. FCMA [Lar80]. February [Ano89a, Bir81c, Fab85, Gre80, Hum87c, PM85, Pow80, Ree89b]. Federal [TY88, CS85]. Federation [War85]. Ferry [Bro80, Bro83a, For80b, Low81a, Pla86]. feud [Wat85b]. Ffooks [Man81]. field [Wah83]. Fields [Law84]. Fifth [Kim88b]. Film [Wat81d, Wat82d, Wat82c]. finalized [Sco82]. Finance [Arc83]. financial [Ano88g]. Finn [Arc83, Mor88a, Ogl84]. FIRM [Bir80c]. FIRM/R [Bir80c]. First [Kwi88b, Low81b, Pul80b, Ram88, Ran82]. Fish [Pau82, CG83, Cru85, GGM86, MBA87, McE82a, WGG82, Wat85b, You82, Gul85, Loc84]. Fisheries [App83a, Bir81a, Due80, God80, Gul84a, Gul89a, Jen89, Mar88a, Pow80, Sco89, Ste84, App85, Ber86b, Bir80c, Bir87c, CC81, Chu83, Chu84, CS85, Cla85, Cop80, Cop81, Cop82a, Cop82b, Cre83, Cru86, Cun80, CW80, Dah88, Dem85, Dow84, FJ85, Goo80, Gul84b, Gul87d, Gul87a, Gul87b, HDG81, HS88a, Hu87, JM87, Joh84, Jos80, Kee83, Law80, Law82, LR83, Law83, Law88, Lei85, LW87, Lof81, Mad85a, MB81, Mat86, MT87, McE83c, McE84, McE89, Mun86, Ols82, Pon88, QJ89, Rob80, RL86, Sto80a, Sto80b, Tie84, WY85, Yah81, Dad85, Mad85b, Mor81a, AC82, Gul86, Cop85, Law85]. fishermen [Bin89, DT86, Hee88, Jen89, Pan80, Sha81, TD86]. Fishery [Bal81, Bal82, CS85, Fin85, Lun80, ST86, SK82, Tur85, Ann83, Bla83, Fra87, Jen85, Lar80, LS85, Lia82, Lie86, Mor80, Sco85, Law86]. Fishing [CW80, Gul83, Gul85, Law86, Meu86, Pea81a, Pow80, Bil86, Chu88, CDW80, For81b, GS80, Gul87e, Kac89, McC80b, QJ89, Smi81, Sto80a, Tei87, Law84]. fishworkers [Cia85]. FL [Wec87]. Flag [War85]. Flags [Met81]. fleet [LP87, Meu86, Sin86, Sma84]. fleets [Dou87, Kac89, Mun85]. flexibility [Gul80a]. Floating [Kin83]. flying [Coh82]. focus [Mar88a]. Food [Gla89e, Gon80, Loc84, Mad85b]. force [Wat87c]. Ford [Len87]. Foreign [Ano88g, Kac89, Nai82, Mar87, Sla84, Sto80a, Sto80b]. foreigners [Moo86]. form [Coh82]. formation [LW87]. former [Ols82, Spe82]. Formulation [Vol86, Sut82a]. Forum [Ano89a, Edm82, Sha82a, Bar85, Mcc80a, Til84, You86]. forward [Seb80]. Fossey [Abr84b]. Foster [Gla89c]. four [Hu88]. Fourth [Met84, Gul89a, Kim87, Man80, For81a]. FR [CS82]. Fragmentation [Men88a]. framework [Böh84, Pen84]. France [Bir80d, Man80, Wat82c].

Frances [App85, Gol87b, Len87]. Francesco [Gla89b]. Francioni [Gla89b]. Francis [Wat80a]. Frankel [Abr83, Stu88]. fraud [CM81]. freedom [Low82a]. French [Car88, Meu86, Moo86, Sha80b, Sha81, SGG85]. fresh [You82]. Fridtjof [Wat84a]. Frozen [Gla89a, Aub84]. fruits [Bro81a]. functional [Fab85]. functions [Fra84]. further [TD86]. future [Bir87a, Edm82, GFGJ83, Gil85, Law84, Len87, McE89, Stu86, TL84, Wat86a, Spe83].

G [Abr83, Alf80, Bar82, Bir82b, Geo89, Gla86, Gul87d, Hum87a, Low81a, Sad81, Sha80b]. G12 [Wat82c]. GA [Ale88]. Gabriel [Stu88]. Galway [Pul82]. gas [Bol89, BL89, Eva86, For80b, Gil87, Man81]. Gavin [Mer83]. General [Kwi89b, Ano84f, Car88]. generated [Eva86]. genetics [Pul82]. Geneva [Kim85a, Kim86b, Kim82b]. Genoa [Fab85]. Gentry [Man81]. genuine [Egi88]. Geoffrey [Cop85, Wat86a]. Geographers [Fab85]. Geographical [Hat88]. geological [Fra85]. geopolitics [Gla89a]. George [Gul86, Low82a, Mor81a, Ran82, Wat87c, Wat81c]. Gerald [Bar82, Bea88, Wil82]. Gerard [Sha82b]. German [Car81]. Germany [CS82, Edm81, TY88]. Gilman [Abr84a]. Ginsburg [Wat82b, Wat85a, Wat86b]. Ginter [Lia82]. Giulio [Wat87h]. Glantz [AC82]. Glasgow [Wat82c]. Glenny [Wat84b]. Global [Mor88a, Bar84, Kim81a, Mer83, Nai82, Sma84]. GLORIA [Ree89a]. Glyn [Len87]. Gold [Ash82, Kwi89a]. golden [Gul84a]. Gollancz [Wat84b]. goods [Gol86, Gol87b]. Goodwin [Fri80]. Gopalakrishnan [For85a]. Gorshkov [Alf80]. Gourlay [Jls89a]. Government [JM87, For81b, Jen89, Sim83, You80, McC80c]. governmental [Bar84]. Gower [Abr84a, Abr84b, Nai81, Wat85b, Bro85]. Graham [Gil86, Mar88b, Mar89a, Wat87a, Mar88c, Pri80]. Grant [Lia82]. Greek [Geo83]. Green [For85a]. Greenwich [Ano89a, Bar85, Edm82, Sha82a, Edm82, You86]. Grotius [Bim86]. groundfish [Bin89, Cru85, Kac84]. Group [For85a, Ran81, Sha82b, Wat87d]. groups [Buz80]. growth [Bir80a, Fri80, Böh89]. guidebook [CS85]. Guiffré [Gla89b]. Guildford [Sha82a]. Guinea [Dou87]. Gulf [Mar88b, Mar88c, MSV85]. Gulland [Loc84].

H [AC82, CS85, Gul86, Hum83b, Law86, McR86, Pen84, Ran81, Van87, Wal86, You85]. habitats [Van85]. Hague [Bir83b, Ogl84, Bir80b, Gil85]. Hakim [Wat81f]. Halifax [Din83, Mun86, Doe89]. Hall [Hum83a]. Hampshire [Len87]. handbook [Wat87d]. Hanks [Nai82]. Hanson [Aub86, Mun86]. Hants [Ash82, Nai81, Wat85b]. hard [Wat81c, Bro81a]. Hardback [Low85, Bir82b, Bir85, Heg81, Roe81, Wat87g, Bim86]. Harmonizing [vHPW85]. haven [Kas87]. Hawaii [Cou89, Edm81, Gla89h, Gla89g, Gul89b, Ram88, Bar82]. Hayuth [Fri80]. hazard [You86]. Hazards [De 83, Cop82a]. hb [Wat87c]. Health [Sal83]. Heaver [Sad81]. held [Ano88g, Pul80a]. Helm [And89a, Bea88, Chu84, Mor88b, Smi82, Stu88, Wat85c, Wat87e, Wat87f].

Henk [Sha82b]. Henry [Gol87b, Wat87d]. Herbert [War85]. Herman [Stu85a]. Herr [Hum83a]. Hershman [Fri80]. Hey [Kwi85]. HI [Mar82a, Mar82b, Ram88]. highly [Tsa86]. Hill [Wat87e]. Historical [CG83, Par88]. history [CL82, Wat84c, McE84, Wat87i]. Hobart [Low84]. Hogwood [McC80c]. Hong [Bir80c]. Honolulu [Bar82, Gla89h, Gla89g, Kim86a, Mar82b, Ram88, Mar82a]. honor [Mun88]. Hoorn [Sha82b]. Hoover [Wat81a]. Hopkins [Gul87d]. Hot [Gil88]. House [Abr83, McC80c, Sha82a]. hybrid [You86]. hydrocarbon [Lon81]. Hydrographic [Bea85].

I. [Low85]. IBM [For83a]. ICLARM [NM84, Pau82]. ICLARM-CSIRO [Pau82]. ICNT [Bal81, Bal82]. II [Low82a, Low85, Mar88c]. IIED [Sut81b]. III [Mer83, Low81a, Bro81a, Buz80, Car88, Cop81, Jin89]. IMO [Bar89]. Impact [Bir86a, FAR81, Gla89h, Cop81, Gul87a, NM84, BC83b]. impacted [Van85]. Imperial [Ree89b]. Implementation [Gri88, Ano88g, Bir86a, Gla89h, Kwi85, LR83, Moo87, Sut82a]. Implementing [Cop82a, Ode84, McR84]. Implications [Mit83, Sco81, Ami82, Hat88, Hu87, KAK88, Ram88, RS80, Sul80, BL89, Ham84, Mil83]. imports [Cru85, HDG81]. Improving [Smi81, Lie86, Lun84]. Inadequacies [Rei81]. incidental [MT82]. incineration [Dit89]. including [Stu85a]. income [TD86, DT86]. incomes [Smi81]. Increasing [MT87, Bar84]. independent [Sul80]. Index [Ano81f, Ano86g, Ano86f, Ano87f, Ano88f, Ano89f, Ros86, You85]. India [Pul80a, For83a]. Indian [Boo81, Buz81, For83b, Kwi88b, Tho81]. Indonesia [Ano84n, Kwi85, Kwi88a, Ham84, McE89]. industrial [McC80c, All82, Bjø88]. Industrializing [SGG85]. industries [GFGJ83]. Industry [Pow80, Bjø88, Coh82, Din83, For81b, GS80, HDG81, McC80b, Shu82, Stu86, Sim83, Stu88]. influence [Mer81c, Nai81, Ste84]. Informal [Buz80]. information [Mil83, Bur88]. INMARSAT [KG88, Lun84, Mer80]. innocent [Jin89]. innovative [Shu82]. Institute [Ano88g, Aub84, Bir86a, Cou89, Din83, Edm81, Gla89h, Gla89g, Gul89a, Kil83, Kim86a, Kim88a, Kwi84, Kwi85, Law84, Low84, Mar82b, Mer83, Mor88a, Mor88b, Nai81, Ram88, Ran82, Sut81a, Wat82b, Wat84c, Wat84a, Wec87, Wec89, You85, Cop82b, Joh84, Mun86, Nai82]. Institution [You86]. Institutional [Mer83]. institutions [Gul88b]. Integrated [Und80, Wat80b, Hoy85, Lev88]. integration [Chu84]. inter [Bro80]. interdiction [Gil89]. interests [Øst86, tHTP82, Hum87a]. Intergovernmental [Gon80]. interim [vH81]. Internationai [Wat82a]. International [Ano84f, Aub84, Bea85, Bir82a, Bir83a, Bir85, Bir86b, Boo80, Cia85, Edm81, Gol87a, Gol89, Gul89a, Hil89, Kim86b, Kwi84, Kwi85, Kwi88a, Law84, Mar82a, Mor88b, Nai81, Phi87, Pla85a, Pul82, Sal85, Stu85b, Sut81a, War85, Wat80a, Wat82b, Wat85c, You87a, And89b, Ano88g, Ano88n, Ano88o, Ash82, Bim86, Car88, Gil85, Gon83, Gul88b, Gul89b, Har87, HS88b, Joh84, Kas89, Kin83, Low81c, Low85, McB84, McE82b, Ogl81, Ogl84, Par88, Pla86, Ram88, Seb88, Sha82a, TL84, Tom84, Wat81e, You81, Ano86h, Ano89g, Arc81b, Bar89, Bir89, BC83a, CS82, Man80, Men84, Nan89, Pul84, Red89, Ric81b, Sch80, Shi88, Bim86, Bir83b, Cou89, Gla89b, Hol88b, Kim88a, Pea81b]. Internationalizing [Bro85]. Interpretation [Dit89]. interwar [Wat81a]. introduction [Ann83, Law85]. investment [Heg81]. IOC [Gon80, Gon83, Mil83, Sul80, Tom82]. Iran [Ami82]. Iran-Iraq [Ami82]. Iraq [Ami82]. Ireland [McB84, Pul82, Sal85, SG83]. ISA [Kim86b]. Island [Aub86, Gon80, Stu86, Wec89, Mor81a, RBA87, ST86, McE83c]. Islands [BB83, All82, Bro80]. Islands/Islas [BB83]. Islas [BB83]. Isles [Jls89b]. issues [And89a, Arc81b, BG87, Kwi85, Kwi89a, Mar82b, Nai81, ST86, Val88, Mor88b, Hum83a]. Italian [Fab85, Sco82]. Italy [Sco82, Seb88]. ITLOS [Kim86b]. IWC [Bir80a, Bir81b, Tom80]. IX [Bar85]. Ixi [Ros86]. Ixv [Van87].

J [Bir87b, Gul86, Lia82, Loc84, Mun86, Nai82, Pea81b, Smi83, Spa87, Wat87e]. J. [AC82, Pri80]. Jack [Gla89a]. Jackson [Ros86]. Jacobson [CS85]. Jacques [Gla86]. Jagota [Bla86]. Jakarta [Kwi85]. Jamaica [Kim84, Kim85b, Kim87, Kim88b, Kim89]. James [Mun88, Ran82, Spe83, Wat87b]. Jan [Chu85]. Jane [Len80, Len80]. Janpath [For83a]. January [Ano80g, Ano81i, Ano82g, Ano83i, Ano84i, Ano85h, Ano86n, Ano87n, Ano88i, Ano89i, Gul83, Kwi88b, Mat86, Mit83, Pau82, Pul80a, SK88, Ano87f]. Japan [Jam82, Nak85]. Japanese [GFGJ83]. Jarashow [Gol87a]. Jay [Lia82]. Jean [Gla89e, Wat82c]. Jersey [McE82b, Nai81]. John [Abr84b, CS85, Hoy85, Law83, Lon81, Van85, Wal86, Wat81b, AC82]. Johns [Gul87d]. Johnson [Bir85]. Johnston [And89a, Low82b]. Joint [App83b, Kac84, Stu85b]. jointly [Ano88g, Stu85b]. Jon [Cou89, Kim86a, Sut81a]. Jonathan [Nai81]. Joseph [Cou89]. Joyner [Gla89d]. Jr [Gla89g, Ran81]. Juda [Pap84]. Judith [Arc81a]. July [Ano80f, Ano81h, Ano82f, Ano83g, Ano84j, Ano85g, Ano86l, Ano87f, Ano87l, Ano88h, Ano89h, Bir82a, Bir83a, Bir86a, Bir86b, Kim88b, Kwi88a, Mad85b, Wat84a, Wec87]. July-21 [Kim88b]. June [Bir87c, Bir89, Din83, Geo82, Mad85b, Met84, Phi87, Wec89, You86]. jurisdiction [Cru86, McD86, QJ89, Ram88, Sco85, Gil86, Mar89a, Mar89b]. Justice [Shi88].

Kailea [Gul89b]. Kailea-Kona [Gul89b]. Kamil [Bir83b]. Kegan [Gul88a]. Ken [Wat87c]. Kent [For85a, Mor81a]. Kerala [Pul80a]. Kester [BC83b]. Kiel [Edm81]. Kildow [Arc81a]. kind [Mcc80a]. King [Kin80a]. Kingdom [Gol87a, Wal86]. Kingston [Aub86, Kim84, Kim85b, Kim87, Kim88b, Kim89]. Kinne [Hoy85]. Kiribati [Tei87]. Kitaoka [Bar82]. Kittichaisaree [Gla89f]. Kluwer [Hum83b].

Koers [Kwi85]. Komer [Til85]. Kona [Gul89b]. Kong [Bir80c]. Kriangsak [Gla89f]. Krieger [Low82a]. Krill [McE82c]. Kronmiller [Low81a]. Kuala [hP81]. Kuwabara [Sal85]. Kuwait [Mat86].

L [Aub86, Bar82, Gla89c, Gol87a, Gul87d, Law86, Len80, Mor89, Ran82, Smi83]. labour [Bro81a, Sla84, Hee88]. Lalwani [You85]. Lamson [Mun86]. Lancaster [Kwi87]. land [Haf81, Sal85, WC88]. land-based [Sal85, WC88]. land-locked [Haf81]. lands [Hol88a]. langoustines [Sha80a]. Lanka [Kwi88b]. last [Chu83]. late [Low85]. Latin [Lun80, Mor81b]. Laursen [Arc83, Mor88a, Ogl84]. Lavigne [Gul86]. Law [Ano88g, Bir86a, CS85, Cou89, Gol87a, Kim84, Kim85a, Kim85b, Kim86b, Kim87, Kwi85, Kwi88a, LT82, Mar82b, Mit83, Moo86, Ram88, Red89, Wat81f, Wat84a, Wec87, Wec89, You85, Ale81, Arc83, Bim86, Bir85, BC83a, Boo80, Bro83a, Buz81, Car88, Cos87, Dek88, Din83, Edm81, GGM86, Gil86, Gil88, Gla89h, Gla89b, Gla89f, Gla89e, Gla89g, Gol89, Gra88, Hag84, Har87, Hum83b, Hum87c, Kim80a, Kim82a, Kim86a, Kim88a, Kwi84, Kwi89a, Kwi89b, LR89, Low82b, Low82a, Low85, Mar89a, Mar89b, Mar89c, McD83, McD85, MT87, Men88a, Moo87, Muk81, Ogl81, Par88, Pla85a, Pla86, Sch80, Sco81, Sco82, Seb80, SG83, Tre83, Van87, Vuk85, Wat81e, Wat87g, Wei84, Wil82, But88b, Kim82b, Mil83, Shi88, Ash82, Bro80, For80b, Gla86, Wal86]. Law [Wat87c]. lawfulness [Low81a]. Lawrence [Pap84]. laws [Low88, Mar88b, Mar88c]. Lawson [App85]. LDC [Mer81a]. LDCs [PK85]. league [App83a]. leasing [Bar82]. Lebensraum [Gla89a]. Lecture [Kwi89a]. legacy [Low81c]. Legal [Bir80d, Gol86, KAK88, Bro80, But88b, Car88, Edm81, Gla89d, Har87, KS85, KG88, Low86, Pen84, RBA87, Sal85, Wat84a, Wat87a, Ami82, Bir87c, Chu88]. Legault [McR86]. legislation [Car81, McD83, Sch80]. Leigh [Chu84]. length [PM85]. length-based [PM85]. lengths [Lav87]. Leonard [McR86, Ran81]. Leone [MB81]. Leslie [Gla89c]. less [Smi83]. Lesser [Kwi89a]. lessons [AC82, Mer83, Yah81]. Letter [CM81, Gin86]. Letters [Ano87g, TS83]. L’Europe [Mor88a]. level [Ber86b, You80]. levels [Böh85]. Lewis [Cou89, Gon80, Hat88]. Lexington [Ash82, Ash82, Wat85b]. liability [KS85, Gri88]. Library [Wat82c]. Libya [Bro83b]. Licensing [Dou87]. lies [Gla89g, Gul84b]. life [Bir81a, Hoy85, SK89]. limit [Cop80]. limitation [CDW80, Gri88, Wat81a]. limitations [SK82]. Limited [Ale85, Pow80, Sal85, Ann83, Lia82]. limits [CG81, Mar89b, RBA87]. Lindsay [Len87]. Liner [Ell83, Ode84, Stu85a, FAR81, Geo88, Pap84, Stu86, Wij80, Moo84]. link [Egi88]. List [Stu85b]. living [TY88]. Lloyd [Ros86, Stu85a, Stu85b, Coh82]. Lloyds [Wal86]. LNG [AW82]. lnterscience [Loc84]. lobster [Ann83, HDG81, Mor80]. Local [Ber86b]. Local-level [Ber86b]. locked [Haf81]. loggerheads [Sha80a]. Lomond [Gla86]. London [Ale85, All82, Ano89a, App85, Arc81a, Aub84, Bar89, Bea88, Bir87b, BC89, Bol89, Chu84, De 83, For81a, Fri80, Gil86, Gol87b, Gol87a, Gre80, Gul86, Gul88a, Law84, Law85, Lei85, Len80, Len87, Lia82, Man81, Mar88b, Mar88c, Mar89a, Mor88b, Pon86, Pri80, Roe81, Ros86, Smi82, Stu85a, Stu85b, Sut81a, Sut81b, Vuk85, Wal86, Wat80a, Wat81c, Wat81b, Wat82b, Wat84b, Wat85c, Wat86a, Wat86b, Wat87a, Wat87b, Wat87c, Wat87e, Wat87d, Wat87g, Wat87f, Wat87i, You86, McE82a, Ree89b]. long [Mun85, Sul80, Bol89]. long-run [Mun85]. long-term [Sul80, Bol89]. longer [Böh86]. Longman [Wat87d]. Looking [Gul84a, Gul87b]. Looseleaf [Vuk85]. losses [Sil85]. low [App83a, Böh85]. Lowe [Wei84]. Ltd [Bim86, Bir85, Gul85, Jls89a, Law85, Pon86, Ros86, Stu85a, Stu85b, Wat87g]. Luc [You87a]. Lumpur [hP81]. Luxembourg [Bir81c]. Lysenko [Wat84b]. Lyster [Bim86].

M [And89a, Aub86, Cou89, Fri80, Gon80, Gul86, Hat88, Hum87a, Kim86a, Low82b, Sco89, Wat80a]. MA [Abr83, Arc81a, Ash82, Kil83, Nai82, Til85, Wat85b]. Macdonald [Len80]. mackerel [WY85]. Macmillan [Spe83, Wat86a, Wat87b]. made [Sha81]. Magnuson [CS85, Fin85, Tur85]. mail [Mar89a, Mar89b, Mar89c]. Maine [MSV85]. Maiolo [Law83]. MAJUIKAN [Yah81]. makers [Cor86]. Making [DT86, And89a, Fra82, Hu87, Mun86, Ric81b, TD86, Wat87g]. Malaysia [hP81, Ham84, MT87, Yah81]. Malaysian [Due80, Oma83]. Maldives [ST86]. Malmö [Phi87]. Malvinas [BB83]. Mammals [Bir80d, Gul86]. manage [Jos80]. managed [Wat87h]. Management [Aub86, For81a, Lar80, Mad85b, Pau82, Pow80, Stu86, WY85, And89b, Ano88g, AC82, Ber86b, Bir80c, CC81, CS85, Cop81, Cop85, Cre83, Cun80, Dah88, Fin85, Fri80, Gul80b, Gul87b, HS88a, HS88b, Hil89, Jen89, LR83, Law88, Lia82, Lie86, MT82, Mat86, Moo87, Mor80, Rob80, SK82, Sul89, Sut81b, Val88, WGG82, Wat80a, You85, Hoy85, Gul87d]. Managing [GO89, Gla86]. Manchester [Wat81f, Wei84, Wat81f]. Manganese [Roe81]. mangroves [Due80]. Manjula [For83a]. Mankabady [Wat85c]. Mann [Wat82b, Wat85a, Wat86b]. Mansell [Law85]. manual [Loc84]. mapping [Fra85]. March [Bar89, Hum87b, Kil83, Kim80a, Kim82a, Kim84, Kim85b, Kim87, Kim88b, Kim89, Man80, Pul82]. margin [McM85]. Marianne [Sha82b]. Maribus [Doe89]. Marine [Alh82, Ano89a, Bar89, Bir80d, BC84, Cha84, De 83, Kwi88b, Mor84, Nan89, Pul80a, Shu84, Sut83, SG83, Tie84, TY88, You85, Ale89, Ano84n, Ash82, Aub86, Bir83b, BC82, BC83b, Cop82a, Cor86, Dah88, Dan88, Eva86, For83b, For85a, GFGJ83, Gon80, Gon83, GO89, HC86, Hum81, Hum83a, Kwi89a, Lam87, Law83, Lev88, Moo87, Ogl84, Pet83, Pon88, Pul80b, Scr83, Seb88, Sil85, SGG85, TL84, TS83, Tom84, Und80, Val88, Vol86, Wal86, Wal89, WC88, You86, CF81, CM81, For80a, Har87, Gla89c, Gul86, Hoy85, Hum83b]. marines [Mor88a]. Maritime [And88, Bar89, Chu85, Lan86, Men88b, Pon86, Wat85c, Wat87i, Ale88, Bir87b, Gla89f, Gol89, Gra88, Gri88, Hum87a, LR89, Low82a, Lun84, Mar88b, Mar88c, Met80, Met84, Nai82, Pap84, Par88, Ram88, Ros86, Wat81c, You80, Hum87b, Stu85b, Ash82, Bea88, Bla86, Til85, Wat87e, Wat87d]. Mark [Spa87, Wat81e, Lei85]. Market [MBA87, Hee88, Kac84]. marketing [McE82a, You82]. Markets [Low84]. Martin [Bir82b, Law85]. Martinus [Bla86, Gil85, Gla89d, Kwi87, Kwi89b, McR86, Mor89, Ogl84, Sco89, Van87, You87a, Bir83b, Kim88a]. Martnus [Gla89c]. Maryland [Gla86]. Mason [Wat80a]. Massachusetts [Kil83]. May [Ale88, Bir89, CS82, Pul84, Sha82a, Stu85b, Wat82c]. Mayen [Chu85]. Mazara [PM85]. MBAI [Pul80a]. McCrea [Fri80]. MD [Gul87d, Wat84c]. Measures [Pow80, Bir80c, Dre88]. Mediterranean [Sal85, Men88b, Sal83, Sco81, Wat82c]. medium [Wat87e]. medunarodno [Kwi84]. Meeting [Bir82a, Bir86b, Bir83a, Bir89, Gul89a, Phi87]. Melbourne [Pow80]. Memoirs [Wat84b]. memory [Gla89e]. Men [Gul86]. mer [Sha80b]. Merchant [Ode84, Scr83, TS83, Ano81p, Sma84]. merican [Ran82]. Metals [For83a]. method [Lan86]. methods [Gul85, Smi82, Loc84]. Methuen [Bir87b, Lei85]. M’Gonigle [Wat81e]. MI [Law83]. Miami [Wec87]. Michael [Ale89, AC82, Chu84, Law85, Wat81e, Wat82a, Wat84b, Law83]. microcosm [Kim81a]. Middle [Wat81f]. migratory [Tsa86]. Milan [Gla89b]. mile [Cop80]. Miles [Mun88, Hu88]. military [Low86, Ano87g]. mineral [Gil83, Gil86, Lan81, Mar89a, Mar89b, Mar89c, Mee83b, RS80]. minerals [SK89]. Mining [CS82, Gre80, Shu82, Bir81c, Bro84, Car81, For83a, Hol85, Kum82, Low81b, Tre83, Low81a, Arc81a]. Minister [Spe82]. ministerial [SK88]. Ministers [Val89]. Ministry [Ano88g, Pow80]. mins [Wat82c]. missed [Bro83b]. missions [Mor82]. mistrust [Low81c]. Mitchell [Aub84, Sut81a, Sut81b]. Mobility [Mul89]. model [Lun80, WGG82]. Models [Jen85]. Modern [Low82a, Edm82]. Modernization [Law83]. Mohamed [Mer83]. Monograph [Hum83a, Nai80a, Wat81a]. Montclair [Nai81]. Montgomery [Pen84]. Moore [Wat81b]. Moorsom [Law84]. mora [Kwi84]. Morales [hP81]. Morgan [Cou89]. Morris [Ale89, Wat82a]. move [Low82c]. Mt [Gla86]. much [CC81, Gul84b]. Multilateral [Nai80a]. multiple [You86]. Multispecies [Pau82]. municipal [Car81]. Mussels [PT84]. must [Nag80]. Myron [Van87].

N [Bol89, For80b, Mor89]. N1 [Law84]. Namibia [Law84]. Nancy [Bol89]. Nansen [Wat84a]. Naples [Sco82]. Narcotics [Gil89]. National [Boo81, McD83, Sal85, Ano84n, Bro80, Gil86, Mar89a, Mar89b, Met84, Sch80, Vol86, LT82, Pow80]. nationalism [Stu86]. Nations [Bro83a, Gla89e, Gon80, Kwi89b, Mer83, Roe81, QJ89, Tei87, Van87]. Natkiel [Wat87i]. natural [Ano84f, Man81, Mun88]. nature [Ano84f]. Naval [Ran81, Wat84c, Moo81, Scr83, Spe83, TS83, Wat81a, Buz81, Boo81]. navies [Dow81, Low80, Mor82, Mor81b, Nai80b]. navigation [Cou89, Sin88]. Navigational [Hat88]. navy [Mcc80a, Spe82, Tho81, Til84]. Nazareth [Pul84]. necessary [HS88b]. need [Ber86a, Dit89, Gul80a, Mar88a, You81, PN84]. needs [FJ85]. negotiating [Buz80, Dup81]. negotiations [Nai80a, Mer83]. net [McE83a]. Netherlands [Bro80, Hum83b, Kim88a, Kwi85, McR86, Mor88a, Roe81, Sco89, Sha82b, tH87, Bir80b, vHPW85]. Newcastle [Met84]. Newfoundland [Lon81, Gil83, Gil84, Gil87, Sin86]. News [Gul85, Law86]. next [Kil83]. Niblett [Len87]. Nice [Moo86]. Nichols [Wat80a]. Nicolson [Wat81b, Wat87i]. Nijhoff [Bir83b, Bla86, Gil85, Gla89d, Gla89c, Kim88a, Kwi87, Kwi89b, McR86, Mor89, Ogl84, Sco89, Van87, You87a]. Nineteenth [Bir86a, Gla89h]. ninth [Kim80a]. Nitze [Ran81]. No [Arc83, Bir80c, Kwi84, Mer83, Sal85, Kwi89a, Sla84, You85]. nodules [For83a, Roe81]. Non [Bar84, KAK88, TY88]. Non-governmental [Bar84]. non-living [TY88]. non-participation [KAK88]. Noordhoff [Bro80]. Nordquist [Van87]. Norman [Wat81b]. North [Bir82b, Sha82a, Sha82b, Wat80a, You85, Sin86, Ben87, Bir80b, BC89, Gul87b, Pee84, PT86, SLB81, WGG82, tHTP82, tH87, vHPW85, Ale89]. North-south [Ale89]. north-west [Sin86, WGG82]. Northern [Arc81b, Mor88b]. Northrup [War85]. northwest [McD86, Sul89, McR86]. Norton [Wat82b, Wat85a, Wat86b]. Norway [Bir87c, Wat84a]. Norwegian [Bjø88, DOS87, Hee88, JM87]. note [Gil88, Hu88, TD86]. Notes [Ano81g, Ano83f, Ano84g, Ano84h, Ano85f, Ano86i, Ano86j, Ano86k, Ano87h, Ano87i, Ano87j, Ano87k, Kwi89a]. Nova [Din83, Doe89, Bin89, MBA87]. November [All82, McE82a, Sco82, Stu86]. Novo [Kwi84]. NS [Mun86]. nuclear [BC83a, Kin83, Van88]. NY [For80b, Low82a, Mer83].

O [Kwi89b]. objectives [Cun80, Cop82a]. obligations [McD83]. obtained [Wat82c]. Occasional [Arc83, Hum87a]. Ocean [Ano88g, Aub86, Bla86, BC83b, Böh83, Böh85, Böh86, Boo81, Bur88, CS82, CS85, Geo82, Gon80, Gon83, Gre80, Hoy85, Kwi88b, Kwi89b, Moo87, Nak85, Roe81, Stu86, Van87, Van88, Bea88, Ber86a, CF81, Din83, Dit89, Doe89, Gla86, Len80, Len87, Mar82b, Pea81b, Ree89b, Sco89, SGG85, Wat87a, Wat87h, Ano84f, Ano88g, Buz81, Gin86, Sut81b, And89a, Bar82, For80b, Wat82b, Wat85a, Wat86b]. Oceana [Bro80, Bro83a, Hol88b, Low81a, Pla86, Vuk85]. OCEANEXPO [Man80]. Oceanic [Sch80, Wat82b, Ale88, For81a, Muk81]. Oceanographic [Gon80]. oceanography [Cor86, Sco82]. Oceans [For80b, Sim83, CF81, Cha84, Edm81, Hoy85, Lon81, Man80, McR84, Mer81c, PT86, Wat80b, Wat87g, Ale85]. O’Connell [Low85]. October [Ano80h, Ano81j, Ano82h, Ano83h, Ano84k, Ano84n, Ano85i, Ano86m, Ano87f, Ano87m, Ano88j, Ano89j, Bir80b, CF81, De 83, Edm81, Gon83, Mar82b, Mer83, tH87, Mar82a, Seb88]. Oda [Kwi89b]. Odeke [Gil85]. Odense [Arc83]. ODR [Sut81b]. OECD [Mer81b]. OETO [Gon80]. Office [Gon80, Roe81]. Offshore [BL89, Hol88a, Bin89, Bol89, Bro81a, Eva86, Gil83, Gil87, Hat88, Haw89, Hum87a, Kas89, Lon81, hP81, RS80, SG83, Lon81]. Ogley [Bro85]. Oil [De 83, Gil87, Spa87, Bir82b, Bol89, BL89, Eva86, For80b, KS85, McC80b, Pri80, Sil85, Van85, Gol87a, Sad81, Wat87f]. old [Dow81, Mor82, Nai81]. only [Sha82a]. Onshore [RS80]. OPCOT [Ree89b]. open [Sha81]. operations [LS85, Sad81]. operations-causes [Sad81]. Opportunities [Gre80, Ree89b, Wat87a]. opportunity [Bro83b]. optimal [Gul80b]. option [Kee83, PT86]. options [MT82, Ran81]. oracle [Mer81b]. oracles’ [Mer81b]. Orbach [Law83]. order [Ogl81, Wat87h, Ogl84]. Ordering [Wat87g]. Oregon [CS85]. ORESU [CS85]. ORESU-H-85-001 [CS85]. Organisation [Bar89, Mad85b]. organised [Stu86]. Organization [Gla89e, Gon80, PT86, Edm82, Mer81c, Wat85c]. Organizations [Lun84, Bar84, Bir83b, Jen89, Kim88a]. organized [Ano88g, Ano89a, Edm82, Stu85b]. Organizing [Cha84]. Oslo [Wat84a]. Osmun [Nai81]. other [Bir82b, Mee83b]. Ottawa [Ano88g]. Otto [Hoy85]. Outer [RBA87, Hol88a]. outlook [Böh86]. overcapacity [Böh82]. overfished [Smi81]. ownership [Dow84, Kee83]. Oxford [Alf80, Bir82b, For83a, Gla89f, Heg81, Jls89b, Low85, hP81, Spa87, Wal89].

P [And89a, Bla86, Low82a, Low85, Wat81c]. PA [War85]. Pacem [Doe89]. Pacific [Mar82b, Mor81a, Ram88, Jos80, Mar82a, Sut82b, Tei87, Til84, Tsa86, For85a]. Packaging [McE83b]. Pages [Ano80i, Ano80g, Ano80f, Ano80h, Ano81k, Ano81i, Ano81h, Ano81j, Ano82i, Ano82g, Ano82f, Ano82h, Ano83j, Ano83i, Ano83g, Ano83h, Ano84i, Ano84j, Ano84k, Ano85j, Ano85h, Ano85g, Ano85i, Ano86o, Ano86n, Ano86l, Ano86m, Ano87o, Ano87n, Ano87l, Ano87m, Ano88k, Ano88i, Ano88h, Ano88j, Ano89k, Ano89i, Ano89h, Ano89j]. Paperback [Low84, Bim86, Bir82b, Wat81c, Heg81, Ran82, Wat87g, Roe81]. Papers [Arc83, Hum87a]. Papua [Dou87]. paradigm [Par88]. Paris [Gon83, Sha80b, Low81b]. parks [Ano84n]. Part [Hoy85, PK85, Bal81, Bal82]. participation [KAK88]. partners [Sim83]. party [Bir80c]. passage [Bor88, Jin89, McD86, McR86, Rei81]. Passenger [Asp88]. Past [Bir87a, McE89]. Patricia [Hol88b, Wat81c]. patterns [CG83]. Paul [Gul88a, Ran81]. pb [Wat87c]. Peace [Hat88, Buz81]. Peaceful [Tsa88, Dek88]. Pealy [Mun88]. Pearson [Abr84b]. Peet [Sha82b]. pelagic [CL82]. Penataran [Kwi88a]. Peninsular [Yah81]. Pennsylvania [War85]. Percy [Sut81b]. Pergamon [Alf80]. perils [Edm82]. perspective [Bar84, Kac89, Mor81b, Sil85, WY85]. perspectives [Ale89, Kin80b, Roe81, Pla86]. Peru [Cos87]. Peruvian [Gul80b]. Perverse [Cop80]. Peter [For80b]. Peterson [Smi83]. petroleum [hP81]. Petros [Mor89]. Pharand [McR86]. Philip [Boo81]. Pierre [Wat82c]. Pinho [Moo86]. Pinter [App85, Gol87b, Len87, Wat80a]. Piracy [Bir87a, Pon86]. place [Sla84]. plan [Lar80]. Planning [Bar82, RS80, SLB81, You85]. plants [Kin83]. plate [McE83a]. platforms [Kas89]. Platzöder [Bro83a]. Plaza [Mer83]. plus [Low85, Ros86]. Plymouth [Bea85]. points [Nai82]. Poisoners [Jls89a]. polar [Bir85, McD86]. policies [Abr83, Böh84, Cru85, For81b, ST86, Hum83a]. Policy [Ano88g, Arc83, Hu88, Hum83a, Mer83, Mor88a, Nai82, Ran82, Van87, You85, Ano80m, Ano80n, Ano81n, Ano81o, Ano81p, Ano82p, Ano82n, Ano82o, Ano83o, Ano83p, Ano83q, Ano84m, Ash82, Asp88, Ber85, Ber86a, Bur88, CF81, Cop82a, Cor86, CW80, DT86, Dit89, Edm81, Ell83, For83b, HDG81, Heg81, Hu87, Kil83, Kin80b, Kwi89a, Lan81, Lei85, Lev88, Lon81, Mar87, McE83c, Meu86, Mun88, Nak85, Pap84, Pet83, Sin86, SGG85, Ste84, Sut82a, Sut83, TD86, TY88, Und80, Vol86, Wat80b, Wil82, You80, vHPW85, Cun80, Ale89, Chu84, Law83]. Political [Hum83a, Bir87b, CC81, JM87, Smi83]. Politics [Wat80a, And89b, Aub86, Ber87, CC81, McC80c, Mor81a, Wat81e, Wat81b, Wat82a]. politique [Mor88a]. pollutants [WC88]. Pollution [Seb80, Alh82, BC82, BC83b, BC84, Boo80, Dre88, Eva86, Gol87a, KS85, Kis88, McD83, Sad81, Sal83, Sal85, Seb88, Sil85, Tom84, Wal86, Wat87f, Pri80, Wat81e, Wat82c]. polymetallic [For83a]. Pontecorvo [Wat87h]. Population [Mor80]. port [Fri80, Ric81a, Heg81]. ports [Ano82p, Edm82, Mer81a, Sid80]. Post [Cre83, Law80, Pla86]. Post-evaluation [Law80]. post-unclos [Pla86]. Potential [hP81, You84]. power [Alf80, Arc83, Boo81, Buz81, Kin83, Nai81, Ran82, Wat81b, Wat86a]. powers [Wat87e]. pp [Ale89, And89a, App85, Arc81a, Ash82, AC82, Bea88, Bim86, Bir82b, Bir83b, Bir87b, Bla86, BC83b, Bro80, Bro83a, Bro85, Cop85, For83a, For80b, Fri80, Geo89, Gla89d, Gla89a, Gla89h, Gla89b, Gol87b, Gol87a, Gul87d, Gul88a, Hat88, Hol88b, Hoy85, Jls89a, Jls89b, Kim88a, Kwi87, Kwi89b, Law83, Law84, Law85, Law86, Len80, Len87, Lia82, Loc84, Low81a, Low85, Man81, Mar88b, Mar88c, McC80c, Moo86, Mor81a, Mor88b, Mun86, Mun88, Nai81, Nai82, Pap84, hP81, Pea81b, Pla86, Ran81, Roe81, Ros86, Sad81, Sha82a, Smi82, Spa87, Spe83, Stu85a, Stu86, Stu88, Sut81a, Van87, Wal89, Wat81f, Wat81c, Wat81b, Wat82a, Wat85b, Wat86b, Wat87a, Wat87b, Wat87c, Wat87e, Wat87d, Wat87h, Wat87g, Wat87f, You85, You87a, Low82b]. pp. [Abr83, Abr84b, Ale85, Arc83, Aub84, Aub86, Bol89, Boo81, Chu84, Cou89, Gla86, Gla89f, Gla89e, Gla89g, Hum83a, Hum83b, Hum87a, Kwi84, Kwi85, Kwi89a, Low82a, Low84, Moo87, Mor88a, Nai80a, Ogl84, Pen84, Sal85, Sco89, Smi83, Sut81b, Til85, Van85, War85, Wat84b, Wat84c, Wat86a, Wei84, Wil82]. Practical [SK82]. Practice [Gol87a, Stu85a, And88, But88a, Red89]. practices [Fri85]. Praeger [BC83b, Gla89a]. prava [Kwi84]. pravo [Kwi84]. precedence [Øst86]. predbeno [Kwi84]. Preface [Din84, EG81]. preliminary [Sto80b]. Premiums [CC81]. Preparatory [Kim88b, Kim89, Kim84, Kim85a, Kim85b, Kim86b, Kim87]. Prepared [Roe81, Gon80, Stu85a]. PrepCom [Ber87, Bro84]. Prescott [Bir87b]. presence [Gil88]. present [Bir87a, Edm82, McE89, Wat81b]. Press [Ale89, Arc81a, Bar85, BC89, Cop85, Fri80, Gla89f, Gul87d, Heg81, Jls89b, Law85, Lia82, Low85, Mor81a, Mun88, hP81, Pea81b, Pon86, Ran81, Ros86, Smi82, Spa87, Stu85a, Stu85b, Stu86, Wal86, Wal89, Wat81f, Wat81e, Wat82a, Wat82b, Wat84c, Wat85a, Wat86a, Wat86b, Wat87h, Wei84, Wil82]. PREST [For80a]. Preston [Wat87i]. preventing [Kis88]. prevention [Pri80]. preview [Pee84]. Price [LW87, Mar88b, Mar88c, Bir82b, Low80]. pricing [Heg81]. Primary [Pow80, Kwi87]. principle [Dek88, Egi88]. Principles [Tsa88]. Prinosi [Kwi84]. priorities [Law82]. Pritchard [Wat87f]. private [AW82]. problem [Ber86b]. problems [Ben87, Dow81, Low86, MT87, Mer81a, Mor82, LR83, Fri80, Smi83, Ran82]. Proceedings [Gla89h, Gla89g, Kwi85, Ano88o, KS85, Stu86, Bar85]. process [Dup81]. processing [MBA87]. Production [CG81, Lan81, Bjø88]. profiles [Ano81a]. Program [Lia82]. Programme [Din83, Moo87, Lof81]. progres [Sim83]. Project [Ano88g, Due80, For80a, Sco82]. projects [Law80]. property [Fri85, Kee83]. proposed [Geo88]. prospect [Edm82]. Prospects [Ano88g, For83a, Sto80a, Böh86, McE84, McE89, Moo87, Pan80, DOS87, Tre83]. protected [Ano84n]. Protecting [WC88]. Protection [Bar89, Hay84, HC86, Lam87, Men84, Sal85]. Protectionism [Gil85]. protest [Sha81]. proucavanje [Kwi84]. provisions [Bal81, Bal82]. Public [Arc83, Hum83a, Ran82, Ano88o, AW82, Asp88, DT86, Gul87c, TD86]. Publications [Ano80j, Ano80k, Ano80l, Ano81l, Ano81m, Ano82j, Ano82k, Ano82l, Ano82m, Ano83k, Ano83l, Ano83m, Ano83n, Ano84l, Ano85k, Ano85l, Ano85m, Ano86p, Ano86q, Ano86r, Ano87p, Ano87q, Ano87r, Ano87s, Ano88l, Ano88m, Ano89l, Ano89m, Ano89n, Bim86, Bro80, Bro83a, For80b, Gla86, Hol88b, Vuk85, Sco89, Bla86, Kwi89b]. published [Nai81]. Publishers [Bla86, Gla89d, Hum83b, Kwi87, McR86, Mor89, Van87, You87a, For83a]. Publishing [Abr83, Bir85, Law85, Sal85]. purposes [Low86]. pursuit [Gil88]. putting [Hum87b].

quality [Bin89, WC88]. quango [God80]. quest [Har87, Wat84c]. question [BB83, Jin89]. Quissac [Bir80d].

R [Bar82, Bir80c, Bir87b, BC83b, Cou89, For83a, Fri80, Gla89h, Gul86, Hum83a, Jls89b, Law83, Law85, Lia82, Low84, War85, Wat81c, Wat81e, Wat87e, Wei84]. R&D [Nak85]. Rabalais [Bol89]. radioactive [BG87, BC86]. Raise [Geo83]. Ramberg [Nai80a]. Rapid [Fra85]. rapporteurs [Ano88n]. rate [Böh85]. rationale [Tho81]. Rationalization [Mer82]. re [Low82a]. re-evalued [Low82a]. reaganomics [Tur85]. reality [GGM86]. Rear [Wat87e]. Reasons [tH87]. Recherche [Wat82c]. Recommendation [Ano84n]. Records [Sha82a]. recoverability [Sil85]. Redefining [Fra84]. redefinition [Sha84]. redistribution [Pon88]. reflections [RL86]. regard [Bir85]. Regime [Gon83, Sco89, Bro80, Edm81, Ham84, Har87, McR84, Muk81, Rei81, Sal85, Sha82a, Sin88, You81, Gla89d, Wat87a]. region [Law82, Sut82b]. Regional [And89a, WGG82, Alh82, BC84, Hay84, JM87, Gon80]. regionalism [Har87, Mor84]. regions [Mor84]. registration [Egi88]. Regulation [Gul83, Alh82, AW82, Bir87c, Hol88b, Tur85, Abr83]. regulations [Fra87, Kin83, Pea81a]. Reidel [Roe81]. Rejoinder [Cab87b, Low87]. related [BG87, hP81, Sal83]. Relations [Kwi84, Law84, But88b]. relationship [Arc81b]. relic [Sla84]. Removal [Kas89]. Renate [Bro83a]. replacement [Meu86]. replies [Kin80a]. reply [Ano87g, LC82, Mad85a]. Report [Gon80, Bir80c, Ano88n, Bir80c]. Reprisals [Par88]. Republic [TY88]. request [Bar82]. Research [Gre80, Mer83, Nan89, PT86, Ran82, Sco82, You85, tHTP82, Ano88g, Dan88, FJ85, Gla86, Hum83b, Law82, Muk81, PN84, Sch80, SG83, TL84]. reservation [Wij80]. resolution [Ano84f]. resort [Sha81]. Resource [Wes82, Arc81b, Mar82a, Mor88b, Mun86, Shu84, TY88, Val88, Wat87a, AC82]. Resources [Gla86, Sal85, Ale88, Ano84f, Aub86, Bir85, CF81, For80a, For81a, Fri85, GGM86, Gil86, Gil87, Lan81, Mar89a, Mar89b, Mar89c, Moo86, Mun88, hP81, RS80, ST86, Smi81, Wat80a, Arc81b, Bea88, Moo87, Sut81a]. response [RBA87]. responsibilities [Dow81]. responsibility [Jen89, Wal89]. ressources [Moo86, Mor88a]. Restoration [Van85]. Restricting [Low82a]. restrictions [Hat88]. resumed [Kim82b]. Ret [Wat87e]. Rettig [Lia82]. revenues [Pea81a]. Review [Abr83, Abr84a, Abr84b, Ale85, Ale89, Alf80, And89a, App85, Arc81a, Arc83, Ash82, Aub84, Aub86, AC82, Bar82, Bar85, Bea88, Bim86, Bir82b, Bir83b, Bir85, Bir87b, Bla86, BC83b, Bol89, Boo81, Bro80, Bro83a, Bro85, Chu84, CS85, Cop85, Cou89, For83a, For85a, For80b, Fri80, Geo89, Gil86, Gla86, Gla89d, Gla89a, Gla89h, Gla89b, Gla89f, Gla89e, Gla89g, Gla89c, Gol87b, Gol87a, Gon80, Gul85, Gul86, Gul87d, Gul88a, Hat88, Heg81, Hol88b, Hoy85, Hum83a, Hum83b, Hum87a, Jls89a, Jls89b, Kim86a, Kim88a, Kwi84, Kwi85, Kwi87, Kwi89a, Kwi89b, Law83, Law85, Law86, Len80, Len87, Lia82, Loc84, Low81a, Low82b, Low82a, Low84, Low85, Man81, Mar88b, Mar88c, Mar89a, Mar89b, Mar89c, McC80c, McR86, Mer83, Moo87, Mor81a]. Review [Mor89, Mor88a, Mor88b, Mun86, Mun88, Nai80a, Nai81, Nai82, Ogl84, Pap84, hP81, Pea81b, Pen84, Pla86, Pon86, Pri80, Ran81, Ran82, Roe81, Ros86, Sad81, Sal85, Sco89, Sha80b, Sha82a, Sha82b, Smi82, Smi83, Spa87, Spe83, Stu85a, Stu86, Stu88, Sut81a, Sut81b, Til85, Van87, Van85, Vuk85, Wal86, Wal89, War85, Wat80a, Wat81a, Wat81f, Wat81c, Wat81e, Wat81b, Wat82a, Wat82b, Wat82c, Wat84b, Wat84c, Wat85b, Wat85c, Wat85a, Wat86a, Wat86b, Wat87a, Wat87b, Wat87c, Wat87e, Wat87d, Wat87h, Wat87g, Wat87f, Wat87i, Wei84, Wil82, You85, You87a, Ber86a, Geo88, Kil83, Wes82, Wij80]. revised [Gul87d]. Rhode [Aub86, Gon80, Stu86, Wec89]. Rhodes [Bar82]. RI [Hat88, Wec89]. Rican [LS85]. Richard [Mar88b, Sut81b, War85, Wat87i, Gol87a, Law84, Mar88c]. Richardson [Gla86]. right [Bor88]. Rights [Ale88, Gil83, Pea81a]. Rijn [Bro80]. Rijswijk [Kwi85]. risk [You86]. RN [Wat87e]. Road [Wat82c]. robbery [Pon86]. Robert [CS85, Kwi87, Low82a, Mun88, Nai82, Til85, Wat81a, Len80]. Robertson [Bir82b]. rock [Ann83, Böh83, Mor80]. Rocks [Cou89]. Roderick [Bro85]. Roger [Pon86, Man81]. Role [Sma84, Ber87, BC83a, Fra89, NM84, PT86, Tho81, Tom82, WC88]. Rome [Gla89e, Gul83, Mad85b, Vuk85, Bir80c]. rooms [Mer83]. Rotterdam [tH87]. Routledge [Gul88a]. Rowan [War85]. Rowena [App85]. Roy [Abr84b]. Royal [Mor88b, Spe82, You86]. Rozakis [Mor89]. Rs [For83a]. rules [Wal89]. run [Mun85]. Ruotsala [Fri80]. Russian [Bir83b, Wat84b].

S [Alf80, Bar82, Geo89, Hum87a, Smi83, You85]. S. [Bla86]. Sachiko [Sal85]. Saetevik [BC89]. Safe [Kas87]. Safety [Bir82b, Edm82, Lam87]. salvage [Coh82]. Samir [Wat85c]. Sandbrook [Sut81b]. Sanger [Wat87g]. Satellite [Cor86, Mer80]. Saunders [And89a]. saving [Hol85]. Saxon [McC80c]. scale [Cla85, Law82, Pan80]. scene [App83b]. schemes [Pla85a, Pla85b]. Schmidt [Low82b]. School [CS85, War85]. Science [And89b, Bol89, For85a, Gre80, Hum83a, Law83, You85, BC83a, Cha84, Gon83, Pul80b, TY88, Wat87a, Gon80]. Scientific [PT86, Sco82, Bir80c, BC82, Dan88, GGM86, Hum83b, SG83]. Scientifique [Wat82c]. scientist [Hum81]. Scotia [Din83, Doe89, MBA87]. Scotian [Bin89, Eva86]. Scotland [Arc81b]. Scottish [Bar85, Wat82c]. Scovazzi [Gla89b]. Scrivener [Mor88b]. Sea [Ano88g, Bir86a, BC89, Cou89, Kim84, Kim85a, Kim85b, Kim86b, Kim87, Kwi85, Kwi88a, Lia82, Mit83, Ram88, Sal85, Sha82a, Sha82b, SLB81, Wat84b, Wat84a, Wec87, Wec89, You85, Abr84b, Ale81, Alf80, Ano89a, Bar85, BG87, Bir85, BC83a, BC86, Bro83a, But88b, Buz81, Cab85, Cos87, DOS87, Dek88, Din83, Dre88, Edm81, Edm82, GGM86, Gil86, Gil89, Gla89h, Gla89f, Gla89e, Gla89g, Gol86, Hag84, Hob83, Hum87c, Kim80a, Kim82a, Kim86a, Kim88a, Kis88, Kum82, Kwi84, Kwi89a, Lan81, LT82, LR89, Low88, Man81, Mar82b, Mar89a, Mar89b, Mar89c, McD85, Men88a, Mer82, Moo87, Muk81, Ogl81, Pla85a, Pla86, Pon86, Ran82, Sco81, Sco82, Sha80b, SG83, Tre83, Van87, Wah83, Wat82a, Wat86a, Ben87, Bir80b, Fab85, God80]. Sea [Gul87b, Kim82b, Kim88a, McD86, Mil83, Pee84, Red89, Shi88, SLB81, Val88, You84, tHTP82, tH87, vHPW85, Bir82b, Wat80a, Arc83, Gol87b, Hum83b, Kwi89b, Low82b, Low84, Low85, Vuk85, Wat81f, Wat81e, Wat87b, Wat87c, Wat87g, Wei84, Gil86, Mar89a, Mar89b, Mar89c, Nai81, Wat81b]. Sea-bed [Gil86, Mar89a, Mar89b, Mar89c]. Sea-use [SLB81]. Seabed [Bro84, Kim86b, Ric81b, Bir81c, Car81, For83a, Low81a, Low81b, Low81c, Nai80a, Shu82, Tre83, Bro85]. seaborne [Böh89]. Seals [Gul87c, Gul87a]. SEAPOL [Ano88n, Ano88o]. seapower [Mcc80a, Wat84c]. searching [Tho81]. seas [Cab87a, Low82a, Low86, Øst86, Ran81, Spa87, Tsa88, You87b, Jls89a]. Seattle [Fri80, Lia82, Mun88]. Second [Bir81c, Gul89b, tH87, Bea85, McB84, Wec89]. Securing [Ran81, Øst86, Ran81]. Security [Mor88b, Arc81b, LT82]. sediments [GO89]. segmentation [MBA87]. Selected [Mar89c, Stu85a]. Seminar [Gre80, Hum87c, McE82a, Pow80, Ano89a, Kwi85, Pul80b, tH87]. Seminary [Wat81a]. Senegal [SK88]. sense [DT86, TD86]. sentinels [PT84]. separation [Pla85a, Pla85b]. September [Arc81b, Bea85, Edm82, Gon83, Kim85a, Kim86b, Kim87, Kim89, Kwi88b, Sim83]. Serebriakov [Bir83b].

Series [Loc84, Nai80a, Sal85, Wat81a]. service [Wat82c]. Services [You86, Tie84]. Session [Bar89, Kim88b, Kim89, Kim80a, Kim81c, Kim82a, Kim87, Kim82b]. setting [MSV85]. settlement [Bro81b, Haw89, Kwi89b]. Seuil [Sha80b]. Sevenoaks [For85a]. Seventh [Bir86b]. Shang [Smi82]. shares [Moo84]. sharing [FAR81, Mar87]. Shearer [Low85]. Shelf [Val89, Bro83b, Eva86, Gil84, Hol88a, McM85, RBA87]. shellfish [McB84, McE82b]. Shetland [Goo80]. Shigeru [Kwi89b]. Ship [Sil85, Boo80]. shipbuilding [McC80c]. Shipping [Fra84, Fra89, Ode84, Stu85b, Ano81p, Ash82, Böh83, Böh84, Böh85, Böh86, Fra82, Gil85, Lam87, Low82c, PK85, Pla86, Ram80, Sha84, Stu88, Ano82p, War85, Abr83, Abr84a, Abr84b, Pea81b, Stu85a, Stu86]. ships [Asp88, Egi88, Ram80, Gol87a]. shoals [Cou89]. shores [Fab85]. Shrimp [Mat86]. Shyam [For83a]. Sicily [PM85]. Siddayao [hP81]. Sidney [Abr84a]. Sierra [MB81]. Sijthoff [Bro80]. Simon [Bim86]. since [Asp88, Pon86]. Singapore [Gla89f, Spa87]. Sino [Hu87]. Sino-Argentine [Hu87]. Sir [Spe83]. situations [hP81]. six [MT82]. Sixth [Kim89]. size [Oma83]. Small [Bim81, Law82, Cla85, Pan80, Sin86, Arc83]. Small-scale [Law82, Cla85, Pan80]. Smith [Kwi87, Low82a, Pri80, Wal89, You85]. Social [Smi83, Bir87c]. Society [Ano89a, CF81, Wat87a, Yat88, Sla84, BC83b]. sociological [Fri85]. softback [Bir85]. Sole [Dow84, Kee83, Lav87]. Solea [Lav87]. Some [Low86, Mer83, RL86, KG88, Mun85]. Sonia [Wat87f]. Soons [Hum83b]. source [Car88, McD83, Sil85]. sources [Sal85]. South [Gla89a, Gla89f, hP81, Mor84, Sut82b, Tei87, Tsa86, You84, Spa87, Ale89]. South-East [Gla89f, hP81, Mor84]. Southeast [Ano88g]. Southern [Sut81b, Tie84, Ano84f]. sovereignty [Tsa86]. Soviet [Nai82, Pla86, Ran81, Ano81p, Bil86, Bla83, But88b, Fra87, Mcc80a, Scr83, Sma84, TS83]. sP10.50 [McC80c]. Space [Dan88, Dek88]. spaces [Fab85]. Spanish [Ric81a]. special [Bir85, Kis88]. species [Tsa86]. Spectre [Böh82]. Spill [De 83]. spills [Van85]. sponsored [Ano89a, Gre80, Pow80, Sco82]. Springer [Kwi89a]. Springer-Verlag [Kwi89a]. Squid [Hu87]. SRC [Gre80]. Sri [Kwi88b]. St [Law85, Lon81]. Stagos [Mor89]. stakes [Aub84]. stand [Buz80, CC81]. standard [Coh82]. standards [Kas89]. Standing [Kwi88b]. State [Pul80a, But88a, Cop80, Kac89, Sin86, Alf80, Wal89]. States [Gol87a, Kim86a, Pap84, Stu86, Wat81f, Ale81, Mun85, Tsa86, BC89, Hat88, Mar88b, Mar88c]. statistics [MB81, Mar88a]. status [tH87]. Stephen [Law85]. Stevens [Gol87a]. still [Böh85, McE82c]. stock [Loc84, PM85, Sul89, WGG82]. Stocks [Pau82, CG83, Tom80]. Stokes [Mun88]. Strait [You87a]. Straits [You87a, Rei81, McR86, Mor89]. Strategic [Boo81, Fra82, Mcc80a, Nai81, Til84, Øst86, WC88]. strategic/economic [Øst86]. strategies [HC86]. strategy [Bir85, LR89, Nai82, Scr83, Sma84, TS83, Til85, Wat87e]. Street [Sut81b]. strong [PN84]. structure [Bir83b, Bjø88]. structures [Wah83]. Studies [Ano88g, Aub86, BC83b, Boo81, Din83, Hum87a, Mer83, Moo87, Mor88a, Nai81, Sad81, Stu86, Kwi89a, Mun86, Mun86]. Study [Kwi84, Ber86b, Bir85, Car81, Cha84, Nai80a, Wil82]. Sturmey [Geo89]. subSaharan [Kac89]. subsidies [JM87]. substandard [Low82c]. successful [Ree89a]. Sudhir [Gla89d]. Suez [Sin88]. Sullivan [Ran81]. Summary [Ano88o]. Sunneva [BC89]. supply [Ram80, Rob80]. support [Ano88g]. Supreme [Gil84]. surface [Mar89a, Mar89b, Mar89c]. surplus [GGM86]. Surrey [Law86, Sha82a]. Surveillance [LS85]. survival [Bir80a, Böh83, Sin86]. Sussex [AC82]. Swan [For80b]. Swanson [Wat84c]. Sweden [Phi87]. Switzerland [Kim82b, Kim85a, Kim86b]. Sydney [Bea88, Wat85c, Hum87c]. symposium [Ano88o, Gul89b, Man80, Pul82, Pul84, Pul80a]. system [Kwi89b, Tie84]. systems [Lie86, Wat84a].

T [Arc81a, Sad81]. tables [Mar89c]. take [PK85]. tale [Wat84a]. tanker [Sad81]. Tankers [Wat81e, Kas87]. Tanzanian [Pet83]. Tasmania [Hum83a]. Tavares [Moo86]. Taxation [Hum83b]. teaching [Wat80b]. Technical [Bea85, Moo81, Ano88n]. techniques [SK82]. technologies [For85b]. Technology [Gon80, Roe81, Yat88, You85, Böh84, CF81, Dan88, Doe89, Edm82, For83b, GFGJ83, Gon83, Kil83, Len80, Ree89b, SGG85, TY88, Wat87a, Len87, Man81]. Technology/The [Yat88]. tendencies [Ric81a]. Tenth [Kim82b, Kim81c]. tenure [Dah88]. term [Böh86, Sul80, Bol89]. terror [Men88b]. text [Stu85a]. Thailand [Ano88g, McD85, MT87, Pan80, Pul80b]. their [Jls89b]. Theodore [Low81a]. theoretical [Cop85]. Theory [Pau82, PM85, Car88]. there [SK89]. Thesis [Moo86]. Theutenberg [Bir85]. Third [Dow81, Kim82b, Gul85, Cla85, Kim80a, Kim82a, Bro83a]. Thirty [Bir86b, Phi87]. Thirty-eighth [Phi87]. Thirty-Seventh [Bir86b]. Thomas [Gla89g]. Thompson [AC82]. Threats [Boo81]. Tiel [Sha82b]. Tilapia [Pul84]. Till [Wat86a]. Tilling [DOS87]. time [Sha84]. Times [Stu86, Ale85]. Tinker [Sut81a, Ano80o]. Tippie [BC83b]. TN15 [For85a]. today [Pon86]. together [Hum87b]. Tomahawks [Hob83]. tool [Cor86, Lia82]. Toronto [Pen84]. Towle [Boo81]. Tozer [CS85]. Trade [Gul89a, Asp88, Böh84, Cru85, FAR81, GS80, Joh84, Mar87, Moo84, Ram80]. Traditional [Dah88, Wah83]. traffic [Pla85a, Sid80, Pla85b]. Training [God80, Mer83, Doe89, Gol89, Tie84]. transboundary [Sul89]. transfer [Doe89, Gon83]. transit [Bor88]. transition [Gol89]. translated [Bir83b, Wat84b]. Transnational [Val88]. Transport [Kwi85, Sad81, War85, AW82, Ash82, Böh89]. transportation [Gol86, You86]. Travails [Coh82]. trawlers [Oma83]. treaties [Bim86]. treatise [Hoy85]. Treaty [PT86, But88a, SK89, Bil86]. trends [KS85, Kim81a, Böh84, ST86]. Tribunal [Kim86b]. Trillo [Len80]. triumph [Kir81]. Tromsø [Bir87c]. Tropical [Pau82, PN84]. Trotman [Gil86, Mar88b, Mar88c, Mar89a, Pri80, Wat87a]. troubled [Spa87]. Tullio [Gla89b]. Tuna [Gul88b, Dou87, HS88b, Hil89, Jos80, LS85, McE89, Mul89].

Tunisia [Bro83b]. Tunisia-Libya [Bro83b]. Turkish [Mor89, Ber86b]. Turner [Bar82]. twentieth [Gla89g, Wec87]. twenty [Hu88, Ram88, Ran82, Hu88, Wec89]. twenty-first [Ram88]. twenty-four [Hu88]. Twenty-second [Wec89]. two [Hol88b, Stu85b, Wat84a]. two-day [Stu85b]. Tycooly [Bir85, Sal85].

UK [Abr84a, Abr84b, And89a, Ano89a, Ash82, AC82, Bim86, Bir82a, Bir83a, Bir86b, Bro85, De 83, Edm82, Gre80, Hoy85, Loc84, Man81, McC80c, McE82b, McE82a, Met84, Pap84, Roe81, Sha82a, Sut81b, Wat81f, Wat85b, You85, Pon86, Bar89, For85a, Gil89, GS80, Gul85, Ode84, Red89, Sid80, Wat82c, WY85, You80, You82]. UK-US [Gil89]. uncertainty [AC82]. UNCLOS [Jin89, Mer83, Pla86, Bir81a, Boo80, Bro81a, Bro84, Buz80, Car88, Cop81, Hum81, Kim80b, Kim81b, Low80, TL84]. UNCTAD [Pap84, Stu86, Mer81c, Ode84, Wij80]. Underwater [Yat88, Wat87a]. UNESCO [Gon83]. Union [Ano84f]. unionism [GS80]. Unit [You85, Gul80a]. UNITAR [Mer83]. United [Bro83a, Gla89e, Gol87a, Gon80, Hat88, Kim86a, Kwi89b, Mer83, Pap84, Roe81, Stu86, You86, Buz80, Van87, Wal86]. University [Ano88g, Arc83, Aub86, Boo81, CS85, Edm81, Gla89h, Gla89f, Gla89g, Gon80, Gon83, Gul87d, Heg81, Hum83a, Kwi84, Lia82, Moo87, Mun86, Nai80a, hP81, Pow80, Ram88, Sad81, Spa87, Stu86, Van87, War85, Wat81a, Wat81f, Wat81e, Wat82b, Wat85a, Wat87h, Wec89, Wei84, You85, Moo86, Wat80b, Fri80, Mun88, Wat86b]. University/IOC/UNESCO [Gon83]. unnoticed [Nai82]. unpromising [Böh86]. Unwin [Gul86, Wat81c, Wat87c]. Update [Ano85n, Kim81c, Pla85b]. upwelling [AC82]. urged [PK85]. US$130 [Bir83b]. US$30 [Wat87h]. US$35.95 [Gla89a]. US$40.00 [Gla89b]. US$85.50 [Gla89d]. USA [Ale88, Arc81a, CF81, CS85, For85a, Hat88, Kil83, Kim80a, Kim88b, Kim89, Mar82b, Nai80a, Ram88, Ran81, Roe81, Stu86, Wec87, Wec89, Bur88, Cru86, Hol88a, Mar82a, Til84, Tsa86]. Use [Fri85, Dek88, Low86, SLB81, Tur85, You85]. used [Tom84]. user [Mar88a]. uses [Ano87g, Sha82b, Tsa88, Wah83, Bir80b, Jls89b]. USSR [Vuk85]. utilization [TY88]. UWIST [Yat88].

V [Bir83b, Bir87b, Sha82a, Wei84]. Valencia [Spa87]. Vallo [PM85]. Vancouver [Sad81]. VanderZwaag [Wat85b]. Variability [Lav87]. Venice [Seb88]. ventures [App83b, Kac84, Stu85b]. Verlag [Kwi89a, Low82b]. Verrier [Sad81]. versus [LT82, Ram80]. vessel [McD83, Oma83]. vessel-source [McD83]. Vessels [Kas87, Sto80a]. Victor [Wat84b]. Victoria [Wat82c]. Victorian [Pow80]. Vietnam [Val89]. view [And88, Kum82]. viewpoint [Haf81]. VII [Edm82, Stu85a]. viii [Stu86]. Vilamoura [PT86]. Villar [Pon86]. violence [Pon86]. Virginia [BC83b, Van87]. visit [Tie84]. Vitali [Bir83b]. Vladil [Wat84b]. Vol [Bla86, Hoy85, Loc84, Mar88c, Nai80a, Sco89, Wat81a]. vols [Low81a]. Volume [Kwi89b, Low85, Mar88b, Mar88c, Ano81f, Ano86g, Ano86f, Ano87f, Ano88f, Ano89f, Gil86, Mar89a, Mar89b, Mar89c, Van87, Wat87a, You87a, Law84, Hol88b, Low81a]. volumes [Hol88b]. voyage [Wat84c]. Vukas [Kwi84].

W [Bir82b, Bro80, Gol87a, Hum83a, Hum87a, Kwi85, Kwi87, Law86, McC80c, Pla86, Sad81, Til85, Vuk85, Wat81e, Wat87d]. W1P [Sut81b]. wake [McD86]. Wales [You85]. Walker [Len87]. Walters [Heg81, Sad81]. war [Ami82, Hol88c, Low88, Hu87, Gul88a]. Ward [Hum87a]. Wardley [Pri80]. Wardley-Smith [Pri80]. warships [Jin89]. Washington [CF81, Fri80, Geo82, Lia82, Mun88, Nai82, Ran82, Sim83, Lia82]. waste [BC83a, BC86, PT86]. wastes [BG87, Van88]. watching [Hol88b]. water [Dou87, Mun85, QJ89, Tei87]. waters [Arc81b, Bil86, Edm82, Hoy85, Jls89b, Mor88b, Spa87]. Watt [Sha82a]. Waugh [Cop85]. way [Abe83, For81a, Seb80]. Weidenfeld [Wat87i, Wat81b]. West [Edm81, Kac89, Moo87, Ram88, Sin86, Car81, LR83, RL86, WGG82]. Westbury [Sha82a]. Western [Ran81, Mor80]. Westport [Gla89a]. Westview [Cop85, Pea81b, Ran81, Smi83, Wat82a, Wil82, Ale89, Mor81a, Smi82]. Whale [Hol82, Hol85, Gul80a, Gul88a, Hol88c, Tom80, Hol88b]. whale-watching [Hol88b]. whales [And89b, Hol88b]. Whaling [Bir82a, Bir83a, Bir86b, Bir89, Phi87, CL82, Hol88b, Sut82a]. Wharton [War85]. Which [Abe83]. Whitmarsh [Law85]. who [CG81]. Wiktor [Gla89c]. Wildlife [Pow80, Bim86, BL89, Bim86]. Wildsmith [Pen84]. Wiley [AC82, Hoy85, Loc84]. Wiley-lnterscience [Loc84]. will [Abe83]. wind [Abe83]. Wise [Lei85]. within [Gil86, Gul84b, Hat88, Mar89a]. workable [McE83c]. worker [War85]. Working [Ran81, Sha82b, Bir80c]. Workshop [Bir80d, Gon83, Pau82, PT86, Ano88n, Bir81c, Bir87c, Lon81, Low81b, Mat86, Stu86, De 83]. World [Ano84n, Dad85, Heg81, Nai80a, Ram80, Rob80, Sha84, Wat81a, Cop81, Gul84b, Mad85a, Pon88, Stu88, Wat80b, Wat84b, Wat87d, Cla85, Dow81, You87a, Bir87b, Gul85, Abr84b]. written [Mar88b].

xii [Len87]. xix [Low81a, Sal85]. xv [Gla86, Gla89f, Gla89c, Wat87h]. xvi [Low84, Doe89]. xviii [Til85]. xx [Bla86, hP81]. xxii [McR86]. xxiii [Bim86, Cop85]. xxvi [Low85]. xxxiii [Wei84]. xxxiipp [Wal86].

Y. [Fri80]. Yamase [Bar82]. Yearbook [Gin86, Wat82b, Wat85a, Kim88a, Wat86b]. Yellow [Val88]. York [Bir87b, BC83b, Bro80, Bro83a, Heg81, Hol88b, Kim80a, Kim82a, Kim87, Kim88b, Kim89, Law85, Low81a, Low82a, Mer83, Pla86, Stu88, Vuk85, Wat80a, Wat87h]. Yugoslavia [Kwi84].

za [Kwi84]. Zacher [Wat81e]. Zagreb [Kwi84]. Zaide [Wat87f]. Zealand [Bir89, Ann83]. Zed [Jls89a, Wat87g]. zone [Buz81, Chu88, Hol88a, Kwi87, Moo86, Moo86, Sco89]. zones [BB83, Wat87a].

References

Abecassis:1983:CBW

[Abe83] D. W. Abecassis. Castillo de Bellver: Which way will the wind blow? Marine Policy, 7(4):238, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900374.

Abrahamsson:1983:BRR

[Abr83] Bernhard J. Abrahamsson. Book review: Regulation and policies of American shipping: by Ernst G. Frankel Auburn House Publishing Company, Boston, MA, 1982, 347 pp. Marine Policy, 7(1):60, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900647.

Abrahamsson:1984:BRC

[Abr84a] Bernhard J. Abrahamsson. Book review: The competitive dynamics of container shipping: by Sidney Gilman, Gower, Aldershot, UK, 1983, 152pp. Marine Policy, 8(3):287–288, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900150.

Abrahamsson:1984:BRW

[Abr84b] Bernhard J. Abrahamsson. Book review: World deep-sea container shipping: by Roy Pearson and John Fossey, Gower, Aldershot, UK, 1983, 288 pp. Marine Policy, 8(3):287–288, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900149.

Ayala-Castanares:1982:BRR

[AC82] Agustín Ayala-Castañares. Book review: Resource management and environmental uncertainty lessons from coastal upwelling fisheries: edited by Michael H. Glantz and J. Dana Thompson, John Wiley, Chichester, Sussex, UK, 1981, 491 pp, £27.05. Marine Policy, 6(2):159–160, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8290077X.

Alexander:1981:DSL

[Ale81] Lewis M. Alexander. The ‘disadvantaged’ states and the law of the sea. Marine Policy, 5(3):185–193, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900464.

Alexander:1985:BRT

[Ale85] Lewis M. Alexander. Book review: The times atlas of the Oceans: editad by Alastair Couper, Times Books Limited, London, 1983, 272 pp. Marine Policy, 9(3):246–247, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900272.

Alexander:1988:ROR

[Ale88] Lewis M. Alexander. Rights to oceanic resources: Deciding and drawing maritime boundaries: Athens, GA, USA, 1–2 May 1987. Marine Policy, 12(1):70–73, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900115.

Alexander:1989:BRN

[Ale89] Lewis M. Alexander. Book review: North-south perspectives on marine policy: edited by Michael A. Morris, Westview Press, Boulder, Co, 1988, 252 pp, $28.50. Marine Policy, 13(3):270–271, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8990064X.

Alford:1980:BRS

[Alf80] Jonathan Alford. Book review: The sea power of the state: S. G. Gorshkov, Pergamon, Oxford, 1979. 290pp. £15.00. Marine Policy, 4(1):82–83, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900342.

Alheritiere:1982:MPC

[Alh82] Dominique Alhéritière. Marine pollution control regulation: Regional approaches. Marine Policy, 6(3):162–174, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900215.

Allen:1982:IIC

[All82] J. E. Allen. Industrial islands conference: London, 17–19 November 1981. Marine Policy, 6(3):241–243, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900288.

Amin:1982:IIW

[Ami82] S. H. Amin. The Iran-Iraq war: Legal implications. Marine Policy, 6(3):193–218, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900239.

Anderson:1988:MDV

[And88] D. H. Anderson. Maritime delimitation: a view of British practice. Marine Policy, 12(3):231–240, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900619.

Andersen:1989:BRO

[And89a] D. H. Andersen. Book review: Ocean boundary making: Regional issues and developments: edited by D. M. Johnston and P. M. Saunders, Croom Helm, Beckenham, UK, 1988, 356 pp, £35.00. Marine Policy, 13(4):359–360, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900262.

Andresen:1989:SPI

[And89b] Steinar Andresen. Science and politics in the international management of whales. Marine Policy, 13(2):99–117, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8990002X.

Annala:1983:ILE

[Ann83] John H. Annala. The introduction of limited entry: the New Zealand rock lobster fishery. Marine Policy, 7(2):101–108, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900210.

Anonymous:1980:Ca

[Ano80a] Anonymous. Calendar. Marine Policy, 4(1):86, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900378.

Anonymous:1980:Cb

[Ano80b] Anonymous. Calendar. Marine Policy, 4(2):ibc, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900494.

Anonymous:1980:Cc

[Ano80c] Anonymous. Calendar. Marine Policy, 4(3):268, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900767.

Anonymous:1980:Cd

[Ano80d] Anonymous. Calendar. Marine Policy, 4(4):340, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900172.

Anonymous:1980:EB

[Ano80e] Anonymous. Editorial Board. Marine Policy, 4(1):ifc, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900184.

Anonymous:1980:PJb

[Ano80f] Anonymous. Pages 158-268 (July 1980). Marine Policy, 4(3):??, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1980:PJa

[Ano80g] Anonymous. Pages 2–86 (January 1980). Marine Policy, 4(1):??, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1980:PO

[Ano80h] Anonymous. Pages 270-340 (October 1980). Marine Policy, 4(4):??, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1980:PA

[Ano80i] Anonymous. Pages 90-156 (April 1980). Marine Policy, 4(2):??, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1980:Pb

[Ano80j] Anonymous. Publications. Marine Policy, 4(1):84–85, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900366.

Anonymous:1980:Pc

[Ano80k] Anonymous. Publications. Marine Policy, 4(2):155–156, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900482.

Anonymous:1980:Pd

[Ano80l] Anonymous. Publications. Marine Policy, 4(3):267, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900755.

Anonymous:1980:RPAa

[Ano80m] Anonymous. Recent policy articles. Marine Policy, 4(3):265–266, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900743.

Anonymous:1980:RPAb

[Ano80n] Anonymous. Recent policy articles. Marine Policy, 4(4):339, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900160.

Anonymous:1980:TC

[Ano80o] Anonymous. Tinker Center. Marine Policy, 4(2):150, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900445.

Anonymous:1981:AP

[Ano81a] Anonymous. Author profiles. Marine Policy, 5(3):290–292, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8190066X.

Anonymous:1981:Ca

[Ano81b] Anonymous. Calendar. Marine Policy, 5(1):95–96, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900920.

Anonymous:1981:Cb

[Ano81c] Anonymous. Calendar. Marine Policy, 5(2):ibc, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8190018X.

Anonymous:1981:Cc

[Ano81d] Anonymous. Calendar. Marine Policy, 5(4):350, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901755.

Anonymous:1981:EB

[Ano81e] Anonymous. Editorial Board. Marine Policy, 5(1):ifc, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900671.

Anonymous:1981:IV

[Ano81f] Anonymous. Index to volume 5. Marine Policy, 5(4):i–iii, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901780.

Anonymous:1981:NA

[Ano81g] Anonymous. Notes for authors. Marine Policy, 5(4):352–353, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901779.

Anonymous:1981:PJb

[Ano81h] Anonymous. Pages 171-292 (July 1981). Marine Policy, 5(3):??, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1981:PJa

[Ano81i] Anonymous. Pages 2–96 (January 1981). Marine Policy, 5(1):??, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1981:PO

[Ano81j] Anonymous. Pages 294-353 (October 1981). Marine Policy, 5(4):??, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1981:PA

[Ano81k] Anonymous. Pages 98-168 (April 1981). Marine Policy, 5(2):??, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1981:Pb

[Ano81l] Anonymous. Publications. Marine Policy, 5(1):92–93, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900890.

Anonymous:1981:Pc

[Ano81m] Anonymous. Publications. Marine Policy, 5(2):166–167, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900166.

Anonymous:1981:RPAa

[Ano81n] Anonymous. Recent policy articles. Marine Policy, 5(1):94, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900907.

Anonymous:1981:RPAb

[Ano81o] Anonymous. Recent policy articles. Marine Policy, 5(4):350– 351, October 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81901767.

Anonymous:1981:SMS

[Ano81p] Anonymous. Soviet merchant shipping policy — recent developments. Marine Policy, 5(1):69–71, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900750.

Anonymous:1982:Ca

[Ano82a] Anonymous. Calendar. Marine Policy, 6(1):88, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8290063X.

Anonymous:1982:Cb

[Ano82b] Anonymous. Calendar. Marine Policy, 6(2):160, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900781.

Anonymous:1982:Cc

[Ano82c] Anonymous. Calendar. Marine Policy, 6(3):255, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900367.

Anonymous:1982:Cd

[Ano82d] Anonymous. Calendar. Marine Policy, 6(4):352, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900203.

Anonymous:1982:EB

[Ano82e] Anonymous. Editorial Board. Marine Policy, 6(1):ifc, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900379.

Anonymous:1982:PJb

[Ano82f] Anonymous. Pages 162-255 (July 1982). Marine Policy, 6(3):??, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1982:PJa

[Ano82g] Anonymous. Pages 2–88 (January 1982). Marine Policy, 6(1):??, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1982:PO

[Ano82h] Anonymous. Pages 258-352 (October 1982). Marine Policy, 6(4):??, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1982:PA

[Ano82i] Anonymous. Pages 90-160 (April 1982). Marine Policy, 6(2):??, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1982:Pb

[Ano82j] Anonymous. Publications. Marine Policy, 6(1):86–87, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900628.

Anonymous:1982:Pc

[Ano82k] Anonymous. Publications. Marine Policy, 6(2):ibc, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900793.

Anonymous:1982:Pd

[Ano82l] Anonymous. Publications. Marine Policy, 6(3):252–253, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900343.

Anonymous:1982:Pe

[Ano82m] Anonymous. Publications. Marine Policy, 6(4):350–351, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900185.

Anonymous:1982:RPAb

[Ano82n] Anonymous. Recent policy articles. Marine Policy, 6(3):253–254, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900355.

Anonymous:1982:RPAc

[Ano82o] Anonymous. Recent policy articles. Marine Policy, 6(4):351– 352, October 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900197.

Anonymous:1982:RPAa

[Ano82p] Anonymous. Recent policy articles: Shipping and ports. Marine Policy, 6(1):76–77, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900525.

Anonymous:1983:Ca

[Ano83a] Anonymous. Calendar. Marine Policy, 7(1):70, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900726.

Anonymous:1983:Cb

[Ano83b] Anonymous. Calendar. Marine Policy, 7(2):137–138, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900350.

Anonymous:1983:Cc

[Ano83c] Anonymous. Calendar. Marine Policy, 7(3):235–236, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900179.

Anonymous:1983:Cd

[Ano83d] Anonymous. Calendar. Marine Policy, 7(4):ibc, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900544.

Anonymous:1983:EB

[Ano83e] Anonymous. Editorial Board. Marine Policy, 7(1):ifc, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900556.

Anonymous:1983:NA

[Ano83f] Anonymous. Notes for authors. Marine Policy, 7(2):139–140, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900362.

Anonymous:1983:PJb

[Ano83g] Anonymous. Pages 142-236 (July 1983). Marine Policy, 7(3):??, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1983:PO

[Ano83h] Anonymous. Pages 238-324 (October 1983). Marine Policy, 7(4):??, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1983:PJa

[Ano83i] Anonymous. Pages 3–72 (January 1983). Marine Policy, 7(1):??, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1983:PA

[Ano83j] Anonymous. Pages 74-140 (April 1983). Marine Policy, 7(2):??, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1983:Pb

[Ano83k] Anonymous. Publications. Marine Policy, 7(1):69–70, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900714.

Anonymous:1983:Pc

[Ano83l] Anonymous. Publications. Marine Policy, 7(2):135–137, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900349.

Anonymous:1983:Pd

[Ano83m] Anonymous. Publications. Marine Policy, 7(3):234–235, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900167.

Anonymous:1983:Pe

[Ano83n] Anonymous. Publications. Marine Policy, 7(4):323–324, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900532.

Anonymous:1983:RPAa

[Ano83o] Anonymous. Recent policy articles. Marine Policy, 7(1):69, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900702.

Anonymous:1983:RPAb

[Ano83p] Anonymous. Recent policy articles. Marine Policy, 7(2):133–134, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900337.

Anonymous:1983:RPAc

[Ano83q] Anonymous. Recent policy articles. Marine Policy, 7(4):323, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900520.

Anonymous:1984:Ca

[Ano84a] Anonymous. Calendar. Marine Policy, 8(1):ibc, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900691.

Anonymous:1984:Cb

[Ano84b] Anonymous. Calendar. Marine Policy, 8(2):199–200, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900940.

Anonymous:1984:Cc

[Ano84c] Anonymous. Calendar. Marine Policy, 8(3):294, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900204.

Anonymous:1984:Cd

[Ano84d] Anonymous. Calendar. Marine Policy, 8(4):ibc, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900393.

Anonymous:1984:EB

[Ano84e] Anonymous. Editorial Board. Marine Policy, 8(1):ifc, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900526.

Anonymous:1984:IUC

[Ano84f] Anonymous. International Union for conservation of nature and natural resources — 1981 general assembly resolution on Antarctica environment and the Southern Ocean. Marine Policy, 8(2):184–197, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900927.

Anonymous:1984:NAa

[Ano84g] Anonymous. Notes for authors. Marine Policy, 8(3):295–296, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900216.

Anonymous:1984:NAb

[Ano84h] Anonymous. Notes for authors. Marine Policy, 8(4):375–376, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900381.

Anonymous:1984:PJa

[Ano84i] Anonymous. Pages 2–80 (January 1984). Marine Policy, 8(1):??, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1984:PJb

[Ano84j] Anonymous. Pages 203-296 (July 1984). Marine Policy, 8(3):??, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1984:PO

[Ano84k] Anonymous. Pages 299-376 (October 1984). Marine Policy, 8(4):??, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1984:P

[Ano84l] Anonymous. Publications. Marine Policy, 8(4):372–374, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8490037X.

Anonymous:1984:RPA

[Ano84m] Anonymous. Recent policy articles. Marine Policy, 8(3):293, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900198.

Anonymous:1984:WNP

[Ano84n] Anonymous. World national parks congress, Bali, Indonesia 11– 22 October 1982: Recommendation on marine and coastal protected areas. Marine Policy, 8(2):198, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900939.

Anonymous:1985:Ca

[Ano85a] Anonymous. Calendar. Marine Policy, 9(1):88, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900934.

Anonymous:1985:Cb

[Ano85b] Anonymous. Calendar. Marine Policy, 9(2):168, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900156.

Anonymous:1985:Cc

[Ano85c] Anonymous. Calendar. Marine Policy, 9(3):253, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900326.

Anonymous:1985:Cd

[Ano85d] Anonymous. Calendar. Marine Policy, 9(4):351, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900557.

Anonymous:1985:EB

[Ano85e] Anonymous. Editorial Board. Marine Policy, 9(1):ifc, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900752.

Anonymous:1985:NA

[Ano85f] Anonymous. Notes for authors. Marine Policy, 9(3):254–255, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900338.

Anonymous:1985:PJb

[Ano85g] Anonymous. Pages 170-255 (July 1985). Marine Policy, 9(3):??, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1985:PJa

[Ano85h] Anonymous. Pages 2–88 (January 1985). Marine Policy, 9(1):??, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1985:PO

[Ano85i] Anonymous. Pages 258-351 (October 1985). Marine Policy, 9(4):??, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1985:PA

[Ano85j] Anonymous. Pages 90-168 (April 1985). Marine Policy, 9(2):??, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1985:Pb

[Ano85k] Anonymous. Publications. Marine Policy, 9(1):87, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900922.

Anonymous:1985:Pc

[Ano85l] Anonymous. Publications. Marine Policy, 9(2):168, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900168.

Anonymous:1985:Pd

[Ano85m] Anonymous. Publications. Marine Policy, 9(4):350–351, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900545.

Anonymous:1985:U

[Ano85n] Anonymous. Update. Marine Policy, 9(1):88, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900946.

Anonymous:1986:A

[Ano86a] Anonymous. Appendix. Marine Policy, 10(1):85, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900485.

Anonymous:1986:Ca

[Ano86b] Anonymous. Calendar. Marine Policy, 10(1):84, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900473.

Anonymous:1986:Cb

[Ano86c] Anonymous. Calendar. Marine Policy, 10(2):163–164, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900254.

Anonymous:1986:Cc

[Ano86d] Anonymous. Calendar. Marine Policy, 10(3):238–239, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900618.

Anonymous:1986:Cd

[Ano86e] Anonymous. Calendar. Marine Policy, 10(4):318–319, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900114.

Anonymous:1986:IVb

[Ano86f] Anonymous. Index to volume 10. Marine Policy, 10(4):323–324, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900138.

Anonymous:1986:IVa

[Ano86g] Anonymous. Index to volume 9. Marine Policy, 10(1):87–88, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900497.

Anonymous:1986:IEB

[Ano86h] Anonymous. International Editorial Board. Marine Policy, 10 (1):ifc, January 1986. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X86900278.

Anonymous:1986:NAa

[Ano86i] Anonymous. Notes for authors. Marine Policy, 10(2):165–166, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900266.

Anonymous:1986:NAb

[Ano86j] Anonymous. Notes for authors. Marine Policy, 10(3):240, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900631.

Anonymous:1986:NAc

[Ano86k] Anonymous. Notes for authors. Marine Policy, 10(4):321–322, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900126.

Anonymous:1986:PJb

[Ano86l] Anonymous. Pages 171-240 (July 1986). Marine Policy, 10(3):??, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1986:PO

[Ano86m] Anonymous. Pages 243-324 (October 1986). Marine Policy, 10(4): ??, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1986:PJa

[Ano86n] Anonymous. Pages 3–88 (January 1986). Marine Policy, 10(1):??, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1986:PA

[Ano86o] Anonymous. Pages 90-166 (April 1986). Marine Policy, 10(2):??, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1986:Pb

[Ano86p] Anonymous. Publications. Marine Policy, 10(2):163, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900242.

Anonymous:1986:Pc

[Ano86q] Anonymous. Publications. Marine Policy, 10(3):239, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8690062X.

Anonymous:1986:Pd

[Ano86r] Anonymous. Publications. Marine Policy, 10(4):318, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900102.

Anonymous:1987:Ca

[Ano87a] Anonymous. Calendar. Marine Policy, 11(1):86, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900522.

Anonymous:1987:Cb

[Ano87b] Anonymous. Calendar. Marine Policy, 11(2):156, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900157.

Anonymous:1987:Cc

[Ano87c] Anonymous. Calendar. Marine Policy, 11(3):253, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900716.

Anonymous:1987:Cd

[Ano87d] Anonymous. Calendar. Marine Policy, 11(4):328, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900303.

Anonymous:1987:EB

[Ano87e] Anonymous. Editorial Board. Marine Policy, 11(1):ifc, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900340.

Anonymous:1987:IVJ

[Ano87f] Anonymous. Index to volume 11 January, April, July and October 1987. Marine Policy, 11(4):333–334, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900339.

Anonymous:1987:LEM

[Ano87g] Anonymous. Letters to the editor: Military uses in the EEZ — a reply. Marine Policy, 11(3):249–250, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900686.

Anonymous:1987:NAa

[Ano87h] Anonymous. Notes for authors. Marine Policy, 11(1):87–88, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900534.

Anonymous:1987:NAb

[Ano87i] Anonymous. Notes for authors. Marine Policy, 11(2):157–158, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900169.

Anonymous:1987:NAc

[Ano87j] Anonymous. Notes for authors. Marine Policy, 11(3):255–256, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900728.

Anonymous:1987:NAd

[Ano87k] Anonymous. Notes for authors. Marine Policy, 11(4):331–332, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900327.

Anonymous:1987:PJb

[Ano87l] Anonymous. Pages 163-256 (July 1987). Marine Policy, 11(3):??, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1987:PO

[Ano87m] Anonymous. Pages 259-334 (October 1987). Marine Policy, 11(4): ??, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1987:PJa

[Ano87n] Anonymous. Pages 3–88 (January 1987). Marine Policy, 11(1):??, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1987:PA

[Ano87o] Anonymous. Pages 90-158 (April 1987). Marine Policy, 11(2):??, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1987:Pb

[Ano87p] Anonymous. Publications. Marine Policy, 11(1):85, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900510.

Anonymous:1987:Pc

[Ano87q] Anonymous. Publications. Marine Policy, 11(2):155, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900145.

Anonymous:1987:Pd

[Ano87r] Anonymous. Publications. Marine Policy, 11(3):252, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900704.

Anonymous:1987:Pe

[Ano87s] Anonymous. Publications. Marine Policy, 11(4):329, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900315.

Anonymous:1988:Ca

[Ano88a] Anonymous. Calendar. Marine Policy, 12(1):79, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900176.

Anonymous:1988:Cb

[Ano88b] Anonymous. Calendar. Marine Policy, 12(2):176, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8890053X.

Anonymous:1988:Cc

[Ano88c] Anonymous. Calendar. Marine Policy, 12(3):323, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900735.

Anonymous:1988:Cd

[Ano88d] Anonymous. Calendar. Marine Policy, 12(4):418, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900346.

Anonymous:1988:EB

[Ano88e] Anonymous. Editorial Board. Marine Policy, 12(1):ifc, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900012.

Anonymous:1988:IV

[Ano88f] Anonymous. Index to volume 12. Marine Policy, 12(4):419–420, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900358.

Anonymous:1988:ODM

[Ano88g] Anonymous. Ocean development and management in Southeast Asia: Prospects for implementation of the UN convention on the Law of the Sea: an international conference jointly organized by the Southeast Asian Project on Ocean Law, Policy and Management, the Institute of Asian Studies, Chulalongkorn University and the Institute of Foreign Affairs, The Ministry of foreign affairs with financial support from the international development research centre, Ottawa, Canada, held in Bangkok, Thailand, 26–30 April 1987. Marine Policy, 12(1):49–54, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900073.

Anonymous:1988:PJb

[Ano88h] Anonymous. Pages 179-323 (July 1988). Marine Policy, 12(3):??, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1988:PJa

[Ano88i] Anonymous. Pages 2–79 (January 1988). Marine Policy, 12(1):??, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1988:PO

[Ano88j] Anonymous. Pages 326-420 (October 1988). Marine Policy, 12(4): ??, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1988:PA

[Ano88k] Anonymous. Pages 82-176 (April 1988). Marine Policy, 12(2):??, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1988:Pb

[Ano88l] Anonymous. Publications. Marine Policy, 12(1):78–79, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900164.

Anonymous:1988:Pc

[Ano88m] Anonymous. Publications. Marine Policy, 12(3):322, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900723.

Anonymous:1988:SICa

[Ano88n] Anonymous. SEAPOL international conference: Report of the rapporteurs on the technical workshop (26–28 April 1987). Marine Policy, 12(1):55–60, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900085.

Anonymous:1988:SICb

[Ano88o] Anonymous. SEAPOL international conference: Summary of proceedings of the public symposium (30 April 1987). Marine Policy, 12(1):61–66, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900097.

Anonymous:1989:BSM

[Ano89a] Anonymous. Britain and the sea: Marine education, seminar sponsored by the Marine Society and organized by the Greenwich Forum: London, UK, 13 February 1989. Marine Policy, 13(4):349–352, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900195.

Anonymous:1989:Ca

[Ano89b] Anonymous. Calendar. Marine Policy, 13(1):84, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900535.

Anonymous:1989:Cb

[Ano89c] Anonymous. Calendar. Marine Policy, 13(2):176, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900134.

Anonymous:1989:Cc

[Ano89d] Anonymous. Calendar. Marine Policy, 13(3):ibc, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900675.

Anonymous:1989:Cd

[Ano89e] Anonymous. Calendar. Marine Policy, 13(4):ibc, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900353.

Anonymous:1989:IV

[Ano89f] Anonymous. Index to volume 13. Marine Policy, 13(4):367–368, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900341.

Anonymous:1989:IEB

[Ano89g] Anonymous. International Editorial Board. Marine Policy, 13 (1):ifc, January 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900365.

Anonymous:1989:PJb

[Ano89h] Anonymous. Pages 178-272 (July 1989). Marine Policy, 13(3):??, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1989:PJa

[Ano89i] Anonymous. Pages 2–84 (January 1989). Marine Policy, 13(1):??, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1989:PO

[Ano89j] Anonymous. Pages 274-368 (October 1989). Marine Policy, 13(4): ??, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1989:PA

[Ano89k] Anonymous. Pages 87-176 (April 1989). Marine Policy, 13(2):??, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:1989:Pb

[Ano89l] Anonymous. Publications. Marine Policy, 13(1):83, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900523.

Anonymous:1989:Pc

[Ano89m] Anonymous. Publications. Marine Policy, 13(2):175, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900122.

Anonymous:1989:Pd

[Ano89n] Anonymous. Publications. Marine Policy, 13(4):365–366, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8990033X.

Appleyard:1983:FAL

[App83a] W. Philip Appleyard. Fisheries aid: low in development league. Marine Policy, 7(3):218–220, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900088.

Appleyard:1983:JVC

[App83b] W. Philip Appleyard. Joint ventures: a changing scene. Marine Policy, 7(4):313–314, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900441.

Appleyard:1985:BRE

[App85] Philip Appleyard. Book review: Economics of fisheries development: Rowena Lawson, Frances Pinter, London, 1984, 283 pp. £18. Marine Policy, 9(1):82, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900879.

Archer:1981:BRD

[Arc81a] A. A. Archer. Book review: Deepsea mining: Edited by Judith T. Kildow, MIT Press, Cambridge, MA, USA and London, 1980, 251 pp, £10.85. Marine Policy, 5(1):89–90, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900865.

Archer:1981:NWR

[Arc81b] Clive Archer. Northern waters: the resource — security relationship: International colloquium on northern waters: Resources and security issues, Ellon, Scotland, 17–19 September 1980. Marine Policy, 5(2):156–157, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900087.

Archer:1983:BRS

[Arc83] Clive Archer. Book review: Small power adaptation: the case of Denmark and the law of the sea: by Finn Laursen, Department of Public Finance and Policy, Odense University, Occasional Papers No 3, 1983, 37 pp. Marine Policy, 7(4):320, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900490.

Ashley:1982:BRM

[Ash82] Andrew Ashley. Book review: Maritime transport the evolution of international marine policy and shipping law: by Edgar Gold, Lexington Books, Lexington, MA, and Farnborough, Hants, UK, 1981, 425 pp, £21.50. Marine Policy, 6(3):251–252, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8290032X.

Aspinwall:1988:PSC

[Asp88] Mark D. Aspinwall. Passenger ships in the coastwise trade: US public policy since 1789. Marine Policy, 12(2):161–164, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900449.

Auburn:1984:BRF

[Aub84] Francis Auburn. Book review: Frozen stakes: B. Mitchell, International Institute for Environment and Development, London, 1983, 135 pp. Marine Policy, 8(3):285–286, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900125.

Auburn:1986:BRA

[Aub86] Francis Auburn. Book review: Antarctic politics and marine resources: Critical choices for the 1980s: edited by L. M. Alexander and L. C. Hanson, Center for Ocean Management Studies, University of Rhode Island, Kingston, 1985, 268 pp. Marine Policy, 10(2):158–159, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900205.

Aronson:1982:UPP

[AW82] Jonathan David Aronson and William Westermeyer. Us public and private regulation of LNG transport. Marine Policy, 6(1): 11–26, January 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900392.

Bainbridge:1982:B

[Bai82] R. C. Bainbridge. Bookwatch. Marine Policy, 6(3):252, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900331.

Bainbridge:1983:Ba

[Bai83a] R. C. Bainbridge. Bookwatch. Marine Policy, 7(1):61–62, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900660.

Bainbridge:1983:Bb

[Bai83b] R. C. Bainbridge. Bookwatch. Marine Policy, 7(2):133, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900325.

Bainbridge:1984:B

[Bai84] R. C. Bainbridge. Bookwatch. Marine Policy, 8(1):73–74, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900642.

Balasubramanian:1981:FPI

[Bal81] S. P. Balasubramanian. Fishery provisions of the ICNT: Part 1. Marine Policy, 5(4):313–321, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81901597.

Balasubramanian:1982:FPI

[Bal82] S. P. Balasubramanian. Fishery provisions of the ICNT: Part 2. Marine Policy, 6(1):27–42, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900409.

Bardach:1982:BRO

[Bar82] John E. Bardach. Book review: Ocean leasing for Hawaii: by Gerald S. Clay with S. Broder, R. Turner, D. S. Kitaoka, G. L. Rhodes and D. K. Yamase, Department of Planning and Economic Development, Honolulu, on request. Marine Policy, 6(1):80–81, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900550.

Barnes:1984:NGO

[Bar84] J. Barnes. Non-governmental organizations: Increasing the global perspective. Marine Policy, 8(2):171–183, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900915.

Barston:1985:BRG

[Bar85] R. P. Barston. Book review: The Greenwich forum IX: Britain and the sea: Scottish Academic Press, Edinburgh, 1984 — Proceedings. Marine Policy, 9(4):343–344, October 1985. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8590048X.

Barston:1989:IMO

[Bar89] R. P. Barston. International Maritime Organisation (IMO), 27th Session Marine Environment Protection Committee: 13–17 March 1989, London, UK. Marine Policy, 13(3):267–268, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900602.

Barston:1983:FII

[BB83] R. P. Barston and P. W. Birnie. The Falkland Islands/Islas Malvinas conflict: a question of zones. Marine Policy, 7(1):14–24, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8390057X.

Boehmer-Christiansen:1982:SBM

[BC82] Sonja Boehmer-Christiansen. The scientific basis of marine pollution control. Marine Policy, 6(1):2–10, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900380.

Boehmer-Christiansen:1983:DNW

[BC83a] S. A. Boehmer-Christiansen. Dumping nuclear waste into the sea: International control and the role of science and law. Marine Policy, 7(1):25–36, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900581.

Boehmer-Christiansen:1983:BRI

[BC83b] Sonja Boehmer-Christiansen. Book review: Impact of marine pollution on society: edited by Virginia K. Tippie and Dana R. Kester, Centre for Ocean Studies, Praeger, New York, 1982, 313 pp, $27.95. Marine Policy, 7(2):130–131, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900295.

Boehmer-Christiansen:1984:MPC

[BC84] Sonia Boehmer-Christiansen. Marine pollution control in Europe: Regional approaches, 1972–80. Marine Policy, 8(1):44–55, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900563.

Boehmer-Christiansen:1986:ERW

[BC86] Sonja Boehmer-Christiansen. An end to radioactive waste disposal ‘at sea’? Marine Policy, 10(2):119–131, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900175.

Boehmer-Christiansen:1989:ECB

[BC89] Sonja A. Boehmer-Christiansen. Environmental cooperation between the North Sea States: Sunneva Saetevik, Belhaven Press, London, 1988. Marine Policy, 13(2):172–173, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900109.

Beazley:1985:SIH

[Bea85] P. B. Beazley. The Second International Hydrographic Technical Conference: Plymouth, 3–7 September 1984. Marine Policy, 9(1):78–79, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900843.

Beazley:1988:BRM

[Bea88] P. B. Beazley. Book review: Maritime boundaries and ocean resources: edited by Gerald Blake Croom Helm, London and Sydney, 256 pp, £27.50. Marine Policy, 12(1):75–77, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900140.

Bentham:1987:NSP

[Ben87] Richard Bentham. The North Sea: problems of de-commissioning. Marine Policy, 11(4):313–315, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900212.

Bergin:1985:RDA

[Ber85] Anthony Bergin. Recent developments in Australia’s Antarctic policy. Marine Policy, 9(3):180–191, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900181.

Bergin:1986:AOP

[Ber86a] Anthony Bergin. Australian ocean policy — the need for review. Marine Policy, 10(2):155–158, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900199.

Berkes:1986:LLM

[Ber86b] Fikret Berkes. Local-level management and the commons problem: a comparative study of Turkish coastal fisheries. Marine Policy, 10(3):215–229, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900540.

Bergin:1987:PPA

[Ber87] Anthony Bergin. The politics of PrepCom: Australia’s role. Marine Policy, 11(4):299–312, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900200.

Bewers:1987:AIR

[BG87] J. M. Bewers and C. J. R. Garrett. Analysis of the issues related to sea dumping of radioactive wastes. Marine Policy, 11(2):105–124, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900030.

Bilger:1986:USF

[Bil86] Cameron Crone Bilger. US–Soviet fishing agreement: Treaty authorizing Soviet fishing in US waters. Marine Policy, 10(1):51–56, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900321.

Bimie:1981:SC

[Bim81] Patricia Bimie. Small cetaceans. Marine Policy, 5(3):277–281, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900610.

Bimie:1986:BRI

[Bim86] Patricia Bimie. Book review: International wildlife law: an analysis of international treaties concerned with the conservation of wildlife: by Simon Lyster, Grotius Publications Ltd, Cambridge, UK, 1985, 470 + xxiii pp, £12 paperback, £25 hardback. Marine Policy, 10(1):75–77, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900400.

Binkley:1989:NSO

[Bin89] Marian Binkley. Nova Scotian offshore groundfish fishermen: Effects of the enterprise allocation and the drive for quality. Marine Policy, 13(4):274–284, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900146.

Birnie:1980:ISG

[Bir80a] Patricia Birnie. IWC — survival and growth. Marine Policy, 4 (1):72–75, January 1980. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X80900299.

Birnie:1980:NSE

[Bir80b] Patricia Birnie. The North Sea and its environment: Uses and conflicts: the Hague, the Netherlands, 25–26 October 1979. Marine Policy, 4(3):254–255, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900652.

Birnie:1980:SBD

[Bir80c] Patricia Birnie. The scientific basis of determining management measures: Report of the ACMRR working party, Hong Kong, 10–15 December 1979 FAO fisheries report No 236 FIRM/R 236 (En), FAO, Rome. Marine Policy, 4(4):336–338, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900159.

Birnie:1980:QWL

[Bir80d] Patricia Birnie. The Quissac Workshop on the Legal Aspects of Conservation of Marine Mammals: Quissac, France, 10–14 December 1979. Marine Policy, 4(3):255–257, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900664.

Birnie:1981:FLA

[Bir81a] Patricia Birnie. Fisheries — life after UNCLOS. Marine Policy, 5(1):2–3, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900683.

Birnie:1981:IBC

[Bir81b] Patricia Birnie. IWC — bargaining and compromise. Marine Policy, 5(1):79–84, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900804.

Birnie:1981:SDS

[Bir81c] Patricia Birnie. Second deep seabed mining workshop: Luxembourg, 10–12 February 1981. Marine Policy, 5(4):334–335, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901639.

Birnie:1982:MIW

[Bir82a] Patricia Birnie. 33rd Meeting of the International Whaling Commission: Brighton, UK, 20–26 July 1981. Marine Policy, 6(1): 74–76, January 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900513.

Birnie:1982:BRO

[Bir82b] Patricia Birnie. Book review: The other price of Britain’s oil: Safety and Control in the North Sea: by W. G. Carson, Martin Robertson, Oxford, 1981, 320 pp, £15 (hardback), £5.95 (paperback). Marine Policy, 6(4):335–337, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900100.

Birnie:1983:MIW

[Bir83a] Patricia Birnie. 34th meeting of the International Whaling Commission: Brighton, UK, 19–24 July 1982. Marine Policy, 7(1): 64–68, January 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900684.

Birnie:1983:BRI

[Bir83b] Patricia Birnie. Book review: International marine organizations essays on structure and activities: Kamil A. Bekiashev and Vitali V. Serebriakov (translated from the Russian by Vitali V. Serebriakov), Martinus Nijhoff, The Hague, 1981, 578 pp, US$130. Marine Policy, 7(3):224–226, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900118.

Birnie:1985:BRE

[Bir85] Patricia Birnie. Book review: The evolution of the law of the sea: a study of resources and strategy with special regard to the polar areas: B. B. Johnson Theutenberg Tycooly International Publishing Ltd, Dublin, 1984, hardback £16.50, softback £8.50. Marine Policy, 9(1):82–84, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900880.

Birnie:1986:CLS

[Bir86a] Patricia Birnie. The UN Convention on the Law of the Sea: Impact and implementation: Nineteenth Annual Conference of the Law of the Sea Institute: Cardiff, 24–27 July 1985. Marine Policy, 10(1):69–74, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900382.

Birnie:1986:TSM

[Bir86b] Patricia Birnie. Thirty-Seventh Meeting of the International Whaling Commission: Bournemouth, UK, 15–19 July 1985. Marine Policy, 10(1):62–66, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900357.

Birnie:1987:PPP

[Bir87a] P. W. Birnie. Piracy: Past, present and future. Marine Policy, 11(3):163–183, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900546.

Birnie:1987:BRM

[Bir87b] Patricia Birnie. Book review: The maritime political boundaries of the world: J. R. V. Prescott, Methuen, London and New York, 1985, 377 pp, £30. Marine Policy, 11(1):79–80, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900443.

Birnie:1987:EWR

[Bir87c] Patricia Birnie. European workshop on the regulation of fisheries: Legal, economic and social aspects: Tromsø, Norway, 2–4 June 1986. Marine Policy, 11(2):150–152, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900108.

Birnie:1989:MIW

[Bir89] Patricia Birnie. 40th meeting of the International Whaling Commission: Auckland, New Zealand, 30 May–3 June 1988. Marine Policy, 13(1):68–72, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900420.

Bjorndal:1988:NAI

[Bjø88] Trond Bjørndal. The Norwegian aquaculture industry: Industrial structure and cost of production. Marine Policy, 12(2):122–142, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900401.

Bolze:1989:OOG

[BL89] D. A. Bolze and M. B. Lee. Offshore oil and gas development: Implications for wildlife in Alaska. Marine Policy, 13(3):231–248, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900572.

Black:1983:SFA

[Bla83] William L. Black. Soviet fishery agreements with developing countries: Benefit or burden? Marine Policy, 7(3):163–174, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900027.

Blake:1986:BRM

[Bla86] G. H. Blake. Book review: Maritime boundary: S. P. Jagota, Publications on Ocean Development Vol 9, Martinus Nijhoff Publishers, Dordrecht, 1985, xx + 388 pp, £48.50. Marine Policy, 10(4):317–318, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900096.

Bohme:1982:SO

[Böh82] Hans Böhme. Spectre of overcapacity. Marine Policy, 6(4):330– 331, October 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900070.

Bohme:1983:OSS

[Böh83] Hans Böhme. Ocean shipping: survival at rock bottom. Marine Policy, 7(3):223–224, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900106.

Bohme:1984:CFS

[Böh84] Hans Böhme. The changing framework of shipping: Trends in trade, technology, and policies. Marine Policy, 8(3):229–238, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900034.

Bohme:1985:OSM

[Böh85] Hans Böhme. Ocean shipping: more cargoes but rate levels still low. Marine Policy, 9(4):337–338, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900430.

Bohme:1986:OSU

[Böh86] Hans Böhme. Ocean shipping: unpromising outlook for 1986 — better prospects for the longer term. Marine Policy, 10(3):230–231, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900552.

Bohme:1989:STA

[Böh89] Hans Böhme. The seaborne transport of automobiles: Growth, change and adjustment. Marine Policy, 13(2):155–165, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900055.

Bolze:1989:BRL

[Bol89] Dorene Bolze. Book review: Long-term environmental effects of offshore oil and gas development: edited by Donald F. Boesch and Nancy N. Rabalais, Elsevier Applied Science, London, 1987, 708 pp. Marine Policy, 13(4):360–361, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900274.

Booth:1980:ISP

[Boo80] Ian J. Booth. International ship pollution law: Recent developments at UNCLOS. Marine Policy, 4(3):215–228, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8090055X.

Booth:1981:BRN

[Boo81] Ken Booth. Book review: Naval power in the Indian Ocean: Threats, bluffs and fantasies: Philip Towle, The Strategic and Defence Studies Centre, Australian National University. Canberra, 1979, 121 pp. Marine Policy, 5(1):86–87, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8190083X.

Bordunov:1988:RTP

[Bor88] V. D. Bordunov. The right of transit passage under the 1982 convention. Marine Policy, 12(3):219–230, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900607.

Brown:1980:BRL

[Bro80] E. D. Brown. Book review: The legal regime of islands in inter national law: Derek W. Bowett, Oceana Publications Inc, Dobbs Ferry, New York, Sijthoff and Noordhoff, Alphen aan den Rijn, Netherlands, 1979, 337 pp, $30.00. Marine Policy, 4(4):333, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900111.

Brown:1981:DOA

[Bro81a] E. D. Brown. Delimitation of offshore areas: Hard labour and bitter fruits at UNCLOS III. Marine Policy, 5(3):172–184, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900452.

Brown:1981:DS

[Bro81b] E. D. Brown. Dispute settlement. Marine Policy, 5(3):282–286, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900634.

Brown:1983:BRT

[Bro83a] E. D. Brown. Book review: Third United Nations conference on the law of the sea: Documents: by Renate Platzöder, Oceana Publications, Dobbs Ferry, New York, 1982, 537 pp, $100. Marine Policy, 7(2):132–133, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900313.

Brown:1983:TLC

[Bro83b] E. D. Brown. The Tunisia-Libya continental shelf case: a missed opportunity. Marine Policy, 7(3):142–162, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900015.

Brown:1984:SMU

[Bro84] E. D. Brown. Seabed mining: From UNCLOS to PrepCom. Marine Policy, 8(2):151–164, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900897.

Brown:1985:BRI

[Bro85] E. D. Brown. Book review: Internationalizing the seabed: Roderick C. Ogley, Gower, Aldershot, UK, 1984, pp 280, £16.50. Marine Policy, 9(3):243–244, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900247.

Burroughs:1988:ODI

[Bur88] R. H. Burroughs. Ocean dumping: Information and policy development in the USA. Marine Policy, 12(2):96–104, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900371.

Butler:1988:CTS

[But88a] W. E. Butler. Custom, treaty, state practice and the 1982 convention. Marine Policy, 12(3):182–186, July 1988. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900553.

Butler:1988:LSA

[But88b] W. E. Butler. The Law of the sea and Anglo — Soviet legal relations. Marine Policy, 12(3):179–181, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900541.

Buzan:1980:UWS

[Buz80] Barry Buzan. ‘United we stand … ’: Informal negotiating groups at UNCLOS III. Marine Policy, 4(3):183–204, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900536.

Buzan:1981:NPL

[Buz81] Barry Buzan. Naval power, the law of the sea, and the Indian Ocean as a zone of peace. Marine Policy, 5(3):194–204, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900476.

Cable:1985:CS

[Cab85] James Cable. Conflict at sea. Marine Policy, 9(4):261–268, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900351.

Cable:1987:CBS

[Cab87a] James Cable. Closing the British seas. Marine Policy, 11(2):90–96, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900017.

Cable:1987:R

[Cab87b] James Cable. Rejoinder. Marine Policy, 11(4):327, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900297.

Caron:1981:DSM

[Car81] David D. Caron. Deep seabed mining: a comparative study of US and West German municipal legislation. Marine Policy, 5(1):

4–16, January 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900695.

Carty:1988:TCT

[Car88] Anthony Carty. Towards a critical theory of general custom as a source of international law: UNCLOS III in recent French legal theory. Marine Policy, 12(3):211–218, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900590.

Chatterton:1981:HMP

[CC81] Lynne Chatterton and Brian Chatterton. How much political compromise can fisheries management stand?: Premiums and politics in closed coastal fisheries. Marine Policy, 5(2):114–134, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900026.

Cunningham:1980:EFE

[CDW80] Stephen Cunningham, Michael Dunn, and David Whitmarsh. Effects of fishing effort limitation. Marine Policy, 4(4):328–329, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900081.

Cameron:1981:OPR

[CF81] Hugh Cameron and Glyn Ford. 1980s’ oceans policy — resources and technology: ‘The decade of the oceans’, 16th annual conference of the Marine Technology Society, Washington, DC, USA, 6–8 October 1980. Marine Policy, 5(2):159–161, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900105.

Cameron:1981:PLW

[CG81] Hugh Cameron and Luke Georghiou. Production limits — who benefits? Marine Policy, 5(3):267–270, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900567.

Caddy:1983:HPF

[CG83] J. F. Caddy and J. A. Gulland. Historical patterns of fish stocks. Marine Policy, 7(4):267–278, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900404.

Charnock:1984:MSO

[Cha84] H. Charnock. Marine science: Organizing the study of the oceans. Marine Policy, 8(2):120–136, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900873.

Churchill:1983:EFA

[Chu83] Robin Churchill. EEC fisheries: agreement at last. Marine Policy, 7(2):74, April 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900180.

Churchill:1984:BRE

[Chu84] R. R. Churchill. Book review: European integration and the common fisheries policy: by Michael Leigh, Croom Helm, London, 1983, 228 pp. Marine Policy, 8(3):288–289, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900162.

Churchill:1985:MDJ

[Chu85] R. R. Churchill. Maritime delimitation in the Jan Mayen area. Marine Policy, 9(1):16–38, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900776.

Churchill:1988:FFZ

[Chu88] R. R. Churchill. The Falklands fishing zone: Legal aspects. Marine Policy, 12(4):343–360, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8890019X.

Ciaireaux:1985:ICF

[Cia85] Henri Ciaireaux. International conference of fishworkers. Marine Policy, 9(4):338, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900442.

Clark:1982:EHA

[CL82] C. W. Clark and R. Lamberson. An economic history and analysis of pelagic whaling. Marine Policy, 6(2):103–120, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900653.

Claireaux:1985:DSS

[Cla85] Henri Claireaux. The dory and small-scale fisheries in the Third World. Marine Policy, 9(3):250–252, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900302.

Cormack:1981:LEM

[CM81] P. Cormack and B. T. Martin. Letter to the editor: Marine fraud. Marine Policy, 5(1):95, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900919.

Cohen:1982:TFD

[Coh82] Mark Cohen. Travails of the flying Dutchmen: Lloyd’s standard form of salvage agreement and the US salvage industry. Marine Policy, 6(4):265–286, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900033.

Copes:1980:BCF

[Cop80] Parzival Copes. British Columbia fisheries and the 200-mile limit: Perverse effects for the coastal state. Marine Policy, 4(3):205–214, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900548.

Copes:1981:IUI

[Cop81] Parzival Copes. The impact of UNCLOS III on management of the world’s fisheries. Marine Policy, 5(3):217–228, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8190049X.

Copes:1982:ICM

[Cop82a] Parzival Copes. Implementing Canada’s marine fisheries policy: Objectives, hazards and constraints. Marine Policy, 6(3):219–235, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900240.

Copes:1982:IFA

[Cop82b] Parzival Copes. The institute of fisheries analysis. Marine Policy, 6(4):331–332, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900082.

Copes:1985:BRF

[Cop85] Parzival Copes. Book review: Fisheries management theoretical developments and contemporary applications: by Geoffrey Waugh, Westview Press, Boulder, Colorado, 1984, xxiii + 247 pp. £18.75. Marine Policy, 9(4):345, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900491.

Cornillon:1986:SON

[Cor86] Peter Cornillon. Satellite oceanography: a new tool for marine policy makers. Marine Policy, 10(1):57–60, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900333.

Costa:1987:PLS

[Cos87] Eduardo Ferrero Costa. Peru and the law of the sea convention. Marine Policy, 11(1):45–57, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900388.

Couper:1989:BRI

[Cou89] A. D. Couper. Book review: International navigation: Rocks and shoals ahead?: edited by Jon M. Van Dyke, Lewis Alexander and Joseph R. Morgan, Law of the Sea Institute, Hawaii, 1988, 435 pp. Marine Policy, 13(4):357–358, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900249.

Crean:1983:PEC

[Cre83] K. Crean. Post experience course in fisheries management. Marine Policy, 7(4):322, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900519.

Crutchfield:1985:CFI

[Cru85] Stephen R. Crutchfield. Coping with fish imports: Alternative policies for Canadian — US groundfish trade. Marine Policy, 9 (3):214–219, July 1985. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8590020X.

Crutchfield:1986:EFJ

[Cru86] Stephen R. Crutchfield. Extended fisheries jurisdiction in the USA: an economic appraisal. Marine Policy, 10(4):271–278, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900035.

Cameron:1982:ICO

[CS82] Hugh Cameron and Jonathan Simnett. International Conference for Ocean Mining: Düsseldorf, FR Germany, 15 May 1981. Marine Policy, 6(2):156–157, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900744.

Cicin-Sain:1985:BRF

[CS85] Biliana Cicin-Sain. Book review: Federal fisheries management: a guidebook to the Magnuson Fishery Conservation and Management Act: edited by John Jacobson, Daniel Conner and Robert Tozer, Ocean and Coastal Law Center, University of Oregon Law School, 1985, £5.00 (Copies of this guidebook (ORESU-H-85-001) are available from: Ocean and Coastal Law Center, University of Oregon Law School, Eugene, Oregon 97403, USA). Marine Policy, 9(4):348–349, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900521.

Cunningham:1980:EFM

[Cun80] Stephen Cunningham. EEC fisheries management: a critique of Common Fisheries Policy objectives. Marine Policy, 4(3):229–235, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900561.

Cunningham:1980:FEF

[CW80] S. Cunningham and D. Whitmarsh. Fishing effort and fisheries policy. Marine Policy, 4(4):309–316, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900044.

Dada:1985:WFC

[Dad85] B. F. Dada. World Fisheries Conference. Marine Policy, 9(3): 252–253, July 1985. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X85900314.

Dahl:1988:TMT

[Dah88] Christopher Dahl. Traditional marine tenure: a basis for artisanal fisheries management. Marine Policy, 12(1):40–48, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900061.

Danilenko:1988:STM

[Dan88] G. M. Danilenko. Space technology and marine scientific research. Marine Policy, 12(3):247–255, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900632.

DeBievre:1983:WCM

[De 83] Aline De Bièvre. Workshop on Combating Marine Oil and Chemical Spill Hazards: London, UK, 26–28 October 1982. Marine Policy, 7(2):125–128, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900271.

Dekanozov:1988:PPU

[Dek88] R. V. Dekanozov. The principle of peaceful use in the law of the sea and space law. Marine Policy, 12(3):271–275, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900668.

Demetropoulos:1985:CF

[Dem85] A. Demetropoulos. Cyprus fisheries. Marine Policy, 9(1):69–72, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900818.

Dingwall:1983:LSO

[Din83] Alastair Dingwall. ‘The law of the sea and ocean industry’, 16th annual conference of the Law of the Sea Institute cosponsored by the Dalhousie Ocean Studies Programme: Halifax, Nova Scotia, Canada, 21–24 June 1982. Marine Policy, 7(1):62–64, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900672.

Dingwall:1984:P

[Din84] Alastair Dingwall. Preface. Marine Policy, 8(2):82, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900836.

Ditz:1989:INU

[Dit89] Daryl Ditz. Interpretation of need in US ocean incineration policy. Marine Policy, 13(1):43–55, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900407.

Doelle:1989:PMX

[Doe89] Meinhard Doelle. Pacem in Maribus XVI, ocean technology, training and transfer: Halifax, Nova Scotia, Canada, 22–28 August 1988. Marine Policy, 13(2):166–167, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900079.

Dale:1987:TSP

[DOS87] Eivind Dale, Jonathan Owens, and Åsmund Stenseth. Tilling the sea: Prospects for Norwegian aquaculture. Marine Policy, 11(3): 229–239, July 1987. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X87900595.

Doulman:1987:LDW

[Dou87] David J. Doulman. Licensing distant-water tuna fleets in Papua New Guinea. Marine Policy, 11(1):16–28, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900364.

Dowdy:1981:TWN

[Dow81] Lee Dowdy. Third World navies — new responsibilities, old problems. Marine Policy, 5(2):147–148, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8190004X.

Dow:1984:SOF

[Dow84] Nik Dow. Sole ownership of fisheries: a comment. Marine Policy, 8(4):356–358, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900307.

Drel:1988:EMA

[Dre88] M. I. Drel’. Enforcement measures against pollution of the sea. Marine Policy, 12(3):297–305, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900693.

Davis:1986:MSD

[DT86] Anthony Davis and Victor Thiessen. Making sense of the dollars: Income distribution among Atlantic Canadian fishermen and public policy. Marine Policy, 10(3):201–214, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900539.

Duerr:1980:FMU

[Due80] Christine Duerr. Fisheries and mangroves: US — Malaysian project. Marine Policy, 4(2):151–152, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900457.

Dupuy:1981:NP

[Dup81] RenéJean Dupuy. The negotiating process. Marine Policy, 5(3): 286–287, July 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900646.

Edmondson:1981:OPL

[Edm81] John Edmondson. 1980s’ oceans policy — the legal regime: ‘The law of the sea in the 1980s’, 14th annual conference of the Law of the Sea Institute, University of Hawaii, cosponsored by the Institute for International Law, University of Kiel, Kiel, West Germany, 20–23 October 1980. Marine Policy, 5(2):158–159, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900099.

Edmondson:1982:GVB

[Edm82] John Edmondson. Greenwich VII, ‘Britain and the sea: modern technology and the organization of safety in ports and coastal and enclosed waters, present perils and future prospects’, organized by the Greenwich Forum: Greenwich, UK, 17 September 1981. Marine Policy, 6(1):71–72, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900495.

Edmondson:1981:P

[EG81] John Edmondson and David Green. Preface. Marine Policy, 5 (3):171, July 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900440.

Egiyan:1988:PGL

[Egi88] G. S. Egiyan. The principle of genuine link and the 1986 UN convention on the registration of ships. Marine Policy, 12(3):314–321, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900711.

Ellsworth:1983:LCE

[Ell83] Robert A. Ellsworth. Liner conferences: Evolution of US policy. Marine Policy, 7(4):254–266, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900398.

Evans:1986:CMP

[Eva86] Shelley M. Evans. Control of marine pollution generated by offshore oil and gas exploration and exploitation: the Scotian shelf. Marine Policy, 10(4):258–270, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900023.

Fabbri:1985:SSN

[Fab85] Paolo Fabbri. Sea and shores: New functional spaces: Association of Italian Geographers (AGel) Conference, Genoa, 19–20 February 1985. Marine Policy, 9(3):242–243, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900235.

Frankel:1981:ICS

[FAR81] E. G. Frankel, J. Arnold, and P. Read. Impact of cargo sharing on US liner trade. Marine Policy, 5(1):23–39, January 1981. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900713.

Fawcett:1982:FC

[Faw82] James Fawcett. Falklands conflict. Marine Policy, 6(4):258, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8290001X.

Finch:1985:FMU

[Fin85] Roland Finch. Fishery management under the Magnuson Act. Marine Policy, 9(3):170–179, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8590017X.

Frost:1985:EBD

[FJ85] Hans Frost and Richard S. Johnston. Economic and biological data needs in fisheries research. Marine Policy, 9(1):62–68, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900806.

Ford:1980:PMR

[For80a] Glyn Ford. PREST — Marine resources project. Marine Policy, 4(3):241–242, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900597.

Forster:1980:BRO

[For80b] Malcorn J. Forster. Book review: Ocean oil and gas drilling and the law: Peter N. Swan, Oceans Publications, Dobbs Ferry, NY, 1979, 462 pp, $45.00. Marine Policy, 4(4):335–336, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900147.

Ford:1981:MOR

[For81a] Glyn Ford. Management of oceanic resources: the way ahead, Fourth Engineering Committee on Oceanic Resources: London, 7–9 April 1981. Marine Policy, 5(4):335–336, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81901640.

Forrest:1981:DEC

[For81b] Janet E. Forrest. Development of the east coast fishing industry: Assessment of Canadian government policies. Marine Policy, 5(4):294–301, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81901573.

Ford:1983:BRM

[For83a] Glyn Ford. Book review: Metals from the seabed: Prospects for mining polymetallic nodules by India: by Manjula R. Shyam, Oxford and IBM, New Delhi, 165 pp, 1982, Rs 72 (available from the publishers at 66 Janpath, New Delhi 110001). Marine Policy, 7(4):317–318, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900465.

Ford:1983:IMT

[For83b] Glyn Ford. Indian marine technology policy. Marine Policy, 7 (2):122–123, April 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900246.

Ford:1985:BRE

[For85a] Glyn Ford. Book review: The emerging marine economy of the Pacific: edited by Chennat Gopalakrishnan Ann Arbor Science, Boston, USA, £42.50 (Distributed in Britain by the Butterworth Group, Borough Green, Sevenoaks, Kent TN15 8PH, UK). Marine Policy, 9(3):249–250, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900296.

Ford:1985:CNT

[For85b] Glyn Ford. Cultivating new technologies. Marine Policy, 9(4): 258–259, October 1985. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8590034X.

Frankel:1982:SDM

[Fra82] Ernst G. Frankel. Strategic decision making in shipping. Marine Policy, 6(4):259–264, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900021.

Frankel:1984:SRF

[Fra84] Ernst G. Frankel. Shipping: Redefining functions. Marine Policy, 8(2):165–170, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900903.

Francis:1985:REG

[Fra85] T. J. G. Francis. Rapid and efficient geological mapping. Marine Policy, 9(3):237–242, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900223.

Franckx:1987:NSF

[Fra87] Erik Franckx. New Soviet fishery regulations concerning the EEZ: an appraisal. Marine Policy, 11(2):125–132, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900042.

Frankel:1989:SRE

[Fra89] Ernst G. Frankel. Shipping and its role in economic development. Marine Policy, 13(1):22–42, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900390.

Friche:1980:BRU

[Fri80] Peter H. Friche. Book review: Under new management port growth and Emerging Coastal Management Problems: M. Hershman, R. Goodwin, A. Ruotsala, M. McCrea and Y. Hayuth, University of Washington Press, Seattle and London, 1979, 185 pp + appendixes, £8.75. Marine Policy, 4(3):263, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900718.

Fricke:1985:USD

[Fri85] Peter Fricke. Use of sociological data in the allocation of common property resources: a comparison of practices. Marine Policy, 9(1):39–52, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900788.

Georghiou:1982:EOE

[Geo82] Luke Georghiou. Eighth Ocean Energy Conference: Washington, DC, 7–11 June 1981. Marine Policy, 6(2):157–158, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900756.

Georgandopoulos:1983:GAR

[Geo83] El A. Georgandopoulos. Greek Association to ‘Raise Environmental Consciousness’. Marine Policy, 7(1):55–56, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900611.

Georgandopoulos:1988:PRC

[Geo88] El. A. Georgandopoulos. The proposed 1988 review conference on the code of conduct for liner conferences. Marine Policy, 12(1): 22–30, January 1988. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X88900048.

Georgandopoulos:1989:BRC

[Geo89] El. A. Georgandopoulos. Book review: The code of conduct and the review conference: by S. G. Sturmey, Athens, 1988, 218 pp. £55.00. Marine Policy, 13(4):363–364, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900316.

Georghiou:1983:JNT

[GFGJ83] Luke Georghiou, Glyn Ford, Michael Gibbons, and Glyn Jones. Japanese new technology: Creating future marine industries. Marine Policy, 7(4):239–253, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900386.

Garcia:1986:NLS

[GGM86] S. Garcia, J. A. Gulland, and E. Miles. The new law of the sea, and the access to surplus fish resources: Bioeconomic reality and scientific collaboration. Marine Policy, 10(3):192–200, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900527.

Gilmore:1983:NOM

[Gil83] William C. Gilmore. Newfoundland offshore mineral rights. Marine Policy, 7(3):175–196, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900039.

Gilmore:1984:NCS

[Gil84] William C. Gilmore. The Newfoundland continental shelf dispute in the Supreme Court of Canada. Marine Policy, 8(4):323–329, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900241.

Gilman:1985:PFI

[Gil85] Sidney Gilman. Protectionism and the future of international shipping: Dr Ademuni-Odeke, Martinus Nijhoff, The Hague. Marine Policy, 9(3):247–249, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900284.

Gilmore:1986:BRS

[Gil86] William C. Gilmore. Book review: Sea-bed energy and mineral resources and the law of the sea volume I: the areas within national jurisdiction: by E. D. Brown, Graham and Trotman, London, 1984, £38.00/US $61.00. Marine Policy, 10(1):80–81, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900448.

Gilmore:1987:OGR

[Gil87] William C. Gilmore. Oil and gas resources offshore Newfoundland. Marine Policy, 11(2):145–147, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900078.

Gilmore:1988:HPC

[Gil88] William C. Gilmore. Hot pursuit and constructive presence in Canadian law enforcement: a case note. Marine Policy, 12(2):105–111, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900383.

Gilmore:1989:NIS

[Gil89] William C. Gilmore. Narcotics interdiction at sea: UK-US cooperation. Marine Policy, 13(3):218–230, July 1989. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900560.

Ginsburg:1986:LEO

[Gin86] Norton Ginsburg. Letter to the Editor: Ocean Yearbook 4. Marine Policy, 10(1):74–75, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900394.

Glassner:1986:BRM

[Gla86] Martin Ira Glassner. Book review: Managing the ocean: Resources, research, law: edited by Jacques G. Richardson, Lomond Publications, Mt Airy, Maryland, 1985, xv + 407 pp. Marine Policy, 10(4):315–316, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900072.

Glassner:1989:BRAb

[Gla89a] Martin Ira Glassner. Book review: Antarctica and South American geopolitics: Frozen Lebensraum: by Jack Child, Praeger, Westport, 1988, 232 pp, US$35.95. Marine Policy, 13(4):361–363, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900304.

Glassner:1989:BRI

[Gla89b] Martin Ira Glassner. Book review: International law for Antarctica: edited by Francesco Francioni and Tullio Scovazzi, Guiffré Editore, Milan, 1987, 532 pp, US$40.00. Marine Policy, 13(4):361, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900286.

Glassner:1989:BRM

[Gla89c] Martin Ira Glassner. Book review: Marine affairs bibliography 1980–1985: Christian L. Wiktor and Leslie A. Foster, Martnus Nijhoff, Dordrecht, 1987, xv + 685pp, $189.00. Marine Policy, 13 (3):271–272, July 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900651.

Glassner:1989:BRAa

[Gla89d] Martin Ira Glassner. Book review: The Antarctic legal regime: edited by Christopher C. Joyner and Sudhir K. Chopra, Martinus Nijhoff Publishers, Dordrecht, 1988, 288 pp, US$85.50. Marine Policy, 13(4):361, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900298.

Glassner:1989:BRLb

[Gla89e] Martin Ira Glassner. Book review: The law and the sea: Essays in memory of Jean Carroz: Food and Agriculture Organization of the United Nations, Rome, 1987, 282 pp. Marine Policy, 13(3):269–270, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900638.

Glassner:1989:BRLa

[Gla89f] Martin Ira Glassner. Book review: The law of the sea and maritime boundary delimitation in South-East Asia: Kriangsak Kittichaisaree, Oxford University Press, Singapore, 1987, xv + 209 pp. Marine Policy, 13(3):269, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900626.

Glassner:1989:BRLc

[Gla89g] Martin Ira Glassner. Book review: The law of the sea: What lies ahead? Proceedings of the twentieth annual conference of the Law of the Sea Institute, 1986: edited by Thomas A. Clingan Jr, The Law of the Sea Institute, University of Hawaii, Honolulu, 1988, 583 pp. Marine Policy, 13(4):355–357, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900237.

Glassner:1989:BRC

[Gla89h] Martin Ira Glassner. Book review: The UN convention on the law of the sea: Impact and implementation. Proceedings of the nineteenth annual conference of the Law of the Sea Institute, 1985: edited by E. D. Brown and R. R. Churchill, The Law of the Sea Institute, University of Hawaii, Honolulu, 1987, 639 pp, $38.50. Marine Policy, 13(4):355, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900225.

Grigalunas:1989:MCM

[GO89] Thomas A. Grigalunas and James J. Opaluch. Managing contaminated marine sediments: Economic considerations. Marine Policy, 13(4):318–333, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900171.

Godman:1980:SFT

[God80] N. A. Godman. Sea Fisheries Training Council — a fisheries quango. Marine Policy, 4(2):147–148, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900421.

Gold:1986:LAT

[Gol86] Edgar Gold. Legal aspects of the transportation of dangerous goods at sea. Marine Policy, 10(3):185–191, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900515.

Gold:1987:BRO

[Gol87a] Edgar Gold. Book review: Oil pollution from ships: International, United Kingdom and United States Law and Practice, 2 ed: by David W. Abecassis and Richard L. Jarashow, Stevens, London, 1985, 619 pp. £75.00. Marine Policy, 11(2):153–154, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900121.

Gold:1987:BRC

[Gol87b] Edgar Gold. Book review: The carriage of dangerous goods by sea: by Cleopatra Elmira Henry, Frances Pinter, London, 1985, 181 pp, £25.00. Marine Policy, 11(3):246, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900650.

Gold:1989:IML

[Gol89] Edgar Gold. International maritime law in transition: New challenges for education and training. Marine Policy, 13(3):178–192, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900547.

Goncalves:1980:BRR

[Gon80] Maria Eduarda Gonçalves. Book review: Regional cooperation in marine science: by Lewis M. Alexander, University of Rhode Island, Report prepared for the Intergovernmental Oceanographic Commission (IOC), the United Nations Ocean Economics and Technology Office (OETO) and the Food and Agriculture Organization (FAO), December 1978. Marine Policy, 4(2):152–154, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900469.

Goncalves:1983:UIU

[Gon83] Maria Eduarda Gonçalves. UN University/IOC/UNESCO Workshop on international cooperation in the development of marine science and the transfer of marine technology in the context of the New Ocean Regime: Paris, 27 September–1 October 1982. Marine Policy, 7(3):229–231, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900131.

Goodlad:1980:SFC

[Goo80] John H. Goodlad. Shetland fisheries — conservation and development. Marine Policy, 4(3):244–245, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900615.

Graham:1988:AML

[Gra88] Alastair Graham. Arabian maritime law. Marine Policy, 12(4): 411, October 1988. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X88900279.

Green:1980:DOM

[Gre80] David Green. Deep Ocean Mining — Opportunities for the UK. Seminar sponsored by the UK Science Research Council (SRC): London, UK, 7 February 1980. Marine Policy, 4(3):257–259, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900676.

Grime:1988:ILC

[Gri88] R. P. Grime. Implementation of the 1976 limitation convention: Liability for maritime claims. Marine Policy, 12(3):306–313, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8890070X.

Godman:1980:CET

[GS80] N. A. Godman and J. D. Stewart. Casual employment and trade unionism in the UK fishing industry. Marine Policy, 4(3):237–241, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900585.

Gulland:1980:BWU

[Gul80a] J. A. Gulland. Blue whale unit — need for flexibility. Marine Policy, 4(3):262, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900706.

Gulland:1980:PAO

[Gul80b] J. A. Gulland. Peruvian anchoveta — optimal management. Marine Policy, 4(1):78–79, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900329.

Gulland:1983:ECR

[Gul83] John Gulland. Expert Consultation on the Regulation of Fishing Effort: Rome, 17–26 January 1983. Marine Policy, 7(3):231, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900143.

Gulland:1984:FLB

[Gul84a] J. A. Gulland. Fisheries: Looking beyond the golden age. Marine Policy, 8(2):137–150, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900885.

Gulland:1984:HMW

[Gul84b] John Gulland. How much of the world’s fisheries lies within EEZs? Marine Policy, 8(4):359, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900319.

Gulland:1985:BRF

[Gul85] John Gulland. Book review: Fish catching methods of the world (third edition): Von Brandt Fishing News Books Ltd, Farnham, UK. Marine Policy, 9(3):244–245, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900259.

Gulland:1986:BRM

[Gul86] John A. Gulland. Book review: Marine mammals and fisheries: edited by J. R. Beddington, R. J. H. Beverton and D. M. Lavigne, George Men and Unwin, London 1985, 354pp, £40. Marine Policy, 10(3):234–236, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8690059X.

Gulland:1987:ISF

[Gul87a] J. A. Gulland. The impact of seals on fisheries. Marine Policy, 11(3):196–204, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8790056X.

Gulland:1987:MNS

[Gul87b] J. A. Gulland. The management of North Sea fisheries: Looking towards the 21st century. Marine Policy, 11(4):259–272, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900170.

Gulland:1987:SP

[Gul87c] John Gulland. Seals and the public. Marine Policy, 11(2):147–148, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8790008X.

Gulland:1987:BRE

[Gul87d] John A. Gulland. Book review: The economics of fisheries management, (revised edition): L. G. Anderson, Johns Hopkins University Press, Baltimore, MD, 1986, 296 pp, £24.40. Marine Policy, 11(1):80–81, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900455.

Gulland:1987:FF

[Gul87e] John A. Gulland. Falklands fishing. Marine Policy, 11(3):240, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900601.

Gulland:1988:BRW

[Gul88a] J. A. Gulland. Book review: The whale war: by David Day Routledge and Kegan Paul, London, 168 pp, £10.95. Marine Policy, 12(1):77–78, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900152.

Gulland:1988:TII

[Gul88b] John Gulland. Tuna and international institutions. Marine Policy, 12(4):408, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900243.

Gulland:1989:FAM

[Gul89a] John Gulland. The fourth annual meeting of the International Institute of Fisheries Economics and Trade, Esbjerg: Denmark, 7–12 August 1988. Marine Policy, 13(1):72–73, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900432.

Gulland:1989:SIB

[Gul89b] John Gulland. The second international billfish symposium: Kailea-Kona, Hawaii, 1–5 August 1988. Marine Policy, 13(1):73–74, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900444.

Hafner:1981:LLV

[Haf81] Gerhard Hafner. The ‘land-locked’ viewpoint. Marine Policy, 5 (3):281–282, July 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900622.

Hage:1984:CLS

[Hag84] Robert Hage. Canada and the law of the sea. Marine Policy, 8(1):2–15, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900538.

Hamzah:1984:IAR

[Ham84] B. A. Hamzah. Indonesia’s archipelagic regime: Implications for Malaysia. Marine Policy, 8(1):30–43, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900551.

Harders:1987:QAL

[Har87] J. Enno Harders. In quest of an Arctic legal regime: Marine regionalism — a concept of international law evaluated. Marine Policy, 11(4):285–298, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900194.

Hatlem:1988:BRN

[Hat88] O. B. Hatlem. Book review: Navigational restrictions within the new LoS context. Geographical implications for the United States: by Lewis M. Alexander offshore consultants, Peace Dale, RI, USA, 1987, 398 pp, $60.00. Marine Policy, 12(1):74–75, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900139.

Haward:1989:AOC

[Haw89] Marcus Haward. The Australian offshore constitutional settlement. Marine Policy, 13(4):334–348, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900183.

Hayward:1984:EPR

[Hay84] Peter Hayward. Environmental protection: Regional approaches. Marine Policy, 8(2):106–119, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900861.

Hirvonen:1986:CSP

[HC86] H. Hirvonen and R. P. Côté. Control strategies for the protection of the marine environment. Marine Policy, 10(1):19–28, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900291.

Hasselback:1981:CFP

[HDG81] Nancy L. Hasselback, Joel B. Dirlam, and John M. Gates. Canadian fisheries policy: Canadian lobster imports and the New England lobster industry. Marine Policy, 5(1):40–51, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900725.

Heen:1988:NFL

[Hee88] Knut Heen. Norwegian fishermen: Labour market behaviour and analysis. Marine Policy, 12(4):396–407, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900231.

Heggie:1981:BRP

[Heg81] Ian Heggie. Book review: Port pricing and investment policy for developing countries: by Esra Bennathan and A. A. Walters, Oxford University Press for the World Bank, Oxford and New York, 1979, 230pp, £5.70 hardback, £2.50 paperback. Marine Policy, 5(2):164–166, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900154.

Hilborn:1989:IMT

[Hil89] R. Hilborn. International management of tuna. Marine Policy, 13(2):166, April 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900067.

Hobbs:1983:TS

[Hob83] David Hobbs. Tomahawks at sea. Marine Policy, 7(1):53–55, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8390060X.

Holt:1982:WC

[Hol82] Sidney J. Holt. Whale conservation. Marine Policy, 6(4):350, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900173.

Holt:1985:WMW

[Hol85] Sidney Holt. Whale mining, whale saving. Marine Policy, 9(3): 192–213, July 1985. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X85900193.

Holser:1988:OLU

[Hol88a] Alexander F. Holser. Offshore lands of the USA: the US exclusive economic zone, continental shelf and outer continental shelf. Marine Policy, 12(1):2–8, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900024.

Holt:1988:BRI

[Hol88b] Sidney J. Holt. Book review: International regulation of whaling: From conservation of whaling to conservation of whales and regulation of whale-watching: by Patricia Birnie, Oceana Publications, New York, 1053 pp in two volumes, 1985, $50.00 per volume. Marine Policy, 12(4):416–417, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900322.

Holt:1988:WW

[Hol88c] Sidney J. Holt. The whale war. Marine Policy, 12(4):410–411, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900267.

Hoydal:1985:BRM

[Hoy85] K. Hoydal. Book review: Marine ecology: a comprehensive integrated treatise on life in oceans and coastal waters Vol 5. Part 2. Ocean Management: edited by Otto Kinne John Wiley, Chichester, UK, 1984, 576 pp, £47.50. Marine Policy, 9(2):161–163, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900119.

Park:1981:BRO

[hP81] Choon ho Park. Book review: The offshore petroleum resources of South-east Asia: Potential conflict situations and related economic considerations: Corazón Morales Siddayao, Oxford University Press, Kuala Lumpur, Malaysia, 1978, xx + 205 pp. £14.95. Marine Policy, 5(1):85–86, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900828.

Hilborn:1988:AMD

[HS88a] Ray Hilborn and John Sibert. Adaptive management of developing fisheries. Marine Policy, 12(2):112–121, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900395.

Hilborn:1988:IMT

[HS88b] Ray Hilborn and John Sibert. Is international management of tuna necessary? Marine Policy, 12(1):31–39, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8890005X.

Hu:1987:SAS

[Hu87] Nien-Tsu Alfred Hu. The Sino-Argentine ‘Squid War’ of 1986: Its implications for fisheries policy making. Marine Policy, 11(2): 133–142, April 1987. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X87900054.

Hu:1988:PNT

[Hu88] Nien-Tsu Alfred Hu. Policy note: Twenty or twenty-four miles? Marine Policy, 12(2):160–161, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900437.

Humphrey:1981:UMS

[Hum81] G. F. Humphrey. UNCLOS and the marine scientist. Marine Policy, 5(3):270–271, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900579.

Humphrey:1983:BRI

[Hum83a] G. F. Humphrey. Book review: Issues in Australia’s marine and Antarctic policies: edited by R. A. Herr, R. Hall and B. W. Davis, Public Policy Monograph, Department of Political Science, University of Tasmania, 1982, 201 pp. Marine Policy, 7(4):319, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900489.

Humphrey:1983:BRM

[Hum83b] George Humphrey. Book review: Marine scientific research and the law of the sea: by Alfred H. A. Soons, Kluwer Law and Taxation Publishers, Deventer, The Netherlands, 1982, 383 pp. Marine Policy, 7(2):129–130, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900283.

Humphrey:1987:BRA

[Hum87a] G. F. Humphrey. Book review: Australia’s offshore maritime interests: edited by W. S. G. Bateman and M. W. Ward, Occasional Papers in Maritime Affairs: 3, Australian Centre for Maritime Studies, Canberra, 1985, 122 pp. Marine Policy, 11(3):246–247, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900662.

Humphrey:1987:MAP

[Hum87b] G. F. Humphrey. Maritime Australia 86 — ‘putting it together’: Canberra, Australia, 14–15 March 1986. Marine Policy, 11(1): 75–76, January 1987. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8790042X.

Humphrey:1987:SAL

[Hum87c] G. F. Humphrey. Seminar on Australia and the law of the sea: Sydney, Australia, 28 February 1987. Marine Policy, 11(3):247–249, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900674.

Jamaguchi:1982:DJ

[Jam82] Kaiji Jamaguchi. The defence of Japan. Marine Policy, 6(2): 151–154, April 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900720.

Jentoft:1985:MFD

[Jen85] Svein Jentoft. Models of fishery development: the cooperative approach. Marine Policy, 9(4):322–331, October 1985. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900405.

Jentoft:1989:FCM

[Jen89] Svein Jentoft. Fisheries co-management: Delegating government responsibility to fishermen’s organizations. Marine Policy, 13(2): 137–154, April 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900043.

Jin:1989:QIP

[Jin89] Shao Jin. The question of innocent passage of warships: After UN-CLOS III. Marine Policy, 13(1):56–67, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900419.

Jlstra:1989:BRP

[Jls89a] Ton I. Jlstra. Book review: Poisoners of the seas: by K. A. Gourlay, Zed Books Ltd, 1988, 256 pp, £7.95. Marine Policy, 13(2):171–172, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900092.

Jlstra:1989:BRW

[Jls89b] Ton I. Jlstra. Book review: The waters around the British Isles; their conflicting uses: by R. B. Clark, Clarendon Press, Oxford, 1987, 386 pp, £32.50. Marine Policy, 13(1):78–80, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8990047X.

Jentoft:1987:GSN

[JM87] Svein Jentoft and Knut H. Mikalsen. Government subsidies in Norwegian fisheries: Regional development or political favouritism? Marine Policy, 11(3):217–228, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900583.

Johnston:1984:IIF

[Joh84] Richard S. Johnston. The international institute of fisheries economics and trade. Marine Policy, 8(3):277, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900083.

Joseph:1980:EPT

[Jos80] James Joseph. Eastern Pacific tuna fisheries — how to manage? Marine Policy, 4(3):259–261, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900688.

Kaczynski:1984:JVE

[Kac84] Vladimir M. Kaczynski. Joint ventures as an export market: US groundfish. Marine Policy, 8(1):16–29, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8490054X.

Kaczynski:1989:FFF

[Kac89] Vlad M. Kaczynski. Foreign fishing fleets in the subSaharan West African EEZ: the coastal state perspective. Marine Policy, 13 (1):2–15, January 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900377.

Kolodkin:1988:LIP

[KAK88] A. L. Kolodkin, V. V. Andrianov, and V. A. Kiselev. Legal implications of participation or non-participation in the 1982 convention. Marine Policy, 12(3):187–191, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900565.

Kasoulides:1987:VDS

[Kas87] George C. Kasoulides. Vessels in distress: ‘safe havens’ for crippled tankers. Marine Policy, 11(3):184–195, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900558.

Kasoulides:1989:ROP

[Kas89] George C. Kasoulides. Removal of offshore platforms and the development of international standards. Marine Policy, 13(3):249–265, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900584.

Keen:1983:CPF

[Kee83] E. A. Keen. Common property in fisheries: Is sole ownership an option? Marine Policy, 7(3):197–211, July 1983. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900040.

Kolodkin:1988:ISL

[KG88] A. L. Kolodkin and E. B. Ganiushkina. INMARSAT: some legal aspects. Marine Policy, 12(3):256–262, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900644.

Kildow:1983:ATP

[Kil83] Judith Kildow. Arctic technology and policy: an assessment and review for the next decade: Massachusetts Institute of Technology, Cambridge, MA, USA, 2–4 March 1983. Marine Policy, 7 (3):232–234, July 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900155.

Kimball:1980:TCL

[Kim80a] Lee Kimball. Third UN conference on the law of the sea, ninth session: New York, USA, 3 March–4 April 1980. Marine Policy, 4(3):246–249, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900627.

Kimball:1980:UAC

[Kim80b] Lee Kimball. UNCLOS — around the corner. Marine Policy, 4 (1):69–72, January 1980. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X80900287.

Kimball:1981:MGT

[Kim81a] Lee Kimball. A microcosm of global trends. Marine Policy, 5 (3):271–273, July 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900580.

Kimball:1981:UDC

[Kim81b] Lee Kimball. UNCLOS — the draft convention. Marine Policy, 5(1):76–79, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900798.

Kimball:1981:UTS

[Kim81c] Lee Kimball. Update — the tenth session. Marine Policy, 5 (3):287–290, July 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900658.

Kimball:1982:TCLb

[Kim82a] Lee Kimball. Third UN conference on the law of the sea, eleventh session, New York, 8 March–30 April 1982. Marine Policy, 6(4): 342–344, October 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900148.

Kimball:1982:TCLa

[Kim82b] Lee Kimball. Third UN Conference on the Law of the Sea, Tenth Session (resumed): Geneva, Switzerland, 3–28 August 1981. Marine Policy, 6(1):69–71, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900483.

Kimball:1984:PCL

[Kim84] Lee Kimball. Preparatory Commission for the Law of the Sea Convention: Kingston, Jamaica, 19 March–13 April 1984. Marine Policy, 8(4):363–366, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900332.

Kimball:1985:PCLa

[Kim85a] Lee Kimball. Preparatory Commission for the Law of the Sea Convention: Geneva, Switzerland, 13 August–5 September 1984. Marine Policy, 9(1):73–77, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8590082X.

Kimball:1985:PCLb

[Kim85b] Lee Kimball. Preparatory Commission for the Law of the Sea Convention: Kingston, Jamaica, 11 March–4 April 1985. Marine Policy, 9(4):340–343, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900478.

Kimball:1986:BRC

[Kim86a] Lee Kimball. Book review: Consensus and confrontation: the United States and the law of the sea convention: edited by Jon M. Van Dyke, Law of the Sea Institute, Honolulu, 1985, 576pp, $29.50. Marine Policy, 10(1):77–78, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900412.

Kimball:1986:PCI

[Kim86b] Lee Kimball. Preparatory Commission for the International Seabed Authority (ISA)/International Tribunal on the Law of the Sea (IT-LOS): Geneva, Switzerland, 12 August–4 September 1985. Marine Policy, 10(1):60–62, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900345.

Kimball:1987:PCL

[Kim87] Lee Kimball. Preparatory Commission for the 1982 Law of the Sea (LOS) convention, fourth session: Kingston, Jamaica, 11 March–17 April, 1986, and New York, 12 August–5 September 1986. Marine Policy, 11(1):69–72, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900406.

Kimball:1988:BRI

[Kim88a] Lee Kimball. Book review: International organizations and the law of the sea documentary yearbook 1985: by The Netherlands Institute for the Law of the Sea, Martinus Nijhoff, Dordrecht, 1987, 645 pp, £00.00. Marine Policy, 12(4):415–416, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900310.

Kimball:1988:FSP

[Kim88b] Lee A. Kimball. The Fifth Session of the Preparatory Commission: Kingston, Jamaica, 30 March–6 April 1987 and New York, USA, 27 July-21 August 1987. Marine Policy, 12(1):67–70, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900103.

Kimball:1989:SSP

[Kim89] Lee Kimball. The Sixth Session of the Preparatory Commission: Kingston, Jamaica, 14 March–8 April 1988 and New York, USA, 15 August–2 September 1988. Marine Policy, 13(2):168–171, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900080.

King:1980:DKR

[Kin80a] Dennis M. King. Dr. King replies. Marine Policy, 4(3):261–262, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8090069X.

King:1980:EPE

[Kin80b] Lauriston R. King. Expanding perspectives — evolution of comparative policy. Marine Policy, 4(1):75–77, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900305.

Kindt:1983:FNP

[Kin83] John Warren Kindt. Floating nuclear power plants: US and international regulations. Marine Policy, 7(2):90–100, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900209.

Kirthisingha:1981:EET

[Kir81] P. N. Kirthisingha. The Enterprise: an expendable triumph. Marine Policy, 5(3):252–256, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8190052X.

Kiselev:1988:SAP

[Kis88] V. A. Kiselev. ‘Special areas’ for preventing pollution of the sea. Marine Policy, 12(3):241–246, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900620.

Kbaier:1985:NTC

[KS85] Rouchdy Kbaier and Viktor Sebek. New trends in compensation for oil pollution damage: Amoco Cadiz legal proceedings and the 1984 diplomatic conference on liability and compensation. Marine Policy, 9(4):269–279, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900363.

Kumar:1982:CVD

[Kum82] Chandra K. Kumar. Commonwealth view of deep sea mining. Marine Policy, 6(3):239–241, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900276.

Kwiatkowska:1984:BRN

[Kwi84] Barbara Kwiatkowska. Book review: Novo pravo mora (the new law of the sea): Edited by B. Vukas, Prinosi za predbeno proucavanje prava i medunarodno pravo (Contribution to the Study of Comparative and International Law) No 17, Institute of International Law and International Relations, University of Zagreb, Yugoslavia, 1982, 308 pp. Marine Policy, 8(1):71, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900629.

Kwiatkowska:1985:BRI

[Kwi85] Barbara Kwiatkowska. Book review: The International Law of the Sea, issues of implementation in Indonesia Proceedings of a seminar at Jakarta, 22–27 August 1983: edited by E. Hey and A. W. Koers, Netherlands Institute of Transport, Rijswijk, The Netherlands, 1984, 181 pp. Marine Policy, 9(2):160–161, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900107.

Kwiatkowska:1987:BRE

[Kwi87] Barbara Kwiatkowska. Book review: Exclusive economic zone claims: an analysis and primary documents: by Robert W. Smith, Martinus Nijhoff Publishers, Dordrecht-Boston-Lancaster, 1986, 501 pp, £112.25. Marine Policy, 11(4):321–323, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8790025X.

Kwiatkowska:1988:PLS

[Kwi88a] Barbara Kwiatkowska. 1st Penataran on The Law of The Sea and International Ecomomic Law: Bandung, Indonesia, 13 July–1 August 1987. Marine Policy, 12(1):73–74, January 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900127.

Kwiatkowska:1988:IOM

[Kwi88b] Barbara Kwiatkowska. The Indian Ocean Marine Affairs Cooperation Conference: Colombo, Sri Lanka, First Conference: 20–28 January 1987; Standing Committee: 7–9 September 1987. Marine Policy, 12(2):170–172, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900498.

Kwiatkowska:1989:BRN

[Kwi89a] Barbara Kwiatkowska. Book review: A new law of the sea for the Caribbean: an examination of marine law and policy issues in the Lesser Antilles. Lecture notes on coastal and estuarine studies no 27: edited by Edgar Gold, Springer-Verlag, Berlin, 1988, 276 pp. Marine Policy, 13(4):353–355, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900213.

Kwiatkowska:1989:BRS

[Kwi89b] Barbara Kwiatkowska. Book review: The system for settlement of disputes under the United Nations convention on the law of the sea: by Andronico O. Adede, Publications on Ocean Development, General Editor Shigeru Oda, Volume 10, Martinus Nijhoff, Dordrecht, 1987, 302 pp, £45.50. Marine Policy, 13(1):81–82, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900511.

Lamson:1987:ASM

[Lam87] Cynthia Lamson. Arctic shipping, marine safety and environmental protection. Marine Policy, 11(1):3–15, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900352.

Langevad:1981:PPD

[Lan81] E. J. Langevad. Production policy for the deep sea mineral resources. Marine Policy, 5(3):264–267, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900555.

Langeraar:1986:MDE

[Lan86] Wijnand Langeraar. Maritime delimitation: the equiratio method — a new approach. Marine Policy, 10(1):3–18, January 1986.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8690028X.

Larkins:1980:MUF

[Lar80] H. A. Larkins. Management under FCMA: Development of a fishery management plan. Marine Policy, 4(3):170–182, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900524.

Lavety:1987:VLD

[Lav87] Jack Lavety. Variability in the lengths of Dover sole (Solea solea) after death. Marine Policy, 11(4):319–320, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900236.

Lawson:1980:PEF

[Law80] Rowena M. Lawson. Post-evaluation of fisheries projects. Marine Policy, 4(1):42–51, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900238.

Lawson:1982:SSF

[Law82] Rowena M. Lawson. Small-scale fisheries in the CECAF region — priorities for research. Marine Policy, 6(1):66–68, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8290046X.

Lawson:1983:BRM

[Law83] Rowena Lawson. Book review: Modernization and marine fisheries policy: edited by John R. Maiolo and Michael K. Orbach, Ann Arbor Science, Ann Arbor, MI, 1982, 330 pp, £25. Marine Policy, 7(2):131–132, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900301.

Lawson:1984:FNV

[Law84] Rowena Lawson. A future for Namibia: Volume 5: Fishing: Richard Moorsom, Catholic Institute for International Relations, 22 Coleman Fields, London N1 7AF, 1984, 123 pp, £2.95. Marine Policy, 8(4):368–369, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900356.

Lawson:1985:BRF

[Law85] Rowena Lawson. Book review: Fisheries economics: an introduction: by Stephen Cunningham, Michael R. Dunn and David Whitmarsh Mansell Publishing Ltd, London and St Martin’s Press, New York, 1985, pp 372, £19.85. Marine Policy, 9(4):345–346, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900508.

Lawson:1986:BRF

[Law86] Rowena Lawson. Book review: Fishery development experiences: by W. H. L. Allsopp, Fishing News Books, Farnham, Surrey, 160 pp, £12.50. Marine Policy, 10(3):233–234, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900588.

Lawson:1988:AFM

[Law88] Rowena Lawson. Artisanal fisheries management. Marine Policy, 12(2):165, April 1988. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X88900450.

Lynne:1982:CR

[LC82] Lynne and Brian Chatterton. The Chattertons reply. Marine Policy, 6(2):147–148, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900690.

Leigh:1985:CFP

[Lei85] Michael Leigh. The common fisheries policy of the European community: Mark Wise, Methuen, London, 1984. Marine Policy, 9(3):245–246, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900260.

Lennard:1980:BRJ

[Len80] Don Lennard. Book review: Jane’s ocean technology 1979–80: edited by Robert L. Trillo, Macdonald and Jane’s, London, 1979, 896 pp. £32.00. Marine Policy, 4(4):334–335, October 1980. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900135.

Lennard:1987:BRF

[Len87] Don Lennard. Book review: The future for ocean technology: by Glyn Ford, Chris Niblett and Lindsay Walker, Frances Pinter, London and New Hampshire, 1987, xii + 139 pp, £17.50. Marine Policy, 11(4):323–324, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900261.

Levy:1988:TIM

[Lev88] Jean-Pierre Levy. Towards an integrated marine policy in developing countries. Marine Policy, 12(4):326–342, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900188.

Liao:1982:BRL

[Lia82] David S. Liao. Book review: Limited entry as a fishery management tool: edited by R. Bruce Rettig and Jay J. C. Ginter, Washington Sea Grant Program, distributed by University of Washington Press, Seattle and London, 1980, 456 pp, $15.00. Marine Policy, 6(1):84–85, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900598.

Lieberman:1986:TIF

[Lie86] Warren H. Lieberman. Towards improving fishery management systems. Marine Policy, 10(1):42–50, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8690031X.

Lockwood:1984:BRF

[Loc84] Stephen J. Lockwood. Book review: Fish stock assessment: a manual of basic methods: J. A. Gulland, FAO/Wiley Series on Food and Agriculture, Vol 1, Wiley-lnterscience, Chichester, UK, 1983, 223 pp, £15.95. Marine Policy, 8(1):68–69, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900605.

Loftas:1981:FEP

[Lof81] Tony Loftas. FAO’s EEZ programme: Assisting a new era in fisheries. Marine Policy, 5(3):229–239, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900506.

Longhurst:1981:OPO

[Lon81] A. R. Longhurst. 1980s’ oceans policy — offshore environment: Offshore environment in the ‘80s’, a workshop on environmental considerations of east coast offshore hydrocarbon development, St John’s, Newfoundland, Canada, 2–4 December 1980. Marine Policy, 5(2):161, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900117.

Lowe:1980:UNA

[Low80] A. V. Lowe. UNCLOS and navies: Accord at a price. Marine Policy, 4(4):270, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900019.

Lowe:1981:BRL

[Low81a] A. V. Lowe. Book review: The lawfulness of deep seabed mining: Theodore G. Kronmiller, 2 vols. Oceana, Dobbs Ferry, New York. 1980, xix + 527 pp, and iii + 460 pp, $40.00 per volume. Marine Policy, 5(1):90–91, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900877.

Lowe:1981:FDS

[Low81b] A. V. Lowe. First deep seabed mining workshop: Paris, 8–10 December 1980. Marine Policy, 5(4):333–334, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81901627.

Lowe:1981:ISL

[Low81c] A. V. Lowe. The international seabed: a legacy of mistrust. Marine Policy, 5(3):205–216, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900488.

Lowe:1982:BRR

[Low82a] A. V. Lowe. Book review: Restricting the concept of the freedom of the seas: Modern maritime law re-evalued, by George P. Smith II, Robert E. Krieger, New York, NY, 1980, 242 pp. Marine Policy, 6(1):79–80, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900549.

Lowe:1982:BRE

[Low82b] A. V. Lowe. Book review: The environmental law of the sea: by Douglas M. Johnston and others Erich Schmidt Verlag, Berlin, 1981, 419 pp. Marine Policy, 6(4):337–338, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900112.

Lowe:1982:MAS

[Low82c] A. V. Lowe. A move against substandard shipping. Marine Policy, 6(4):326–330, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900069.

Lowe:1984:BRM

[Low84] A. V. Lowe. Book review: Markets under the sea: by D. R. Denman, Hobart Paperback 17, Institute of Economic Affairs, 1984, xvi and 80 pp. Marine Policy, 8(3):286–287, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900137.

Lowe:1985:BRI

[Low85] A. V. Lowe. Book review: The international law of the sea: Volume II: the late D. P. O’Connell, edited by I. A. Shearer, Clarendon Press, Oxford, 1984, xxvi and 524 pp, plus 43 pp, Hardback £40. Marine Policy, 9(1):79–80, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900855.

Lowe:1986:SLP

[Low86] A. V. Lowe. Some legal problems arising from the use of the seas for military purposes. Marine Policy, 10(3):171–184, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900503.

Lowe:1987:R

[Low87] A. V. Lowe. Rejoinder. Marine Policy, 11(3):250–252, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900698.

Lowe:1988:LWS

[Low88] A. V. Lowe. The laws of war at sea and the 1958 and 1982 conventions. Marine Policy, 12(3):286–296, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900681.

Lee:1987:ECF

[LP87] M. Lee and R. Pearson. Expansion of China’s fleet. Marine Policy, 11(4):315–319, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900224.

Lawson:1983:AFW

[LR83] Rowena Lawson and Michael Robinson. Artisanal fisheries in West Africa: Problems of management implementation. Marine Policy, 7(4):279–290, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900416.

Larson:1989:LSM

[LR89] David L. Larson and Michael W. Roth. The law of the sea and maritime strategy. Marine Policy, 13(3):193–217, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900559.

Lepiz:1985:SEO

[LS85] Luis Guillermo Lepiz and Jon G. Sutinen. Surveillance and enforcement operations in the Costa Rican tuna fishery. Marine Policy, 9(4):310–321, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900399.

Larson:1982:LSA

[LT82] David Larson and Peter Tarpgaard. Law of the sea and ASW: National security versus arms control. Marine Policy, 6(2):90–102, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900641.

Luna:1980:FDL

[Lun80] Julio Luna. Fishery development — the Latin American model. Marine Policy, 4(2):148–150, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900433.

Lundberg:1984:OII

[Lun84] Olof Lundberg. Organizations: INMARSAT: improving maritime communications. Marine Policy, 8(4):337–346, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900265.

Loannides:1987:PFF

[LW87] Chris Loannides and David Whitmarsh. Price formation in fisheries. Marine Policy, 11(2):143–145, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900066.

Madeley:1985:FWF

[Mad85a] John Madeley. FAO world fisheries conference — a reply. Marine Policy, 9(4):338, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900454.

Madeley:1985:FMD

[Mad85b] John Madeley. ‘Fisheries Management and Development’, UN Food and Agriculture Organisation (FAO) Conference: Rome, 27 June–6 July 1984. Marine Policy, 9(1):77–78, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900831.

Mankabady:1980:OFI

[Man80] S. Mankabady. OCEANEXPO 80, fourth International exhibition and symposium on the exploitation of the oceans: Bordeaux, France, 4–8 March 1980. Marine Policy, 4(4):330–332, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8090010X.

Mankabady:1981:BRN

[Man81] Samir Mankabady. Book review: Natural gas by sea: the development of a new technology: by Roger Ffooks, Gentry Books, London, UK, 1979, 234 pp, £18.50. Marine Policy, 5(4):346–347, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901706.

Marjoram:1982:ICR

[Mar82a] Tony Marjoram. International conference on resource development in the Pacific: Honolulu, HI, USA, 13–15 October 1981. Marine Policy, 6(3):248–251, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900318.

Marjoram:1982:LSO

[Mar82b] Tony Marjoram. ‘The Law of the sea and ocean development issues in the Pacific basin’: 15th Annual Conference of the Law of the Sea Institute Honolulu, HI, USA, 5–8 October 1982. Marine Policy, 6(3):245–248, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900306.

Marti:1987:CST

[Mar87] Bruce E. Marti. Cargo-sharing through a policy of bilateralism: Effects on US containerized foreign trade. Marine Policy, 11(2): 97–104, April 1987. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X87900029.

Marriott:1988:FSN

[Mar88a] Alan Marriott. Fisheries statistics: the need for user focus. Marine Policy, 12(4):409, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900255.

Marston:1988:BRMa

[Mar88b] Geoffrey Marston. Book review: The maritime laws of the Arabian Gulf Cooperation Council States: Volume I, written and edited by Richard Price; Graham & Trotman, London, 1986: Volume I, 267 pp. £73.50. Marine Policy, 12(2):174–175, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900516.

Marston:1988:BRMb

[Mar88c] Geoffrey Marston. Book review: The maritime laws of the Arabian Gulf Cooperation Council States: Volume II, edited by Richard Price, Graham & Trotman, London, 1986: Vol II, 850 pp, £288.50. Marine Policy, 12(2):174–175, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900528.

Marston:1989:BRSa

[Mar89a] Geoffrey Marston. Book review: Sea-bed energy and mineral resources and the law of the sea volume 1: the areas within national jurisdiction: by E. D. Brown, Graham and Trotman, London, 1984– 1986, Volume 1 £36.50/US $64.00 (surface mail); £41.00/US $72.00 (airmail). Marine Policy, 13(1):80, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900481.

Marston:1989:BRSb

[Mar89b] Geoffrey Marston. Book review: Sea-bed energy and mineral resources and the law of the sea volume 2: the area beyond the limits of national jurisdiction: by E. D. Brown, (airmail); Volume 2 £46.50/US $82.00 (surface mail); £51.00/US $90.00 (airmail). Marine Policy, 13(1):80, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900493.

Marston:1989:BRSc

[Mar89c] Geoffrey Marston. Book review: Sea-bed energy and mineral resources and the law of the sea volume 3: Selected documents, tables and bibliography: by E. D. Brown, (airmail); Volume 3 £56.50/US $99.00 (surface mail); £61.00/US $107.00 (airmail). Marine Policy, 13(1):80–81, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8990050X.

Mathews:1986:SFM

[Mat86] C. P. Mathews. Shrimp and fisheries management workshop, Kuwait, 11–13 January 1986. Marine Policy, 10(3):231–232, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900564.

Marriott:1981:SLN

[MB81] Alan Marriott and T. R. Brainerd. Sierra Leone — new directions in fisheries statistics. Marine Policy, 5(1):72–73, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900762.

Mazany:1987:MSN

[MBA87] R. L. Mazany, L. G. Barrett, and R. A. Apostle. Market segmentation: Nova Scotia fish processing and the US market. Marine Policy, 11(1):29–44, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900376.

McBroy:1984:SIS

[McB84] J. K. McBroy. Second international shellfish conference: Dublin, Ireland, 19–21 April 1983. Marine Policy, 8(1):75–77, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900666.

MccGwire:1980:SFS

[Mcc80a] Michael MccGwire. Strategic forum: Soviet seapower — a new kind of navy. Marine Policy, 4(4):317–322, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900056.

McColl:1980:CFI

[McC80b] R. A. McColl. Consequences for the fishing industry of oil industry debris. Marine Policy, 4(1):61–63, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900251.

McConville:1980:BRG

[McC80c] J. McConville. Book review: Government and shipbuilding. The politics of industrial change: Brian W. Hogwood Saxon House, Farnborough, UK, 1979, 302 pp, sP10.50. Marine Policy, 4(3):264–265, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900731.

McDorman:1983:NLC

[McD83] Ted L. McDorman. National legislation and Convention obligations: Canadian vessel-source pollution law. Marine Policy, 7(4): 302–312, October 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8390043X.

McDorman:1985:TLS

[McD85] Ted L. McDorman. Thailand and the 1982 law of the sea convention. Marine Policy, 9(4):292–309, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900387.

McDorman:1986:WPS

[McD86] Ted L. McDorman. In the wake of the Polar Sea: Canadian jurisdiction and the northwest passage. Marine Policy, 10(4):243–257, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900011.

McElroy:1982:SMF

[McE82a] J. K. McElroy. Seminar on the marketing of fish: London, UK, November 1981. Marine Policy, 6(2):154–156, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900732.

McElroy:1982:SIC

[McE82b] J. K. McElroy. Shellfish international conference, Jersey, UK, 27–28 April 1982. Marine Policy, 6(4):347–350, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900161.

McElroy:1982:KSE

[McE82c] James McElroy. Krill — still an enigma. Marine Policy, 6(3): 238–239, July 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900264.

McElroy:1983:NP

[McE83a] J. K. McElroy. From the net to the plate. Marine Policy, 7(4): 314–317, October 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900453.

McElroy:1983:PF

[McE83b] J. K. McElroy. Packaging the Falklands. Marine Policy, 7(2): 124–125, April 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8390026X.

McElroy:1983:WPF

[McE83c] J. K. McElroy. A workable policy for Falkland Island fisheries. Marine Policy, 7(3):215–218, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900076.

McElroy:1984:AFH

[McE84] James K. McElroy. Antarctic fisheries: History and prospects. Marine Policy, 8(3):239–258, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900046.

McElroy:1989:ITF

[McE89] James Kevin McElroy. Indonesia’s tuna fisheries: Past, present and future prospects. Marine Policy, 13(4):285–308, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900158.

McMillan:1985:ECS

[McM85] Duncan J. McMillan. The extent of the continental shelf: Factors affecting the accuracy of a continental margin boundary. Marine Policy, 9(2):148–156, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900065.

McRae:1984:NOR

[McR84] D. M. McRae. The new oceans regime: Implementing the convention. Marine Policy, 8(2):83–94, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900848.

McRae:1986:BRN

[McR86] Donald M. McRae. Book review: Northwest passage: Arctic straits: by Donat Pharand, in association with Leonard H. Legault, Martinus Nijhoff Publishers, Dordrecht, The Netherlands, 1984, xxii + 199pp. Dfl 120. Marine Policy, 10(1):78–79, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900424.

Meese:1983:CD

[Mee83a] Sally Meese. The case for the defence. Marine Policy, 7(1): 56–57, January 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900623.

Meese:1983:OM

[Mee83b] Sally Meese. LOS and ‘other minerals’. Marine Policy, 7(2):123–124, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900258.

Mensah:1984:EPI

[Men84] Thomas A. Mensah. Environmental protection: International approaches. Marine Policy, 8(2):95–105, April 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8490085X.

Mendelson:1988:FLS

[Men88a] M. H. Mendelson. Fragmentation of the law of the sea. Marine Policy, 12(3):192–200, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900577.

Menefee:1988:MTE

[Men88b] Samuel Pyeatt Menefee. Maritime terror in Europe and the Mediterranean. Marine Policy, 12(2):143–152, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900413.

Meredith:1980:ISC

[Mer80] Brian Meredith. INMARSAT — satellite communications. Marine Policy, 4(4):329–330, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900093.

Meredith:1981:DPL

[Mer81a] Brian Meredith. Development problems of LDC ports. Marine Policy, 5(4):338–340, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81901664.

Meredith:1981:OOO

[Mer81b] Brian Meredith. OECD — the oracles’ oracle. Marine Policy, 5 (1):73–75, January 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900774.

Meredith:1981:UOI

[Mer81c] Brian Meredith. UNCTAD and the oceans — influence and organization. Marine Policy, 5(2):152–154, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900063.

Meredith:1982:RS

[Mer82] Brian Meredith. Rationalization at sea. Marine Policy, 6(4): 332–335, October 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900094.

Meredith:1983:BRC

[Mer83] Brian Meredith. Book review: Crowded agendas, crowded rooms: Institutional arrangements at UNCLOS III. Some lessons in global negotiations: by Mohamed El Baradi and Chleo Gavin, UNITAR Policy and Efficacy Studies, No 3, United Nations Institute for Training and Research, 801 UN Plaza, New York, NY 10017, October 1981. Marine Policy, 7(1):61, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900659.

Metaxas:1980:DME

[Met80] B. N. Metaxas. The development of maritime economics. Marine Policy, 4(3):242–244, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900603.

Metaxas:1981:FC

[Met81] B. N. Metaxas. Flags of convenience. Marine Policy, 5(1):52– 66, January 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900737.

Metaxas:1984:FNM

[Met84] B. N. Metaxas. Fourth national maritime economists conference: Newcastle, UK, 28–29 June 1983. Marine Policy, 8(1):80, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8490068X.

Meuriot:1986:FFR

[Meu86] Eric Meuriot. Fishing fleet replacement: the French policy from 1945 to 1983. Marine Policy, 10(4):294–309, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900059.

Miles:1983:IDI

[Mil83] Edward Miles. IOC data and information exchange: Implications of the Law of the Sea Convention. Marine Policy, 7(2):75–89, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900192.

Mitchell:1983:CIN

[Mit83] Barbara Mitchell. Conference on the Implications of the New Law of the Sea for the European Environment: Brussels, 27–29 January 1983. Marine Policy, 7(3):226–228, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8390012X.

Moore:1981:TDN

[Moo81] J. E. Moore. Technical developments in naval affairs. Marine Policy, 5(1):67–69, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900749.

Moore:1984:LCS

[Moo84] K. A. Moore. UN Liner code and shares of trade. Marine Policy, 8(3):274–276, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900071.

Moore:1986:LEE

[Moo86] Gerald Moore. L’accès des étrangers à l’exploitation des ressources biologiques de la zone économique exclusive: Antonio Tavares de Pinho, Thesis for the Doctorate in Law, university of Nice, December 1983, pp 574. (French) [Access by foreigners to the exploitation of the biological resources of the Exclusive Economic Zone]. Marine Policy, 10(2):160–161, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900229.

Moore:1987:BRI

[Moo87] Gerald K. Moore. Book review: The implementation of the new law of the sea in West Africa — prospects for the development and management of marine resources: Dalhousie Ocean Studies Programme, Dalhousie University, 1985, 132 pp. Marine Policy, 11(1):83–85, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900509.

Morgan:1980:PDM

[Mor80] G. R. Morgan. Population dynamics and management of the western rock lobster fishery. Marine Policy, 4(1):52–60, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8090024X.

Morgan:1981:BRP

[Mor81a] Joseph R. Morgan. Book review: The politics of Pacific island fisheries: by George Kent, Westview Press, Boulder, CO, 1980, 191 pp, $22. Marine Policy, 5(4):344–345, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81901688.

Morris:1981:LAN

[Mor81b] Michael A. Morris. Latin American navies in comparative perspective. Marine Policy, 5(2):149–152, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900051.

Morgan:1982:ANN

[Mor82] Joseph R. Morgan. ASEAN navies: new missions and old problems. Marine Policy, 6(3):236–237, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900252.

Morgan:1984:MRR

[Mor84] Joseph Morgan. Marine regions and regionalism in South-east Asia. Marine Policy, 8(4):299–310, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900228.

Morris:1988:BRL

[Mor88a] Michael A. Morris. Book review: L’Europe bleue: La politique communautaire des ressources marines: by Finn Laursen, Institute for Global Policy Studies, Amsterdam, The Netherlands, 1986, 158 pp. Marine Policy, 12(4):414, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900309.

Morris:1988:BRN

[Mor88b] Michael A. Morris. Book review: Northern waters: Security and resource issues: edited by Clive Archer and David Scrivener, The Royal Institute of International Affairs, Croom Helm, London, 1986, 240 pp £25.00. Marine Policy, 12(4):412–413, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900292.

Morgan:1989:BRT

[Mor89] Joseph R. Morgan. Book review: The Turkish straits: Christos L. Rozakis and Petros N. Stagos, Martinus Nijhoff Publishers, Dordrecht, 1987. Marine Policy, 13(2):173–174, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900110.

McDorman:1985:GMB

[MSV85] Ted L. McDorman, Phillip M. Saunders, and David L. VanderZwaag. The Gulf of Maine boundary: Dropping anchor or setting a course? Marine Policy, 9(2):90–107, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900016.

Marasco:1982:CIC

[MT82] Richard J. Marasco and Joseph M. Terry. Controlling incidental catch: an economic analysis of six management options. Marine Policy, 6(2):131–139, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900677.

McDorman:1987:IPT

[MT87] Ted L. McDorman and Panat Tasneeyanond. Increasing problems for Thailand’s fisheries: Malaysia’s new fisheries law. Marine Policy, 11(3):205–216, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900571.

Mukherjee:1981:CRO

[Muk81] P. K. Mukherjee. The consent regime of oceanic research in the new law of the sea. Marine Policy, 5(2):98–113, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900014.

Mullen:1989:MT

[Mul89] A. J. Mullen. Mobility of tuna. Marine Policy, 13(1):77–78, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900468.

Munro:1985:CSD

[Mun85] Gordon R. Munro. Coastal states, distant-water fleets and EFJ: Some long-run considerations. Marine Policy, 9(1):2–15, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900764.

Munro:1986:BRA

[Mun86] Gordon R. Munro. Book review: Atlantic fisheries and coastal communities: Fisheries decision-making case studies: edited by Cynthia Lamson and Arthur J. Hanson, Institute for resource and environmental studies, Dalhousie University, Halifax, NS, 1984, 262 pp, $18.50. Marine Policy, 10(2):159–160, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900217.

Munro:1988:BRN

[Mun88] Gordon R. Munro. Book review: Natural resources economics and policy applications: Essays in honor of James A. Crutchfield: edited by Edward Miles, Robert Pealy and Robert Stokes, University of Washington Press, Seattle, 1987, 456 pp, $30.00. Marine Policy, 12(4):417, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900334.

Nagelmackers:1980:AAC

[Nag80] Henri-G. Nagelmackers. Aftermath of the Amoco Cadiz: Why must the European community act? Marine Policy, 4(1):3–18, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900202.

Nailor:1980:BRS

[Nai80a] Peter Nailor. Book review: The seabed arms control negotiations: a study of Multilateral Arms Control Conference Diplomacy: B. Ramberg, University of Denver Monograph Series in World Affairs, Vol 15, Book 2, University of Denver, USA, 1978, 135 pp. Marine Policy, 4(3):263–264, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8090072X.

Nailor:1980:NDA

[Nai80b] Peter Nailor. Navies — the decisions ahead. Marine Policy, 4 (3):236–237, July 1980. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X80900573.

Nailor:1981:BRS

[Nai81] Peter Nailor. Book review: Sea power and influence: old issues and new challenges, edited by Jonathan Alford, published for the International Institute for Strategic Studies by Gower and Allanheld, Osmun, Farnborough, Hants and Montclair, New Jersey, 1980, 220 pp. £9.50. Marine Policy, 5(2):162–163, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900130.

Nailor:1982:BRU

[Nai82] Peter Nailor. Book review: The unnoticed challenge: Soviet maritime strategy and the global choke points, by Robert J. Hanks, Institute for Foreign Policy Analysis, Cambridge, MA, and Washington, DC, 1980, 66 pp, $6.50. Marine Policy, 6(1):85–86, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900616.

Nakahara:1985:OPR

[Nak85] Hiroyuki Nakahara. Ocean policy and R&D in Japan. Marine Policy, 9(4):333–336, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900429.

Nanlai:1989:CAI

[Nan89] Liu Nanlai. Chinese Association for International Marine Affairs Research (CAIMAR). Marine Policy, 13(3):272, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900663.

Neal:1984:RII

[NM84] R. A. Neal and J. L. Maclean. The role and impact of ICLARM. Marine Policy, 8(3):278–282, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900095.

Odeke:1984:IUC

[Ode84] Ademuni Odeke. Implementing the UNCTAD Code of Conduct: the UK Merchant Shipping (Liner Conferences) Act, 1982. Marine Policy, 8(1):56–63, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900575.

Ogley:1981:LSD

[Ogl81] Roderick C. Ogley. The law of the sea draft convention and the new international economic order. Marine Policy, 5(3):240–251, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900518.

Ogley:1984:BRT

[Ogl84] Roderick Ogley. Book review: Towards a new international marine order: edited by Finn Laursen, Martinus Nijhoff, The Hague, 1982, 198 pp. Marine Policy, 8(3):290–291, July 1984. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900174.

Olsen:1982:CFD

[Ols82] A. M. Olsen. Criticism from former director of fisheries. Marine Policy, 6(2):140–147, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900689.

Omar:1983:MTE

[Oma83] Ishak Haji Omar. Malaysian trawlers: economics of vessel size. Marine Policy, 7(3):220–222, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8390009X.

Ostreng:1986:DAA

[Øst86] Willy Østreng. Delimitation arrangements in Arctic seas: Cases of precedence or securing of strategic/economic interests? Marine Policy, 10(2):132–154, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900187.

Panayotou:1980:ECP

[Pan80] Theodore Panayotou. Economic conditions and prospects of smallscale fishermen in Thailand. Marine Policy, 4(2):142–146, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8090041X.

Pappus:1984:BRU

[Pap84] Florence A. Pappus. Book review: The UNCTAD liner code: United States maritime policy at the crossroads: Lawrence Juda, Bowker, Epping, Essex, UK, 1983, 234 pp, £17.50. Marine Policy, 8(4):366–368, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900344.

Parkinson:1988:RIM

[Par88] F. Parkinson. Reprisals in international maritime law: a historical paradigm. Marine Policy, 12(3):276–285, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8890067X.

Pauly:1982:ICW

[Pau82] Daniel Pauly. ICLARM-CSIRO Workshop on the Theory and Management of Tropical Multispecies Fish Stocks: Cronulla, Australia, 12–23 January 1981. Marine Policy, 6(1):72–74, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900501.

Pearse:1981:FRR

[Pea81a] Peter H. Pearse. Fishing rights, regulations and revenues. Marine Policy, 5(2):135–146, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900038.

Pearson:1981:BRI

[Pea81b] Roy Pearson. Book review: International ocean shipping: by Bernhard J. Abrahamson, Westview Press, Boulder, CO, 1980, 232 pp, $27.50. Marine Policy, 5(4):346, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8190169X.

Peet:1984:NSC

[Pee84] Gerard Peet. The 1984 North Sea conference: a preview. Marine Policy, 8(3):259–270, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900058.

Penick:1984:BRA

[Pen84] F. V. W. Penick. Book review: Aquaculture the legal framework: Bruce H. Wildsmith, Emond-Montgomery, Toronto, 1982, 313 pp. Marine Policy, 8(1):72–73, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900630.

Peter:1983:TMP

[Pet83] C. M. Peter. Tanzanian marine policy. Marine Policy, 7(1): 58–59, January 1983. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X83900635.

Phillips:1987:TEM

[Phi87] Cassandra Phillips. Thirty-eighth meeting of the International Whaling Commission: Malmö, Sweden, 9–13 June 1986. Marine Policy, 11(1):76–78, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900431.

Peter:1985:LUT

[PK85] C. M. Peter and P. J. Kabudi. LDCs urged to take more active part in shipping. Marine Policy, 9(2):166–167, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900144.

Plant:1985:ITS

[Pla85a] G. Plant. International traffic separation schemes in the new law of the sea. Marine Policy, 9(2):134–147, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900053.

Plant:1985:UTS

[Pla85b] G. Plant. Update: Traffic separation schemes in the EEZ. Marine Policy, 9(4):332–333, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900417.

Plant:1986:BRL

[Pla86] Glen Plant. Book review: The law of the sea and international shipping: Anglo–Soviet post-unclos perspectives: edited by W. E. Butler, Oceana, Dobbs Ferry, New York, 1985, 411 pp. $45.00. Marine Policy, 10(3):236–238, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900606.

Pauly:1985:TAL

[PM85] Daniel Pauly and Gary R. Morgan. Theory and application of length-based stock assessments: Mazara del Vallo, Sicily, 11–16 February 1985. Marine Policy, 9(4):339–340, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900466.

Pullin:1984:TAN

[PN84] Roger S. Pullin and Richard A. Neal. Tropical aquaculture: Need for a strong research base. Marine Policy, 8(3):217–228, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900022.

Pontecorvo:1986:BRP

[Pon86] Giulio Pontecorvo. Book review: Piracy today: robbery and violence at sea since 1980: by Roger Villar, Conway Maritime Press Ltd, London, Uk, 1985, £12.50. Marine Policy, 10(3):232–233, July 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900576.

Pontecorvo:1988:EMC

[Pon88] Giulio Pontecorvo. The enclosure of the marine commons: Adjustment and redistribution in world fisheries. Marine Policy, 12(4):361–372, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900206.

Pownall:1980:NSE

[Pow80] Peter Pownall. National Seminar on Economic Aspects of Limited Entry and Associated Fisheries Management Measures, sponsored by Fisheries Division, Australian Department of Primary Industry, University of Melbourne, Australian Fishing Industry Council and Fisheries and Wildlife Division, Victorian Ministry for Conservation,: University of Melbourne, Australia, 6–8 February 1980. Marine Policy, 4(3):249–251, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900639.

Pritchard:1980:BRP

[Pri80] S. Z. Pritchard. Book review: The prevention of oil pollution: edited by J. Wardley-Smith, Graham & Trotman, London, 1979, 309pp, £15.00. Marine Policy, 4(1):83–84, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900354.

Peterson:1984:MS

[PT84] Susan Peterson and Bruce Tripp. Mussels as sentinels. Marine Policy, 8(3):291–292, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900186.

Peterson:1986:SBR

[PT86] Susan Peterson and John Teal. Scientific basis for the role of the oceans as a waste disposal option: North Atlantic Treaty Organization Advanced Research Workshop, Vilamoura, April 1985. Marine Policy, 10(1):68–69, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900370.

Pullin:1980:SCA

[Pul80a] R. S. V. Pullin. Symposium on Coastal Aquaculture, held by the Marine Biological Association of India (MBAI): Cochin, Kerala State, India, 12–18 January 1980. Marine Policy, 4(3):251–253, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900640.

Pullin:1980:TFS

[Pul80b] Roger Pullin. Thailand’s first seminar on marine science. Marine Policy, 4(1):77–78, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900317.

Pullin:1982:ISG

[Pul82] R. S. V. Pullin. International symposium on genetics in aquaculture, Galway, Ireland, 29 March–2 April 1982. Marine Policy, 6(4):345–347, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8290015X.

Pullin:1984:IST

[Pul84] R. S. V. Pullin. International symposium on Tilapia in aquaculture: Nazareth, 8–13 May 1983. Marine Policy, 8(1):78–80, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900678.

Queirolo:1989:DWF

[QJ89] Lewis E. Queirolo and Richard S. Johnston. Distant water fishing nations and extended fisheries jurisdiction. Marine Policy, 13 (1):16–21, January 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900389.

Ramsay:1980:WTV

[Ram80] R. A. Ramsay. World trade versus the supply of shipping and ships. Marine Policy, 4(1):63–66, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900263.

Ramakrishna:1988:IIE

[Ram88] Kilaparti Ramakrishna. The international implications of extended maritime jurisdiction in the Pacific, twenty-first annual conference of the Law of the Sea Institute, University of Hawaii, cosponsored by the East-West Center and the Hawaii Maritime Center: Honolulu, HI, USA, 3–6 August 1987. Marine Policy, 12(2):168–170, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900486.

Ranft:1981:BRS

[Ran81] Bryan Ranft. Book review: Securing the seas: the Soviet Naval Challenge and Western Alliance Options: Paul H. Nitze, Leonard Sullivan, Jr. and the Atlantic Council Working Group on Securing the Seas, Westview Press, Boulder, CO, USA, 1979, 464 pp, £16.25. Marine Policy, 5(1):87–88, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900841.

Ranft:1982:BRP

[Ran82] Bryan Ranft. Book review: Problems of sea power as we approach the twenty first century: edited by James L. George, merican Enterprise Institute for Public Policy Research, Washington, DC, 1978, $11.75 cloth, $5.75 paperback. Marine Policy, 6(1):85, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900604.

Ramakrishna:1987:OLC

[RBA87] Kilaparti Ramakrishna, Robert E. Bowen, and Jack H. Archer. Outer limits of continental shelf: a legal analysis of Chilean and Ecuadorian island claims and US response. Marine Policy, 11(1): 58–68, January 1987. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8790039X.

Redgwell:1989:CDU

[Red89] Catherine Redgwell. Current developments: UK practice in the International Law of the Sea. Marine Policy, 13(3):266, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900596.

Rees:1989:GSB

[Ree89a] A. I. Rees. GLORIA — a successful business? Marine Policy, 13(4):309–317, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8990016X.

Rees:1989:OCO

[Ree89b] A. I. Rees. Opportunities for collaboration in ocean technology — an Anglo-Canadian conference, OPCOT 89: 6–7 February 1989, Imperial College, London. Marine Policy, 13(3):268–269, July 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900614.

Reisman:1981:ISP

[Rei81] W. Michael Reisman. Inadequacies of the straits’ passage regime in the LOS draft. Marine Policy, 5(3):276–277, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900609.

Richards:1981:SPT

[Ric81a] Robert Richards. Spanish port tendencies. Marine Policy, 5(4): 336–338, October 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81901652.

Richardson:1981:DMI

[Ric81b] Elliot L. Richardson. Decision making in the International Seabed Authority. Marine Policy, 5(3):256–260, July 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900531.

Robinson:1986:SRA

[RL86] M. A. Robinson and Rowena Lawson. Some reflections on aid to fisheries in West Africa. Marine Policy, 10(2):101–110, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900151.

Robinson:1980:WFS

[Rob80] M. A. Robinson. World fisheries to 2000: Supply, demand and management. Marine Policy, 4(1):19–32, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900214.

Roels:1981:BRM

[Roe81] Oswald A. Roels. Book review: Manganese nodules: dimensions and perspectives: Prepared by the United Nations Ocean Economics and Technology Office, D. Reidel, Dordrecht, the Netherlands, Boston, USA, London, UK, 1979. 250 pp. $34.00, hardback, $14.95, paperback. Marine Policy, 5(1):91–92, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900889.

Rose:1986:BRE

[Ros86] F. D. Rose. Book review: Enforcement of maritime claims: D. C. Jackson, Lloyd’s of London Press Ltd, London, 1985, Ixi and 376 pp, plus 93 pp appendices and 14 pp index, £39.50. Marine Policy, 10(4):316–317, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900084.

Roberts:1980:OPI

[RS80] Peter W. Roberts and Timothy Shaw. Onshore planning implications of the offshore development of mineral resources. Marine Policy, 4(2):128–141, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900408.

Sadler:1981:BRO

[Sad81] P. G. Sadler. Book review: Oil pollution from tanker operationscauses, costs, controls: by W. G. Walters, T. D. Heaver, T. Verrier, Centre for Transport Studies, University of British Columbia, Vancouver, 1980, 216 pp, $22.00. Marine Policy, 5(4):347–348, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901718.

Saliba:1983:MPH

[Sal83] Louis J. Saliba. Mediterranean pollution: Health related aspects. Marine Policy, 7(2):109–117, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900222.

Saliba:1985:BRL

[Sal85] Louis J. Saliba. Book review: The legal regime of the protection of the Mediterranean Sea against pollution from land-based sources: by Sachiko Kuwabara, National Resources and the Environment Series No 15, Tycooly International Publishing Limited, Dublin, Ireland, 1984, xix + 227 pp. Marine Policy, 9(2):164–165, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900132.

Scholz:1980:ORI

[Sch80] Wesley S. Scholz. Oceanic research: International law and national legislation. Marine Policy, 4(2):91–127, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900391.

Scovazzi:1981:INL

[Sco81] Tullio Scovazzi. Implications of the new law of the sea for the Mediterranean. Marine Policy, 5(4):302–312, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901585.

Scovazzi:1982:NLS

[Sco82] Tullio Scovazzi. The new law of the sea and Italy: sponsored by the finalized project on oceanography of the Italian Council of Scientific Research Naples 12–14 November 1981. Marine Policy, 6(3):243–244, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8290029X.

Scovazzi:1985:EEF

[Sco85] Tullio Scovazzi. Explaining exclusive fishery jurisdiction. Marine Policy, 9(2):120–125, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8590003X.

Scovazzi:1989:BRF

[Sco89] Tullio Scovazzi. Book review: The Fisheries Regime of the Exclusive Economic Zone: M. Dahmani, publications on ocean development, Vol 11, Martinus Nijhoff, Dordrecht, The Netherlands, 1987, 188 pp. Marine Policy, 13(4):358–359, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900250.

Scrivener:1983:MMS

[Scr83] David Scrivener. Merchant marine in Soviet naval strategy. Marine Policy, 7(2):118–122, April 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900234.

Sebek:1980:PLW

[Seb80] Viktor Sebek. Pollution law — the way forward. Marine Policy, 4(3):158, July 1980. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X80900500.

Sebek:1988:AIC

[Seb88] Viktor Sebek. ACOPS international conference on pollution of the marine environment: Venice, Italy, 27–30 October 1987. Marine Policy, 12(2):165–166, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900462.

Symmons:1983:MSR

[SG83] Clive R. Symmons and Piers R. R. Gardiner. Marine scientific research in offshore areas: Ireland and the law of the sea convention. Marine Policy, 7(4):291–301, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900428.

Spagni:1985:FMT

[SGG85] Daniel Spagni, Luke Georghiou, and Michael Gibbons. French marine technology policy: Industrializing ocean depths. Marine Policy, 9(4):280–291, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900375.

Shackleton:1980:LL

[Sha80a] Michael Shackleton. At loggerheads over langoustines. Marine Policy, 4(1):2, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900196.

Shackleton:1980:BRM

[Sha80b] Michael Shackleton. Book review: La mer confisquée: G. Chouraqui, Editions du Seuil, Paris, 1979. (French) [The confiscated sea]. Marine Policy, 4(4):333–334, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900123.

Shackleton:1981:WMF

[Sha81] Michael Shackleton. What made the French fishermen resort to open protest? Marine Policy, 5(4):340–343, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81901676.

Shackleton:1982:BRNa

[Sha82a] Michael Shackleton. Book review: The North Sea: a new international regime?: Records of Greenwich Forum V, 2–4 May 1979 edited by D. C. Watt, Westbury House, Guildford, Surrey, UK, 1980, 260 pp, £20, UK only. Marine Policy, 6(1):82–83, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900562.

Shackleton:1982:BRNb

[Sha82b] Michael Shackleton. Book review: The North Sea and its environment, uses and conflicts: edited by Henk van Hoorn, Gerard Peet and Marianne Tiel, Netherlands North Sea Working Group, Amsterdam, 1980. Marine Policy, 6(1):82–83, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900574.

Shah:1984:WST

[Sha84] M. J. Shah. World shipping: a time for redefinition? Marine Policy, 8(4):347–350, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900277.

Shinkaretskaia:1988:ICJ

[Shi88] G. Shinkaretskaia. The International Court of Justice and the development of the Law of the Sea. Marine Policy, 12(3):201–210, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X88900589.

Shusterich:1982:MDS

[Shu82] Kurt Shusterich. Mining the deep seabed: a complex and innovative industry. Marine Policy, 6(3):175–192, July 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900227.

Shusterich:1984:MRD

[Shu84] Kurt M. Shusterich. Marine resource development in Belize. Marine Policy, 8(4):369–372, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900368.

Sidey:1980:UPB

[Sid80] J. MacN. Sidey. UK ports — the battle for traffic. Marine Policy, 4(2):90, April 1980. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8090038X.

Silverstone:1985:SSO

[Sil85] Deana Silverstone. Ship source oil pollution damage: a Canadian perspective on recoverability of economic losses and damage to the marine environment. Marine Policy, 9(2):108–119, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900028.

Simnet:1983:IGA

[Sim83] Jon Simnet. ‘Industry, government and academia: partners in progress’ Oceans 82: Washington, DC, 20–22 September 1982. Marine Policy, 7(1):68, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900696.

Sinclair:1986:SSC

[Sin86] Peter R. Sinclair. The survival of small capital: State policy and the dragger fleet in north-west Newfoundland. Marine Policy, 10(2):111–118, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900163.

Sintsova:1988:RNS

[Sin88] T. A. Sintsova. The regime of navigation in the Suez Canal. Marine Policy, 12(3):263–270, July 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900656.

Sissenwine:1982:FMT

[SK82] M. P. Sissenwine and J. E. Kirkley. Fishery management techniques: Practical aspects and limitations. Marine Policy, 6(1):43–57, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900410.

Sebek:1988:DMC

[SK88] Viktor Sebek and Rouchdy Kbaier. Dakar ministerial conference on environment and development: Dakar, Senegal, 14–15 January 1988. Marine Policy, 12(2):166–168, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900474.

Scully:1989:ATL

[SK89] R. Tucker Scully and Lee A. Kimball. Antarctica: is there life after minerals?: the minerals treaty and beyond. Marine Policy, 13(2):87–98, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900018.

Slater:1984:CFL

[Sla84] Jim Slater. Cheap foreign labour: empire relic with no place in civilized society. Marine Policy, 8(1):64–65, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900587.

Smith:1981:SUP

[SLB81] H. D. Smith, C. S. Lalwani, and I. C. Brookfield. Sea-use planning and the North Sea. Marine Policy, 5(1):75–76, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900786.

Smalley:1984:RMF

[Sma84] G. Smalley. Role of merchant fleet In Soviet global strategy. Marine Policy, 8(1):65–68, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900599.

Smith:1981:IFI

[Smi81] Ian R. Smith. Improving fishing incomes when resources are overfished. Marine Policy, 5(1):17–22, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900701.

Smith:1982:BRA

[Smi82] Ian R. Smith. Book review: Aquaculture economics basic concepts and methods of analysis: by Y. C. Shang, Westview Press, Boulder, CO, and Croom Helm, London, 1981, 124 pp, $17.50. Marine Policy, 6(2):158–159, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900768.

Smith:1983:BRA

[Smi83] Ian R. Smith. Book review: Aquaculture development in less developed countries: Social, economic and political problems: edited by L. J. Smith and S. Peterson, Westview, Boulder, CO, 1982, 152 pp. Marine Policy, 7(4):321–322, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900507.

Spagni:1987:BRS

[Spa87] Daniel A. Spagni. Book review: South East Asian seas: Oil under troubled waters: by Mark J. Valencia, Oxford University Press, Singapore, 155 pp, £17.00. Marine Policy, 11(2):154–155, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900133.

Speed:1982:RNC

[Spe82] Keith Speed. Royal Navy cuts — former Minister comments. Marine Policy, 6(2):150–151, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900719.

Speed:1983:BRB

[Spe83] Keith Speed. Book review: Britain’s naval future: Sir James Cable, Macmillan, 1983, 220 pp, £20. Marine Policy, 7(4):318–319, October 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900477.

Sathiendrakumar:1986:FRP

[ST86] R. Sathiendrakumar and C. Tisdell. Fishery resources and policies in the Maldives: Trends and issues for an island developing country. Marine Policy, 10(4):279–293, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900047.

Steel:1984:FPE

[Ste84] David I. A. Steel. Fisheries policy and the EEC: the democratic influence. Marine Policy, 8(4):350–353, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900289.

Stokes:1980:PFF

[Sto80a] Robert L. Stokes. Prospects for foreign fishing vessels in US fisheries development. Marine Policy, 4(1):33–41, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900226.

Stokes:1980:UFC

[Sto80b] Robert L. Stokes. US — foreign cooperation for fisheries development — a preliminary economic analysis. Marine Policy, 4(4): 326–328, October 1980. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8090007X.

Sturmey:1984:ACC

[Stu84] S. G. Sturmey. Administering the code of conduct. Marine Policy, 8(3):203–216, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900010.

Sturmey:1985:BRS

[Stu85a] S. G. Sturmey. Book review: Shipping conferences: Amos Herman Lloyd’s of London Press Ltd, 1983, pp vii + 255, including text of the Convention on a Code of Conduct for Liner Conferences, text of Code of Practice for Conferences, prepared by CENSA and ESC and selected bibliography. Marine Policy, 9(1):84–85, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900892.

Sturmey:1985:MJV

[Stu85b] S. G. Sturmey. Maritime joint ventures: a two-day conference organized jointly by the International Chamber of Commerce and Lloyd’s of London Press Ltd in association with the International Chamber of Shipping and Lloyd’s List, May 1984. Marine Policy, 9(2):158–159, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900089.

Sturmey:1986:BRS

[Stu86] S. G. Sturmey. Book review: Shipping nationalism and the future of the United States liner industry: the UNCTAD code and bilateralism; proceedings of a workshop organised by the Center for Ocean Management Studies, University of Rhode Island, USA, November, 1983: Times Press, 1984, pp 164 + viii. Marine Policy, 10(2):161–162, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900230.

Sturmey:1988:BRW

[Stu88] S. G. Sturmey. Book review: The world shipping industry: by Ernst Gabriel Frankel, Croom Helm, New York, 1987, 261 pp. £35.00. Marine Policy, 12(2):173–174, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900504.

Sullivan:1980:CIM

[Sul80] William L. Sullivan. Constituting the IOC as a more autonomous or independent body: the long-term implications. Marine Policy, 4(4):290–308, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900032.

Sullivan:1989:CMN

[Sul89] Karl M. Sullivan. Conflict in the management of a Northwest Atlantic transboundary cod stock. Marine Policy, 13(2):118–136, April 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X89900031.

Suter:1981:BRA

[Sut81a] Keith D. Suter. Book review: Antarctica and its resources: by Barbara Mitchell and Jon Tinker, Earthscan/International Institute for Environment and Development, London, 1979, 94 pp and appendix. Marine Policy, 5(2):163–164, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900142.

Suter:1981:BRM

[Sut81b] Keith D. Suter. Book review: The management of the Southern Ocean: Barbara Mitchell and Richard Sandbrook, IIED, 10 Percy Street, London W1P ODR, UK, 1980, 162 pp. Marine Policy, 5(1):88–89, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X81900853.

Suter:1982:ANW

[Sut82a] Keith D. Suter. Australia’s new whaling policy: Formulation and implementation. Marine Policy, 6(4):287–302, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900045.

Suter:1982:SPN

[Sut82b] Keith D. Suter. The South Pacific — a new region for confrontations? Marine Policy, 6(1):58–60, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900422.

Suter:1983:MPD

[Sut83] Keith D. Suter. Marine policy and disarmament. Marine Policy, 7(3):213–215, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X83900064.

Thiessen:1986:FNM

[TD86] Victor Thiessen and Anthony Davis. A further note to making sense of the dollars: income distribution among Atlantic Canadian fishermen and public policy. Marine Policy, 10(4):310–314, October 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900060.

Teiwaki:1987:AAS

[Tei87] Roniti Teiwaki. Access agreements in the South Pacific: Kiribati and the distant water fishing nations 1979–1986. Marine Policy, 11(4):273–284, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900182.

tenHallers:1987:SNS

[tH87] Cato ten Hallers. The second North Sea seminar ’86. The status of the North Sea environment: Reasons for concern: Rotterdam, The Netherlands, 1–3 October 1986. Marine Policy, 11(2):148–150, April 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900091.

Thomas:1981:INS

[Tho81] Raju G. C. Thomas. Indian Navy — searching for a role and rationale. Marine Policy, 5(4):331–332, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81901615.

tenHallers-Tjabbes:1982:RIN

[tHTP82] Cato C. ten Hallers-Tjabbes and Gerard M. Peet. Research interests in the North Sea. Marine Policy, 6(1):63–66, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900458.

Tietze:1984:MFE

[Tie84] U. Tietze. Marine fisheries extension services: a training and visit system for southern Asia. Marine Policy, 8(4):330–336, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900253.

Till:1984:SFU

[Til84] G. Till. Strategic forum: the USA, its Navy and the Pacific. Marine Policy, 8(3):271–274, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8490006X.

Till:1985:BRM

[Til85] Geoffrey Till. Book review: Maritime strategy or coalition defense?: Robert W. Komer, Abt Books, Cambridge, MA, 1984, xviii + 109 pp. Marine Policy, 9(2):159–160, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900090.

Tambs-Lyche:1984:AUF

[TL84] Hans Tambs-Lyche. After UNCLOS: the future of international marine research. Marine Policy, 8(4):353–356, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X84900290.

Tomczak:1980:IDW

[Tom80] Matthias Tomczak. The IWC and the decimation of whale stocks. Marine Policy, 4(1):79–82, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900330.

Tomczak:1982:NRI

[Tom82] Matthias Tomczak. New role for IOC. Marine Policy, 6(1):60, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900434.

Tomczak:1984:DMP

[Tom84] M. Tomczak. Defining marine pollution: a comparison of definitions used by international conventions. Marine Policy, 8(4):311–322, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8490023X.

Treves:1981:DC

[Tre81] Tullio Treves. Drafting the LOS Convention. Marine Policy, 5 (3):273–276, July 1981. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X81900592.

Treves:1983:ALS

[Tre83] Tullio Treves. The adoption of the law of the sea convention: Prospects for seabed mining. Marine Policy, 7(1):3–13, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900568.

Till:1983:MMS

[TS83] G. Till and G. Smalley. Merchant marine in Soviet naval strategy: letters. Marine Policy, 7(3):212–213, July 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X83900052.

Tsamenyi:1986:SPS

[Tsa86] B. Martin Tsamenyi. The South Pacific states, the USA and sovereignty over highly migratory species. Marine Policy, 10(1): 29–41, January 1986. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X86900308.

Tsarev:1988:PUS

[Tsa88] V. F. Tsarev. Peaceful uses of the seas: Principles and complexities. Marine Policy, 12(2):153–159, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900425.

Turgeon:1985:FRU

[Tur85] Donna D. Turgeon. Fishery regulation: Its use under the Magnuson Act and reaganomics. Marine Policy, 9(2):126–133, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900041.

Tiler:1988:MST

[TY88] Christine Tiler and John Yates. Marine science and technology in the Federal Republic of Germany: a policy for non-living resource utilization. Marine Policy, 12(4):373–381, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900218.

Underdal:1980:IMP

[Und80] Arild Underdal. Integrated marine policy: What? Why? How? Marine Policy, 4(3):159–169, July 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900512.

Valencia:1988:YST

[Val88] Mark J. Valencia. The Yellow Sea: Transnational marine resource management issues. Marine Policy, 12(4):382–395, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8890022X.

Valencia:1989:DCS

[Val89] Mark J. Valencia. Discussions with Continental Shelf Committee of the Council of Ministers of Vietnam. Marine Policy, 13(4): 364–365, October 1989. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X89900328.

Vandermeulen:1985:BRR

[Van85] J. H. Vandermeulen. Book review: Restoration of habitats impacted by oil spills: edited by John Cairns and Arthur Buikema, 1984, 182 pp. Marine Policy, 9(4):349–350, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900533.

VanDyke:1987:BRU

[Van87] Jon M. Van Dyke. Book review: United nations convention on the law of the sea 1982: a commentary volume 1: edited by Myron H. Nordquist, Center for Ocean Law and Policy, University of Virginia, Martinus Nijhoff Publishers, Dordrecht, 1985, Ixv + 467 pp, $100. Marine Policy, 11(3):241–242, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900613.

VanDyke:1988:ODN

[Van88] Jon M. Van Dyke. Ocean disposal of nuclear wastes. Marine Policy, 12(2):82–95, April 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8890036X.

vanHegelsom:1981:AIA

[vH81] G. J. F. van Hegelsom. The argument for an interim arrangement. Marine Policy, 5(3):260–264, July 1981. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900543.

vanHoorn:1985:HNS

[vHPW85] Henk van Hoorn, Gerard Peet, and Koos Wieriks. Harmonizing North Sea policy in the Netherlands. Marine Policy, 9(1):53–61, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8590079X.

Volckaert:1986:FNM

[Vol86] Filip Volckaert. Formulation of national marine policy: the case of Belgium. Marine Policy, 10(2):90–100, April 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X8690014X.

Vukas:1985:BRU

[Vuk85] Budislav Vukas. Book review: The USSR, Eastern Europe and the development of the law of the sea: W. E. Butler, Oceana Publications, Inc. London, Rome, New York, 1983, Looseleaf, $100. Marine Policy, 9(1):85–86, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900909.

Wahiche:1983:AST

[Wah83] Jean-Dominique Wahiche. Artificial structures and traditional uses of the sea: the field of conflict. Marine Policy, 7(1):37–52, January 1983. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X83900593.

Walkerden:1982:AA

[Wal82] R. D. Walkerden. Administrative aspects. Marine Policy, 6(2): 148–149, April 1982. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X82900707.

Walder:1986:BRU

[Wal86] C. A. Walder. Book review: United Kingdom marine pollution law: by John H. Bates, Lloyds of London Press, London, 1985, 461 + xxxiipp, £35.00. Marine Policy, 10(1):81–82, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8690045X.

Walder:1989:BRS

[Wal89] C. A. Walder. Book review: State responsibility and the marine environment — The rules of decision: by Brian D. Smith, Clarendon Press, Oxford, 1988, 281 pp. £30.00. Marine Policy, 13(4):352–353, October 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X89900201.

Warner:1985:BRI

[War85] David W. Warner. Book review: The International Transport workers’ Federation and Flag of Convenience Shipping: Herbert R. Northrup and Richard I. Rowan, The Wharton School, University of Pennsylvania, PA, 1983, 251 pp. Marine Policy, 9(1):80–81, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900867.

Watt:1980:BRE

[Wat80a] D. C. Watt. Book review: The effective management of resources: the International Politics of the North Sea, edited by C. M. Mason, Francis Pinter, London, Nichols, New York, 1979, 268pp, £10.00. Marine Policy, 4(2):154–155, April 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900470.

Watt:1980:IPO

[Wat80b] D. C. Watt. Integrated policy for the oceans — teaching in the university world. Marine Policy, 4(1):67–68, January 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X80900275.

Watt:1981:BRA

[Wat81a] D. C. Watt. Book review: Arms control: the interwar naval limitation agreements: by Robert A. Hoover, Vol 17, Books, Monograph Series in World Affairs, University of Denver (Colorado Seminary), 1980, 124pp, $3.90. Marine Policy, 5(4):349, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901743.

Watt:1981:BRS

[Wat81b] D. C. Watt. Book review: Sea power and politics from the Norman conquest to the present day: by Captain John Moore, Weidenfeld and Nicolson, London, 1979, 184 pp, £7.50. Marine Policy, 5(4):349, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901731.

Watt:1981:BRMb

[Wat81c] D. C. Watt. Book review: The maritime dimension: edited by R. P. Barston and Patricia Birnie, George Allen and Unwin, London, 1980, 194 pp, hard cover £11.95, paperback £5.50. Marine Policy, 5(4):348–349, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X8190172X.

Watt:1981:F

[Wat81d] D. C. Watt. Film. Marine Policy, 5(2):167–168, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900178.

Watt:1981:BRP

[Wat81e] Donald C. Watt. Book review: Pollution, politics and international law: Tankers at sea, by R. Michael M’Gonigle and Mark W. Zacher, University of California Press, Berkeley and Los Angeles, 1979, 394pp, £9.50. Marine Policy, 5(2):162, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900129.

Watt:1981:BRMa

[Wat81f] U. C. Watt. Book review: The Middle Eastern States and the Law of the sea: Ali A. El-Hakim, Manchester University Press. Manchester, UK. 1979, 233 pp. £19.50. Marine Policy, 5(1):84–85, January 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81900816.

Watt:1982:BRI

[Wat82a] Donald C. Watt. Book review: Internationai politics and the sea: the case of Brazil, by Michael A. Morris, Westview Press, Boulder, CO, 1979, 291 pp, $22. Marine Policy, 6(1):83–84, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900586.

Watt:1982:BRO

[Wat82b] Donald C. Watt. Book review: Ocean Yearbook 2: edited by Elisabeth Mann Borgese and Norton Ginsburg, University of Chicago Press for International Oceanic Institute, Chicago and London, 1980, 713pp. Marine Policy, 6(1):78–79, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900537.

Watt:1982:BRP

[Wat82c] Donald C. Watt. Book review: Pollution in the Mediterranean: Directed by Jean Pierre Baux La service de film de Recherche Scientifique CNDP, France, 1980, 30 mins. This film may be obtained from the Scottish Central Film Library, Dowanhill, 74 Victoria Crescent Road, Glasgow G12 9JN, UK. Marine Policy, 6(4):341–342, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900136.

Watt:1982:F

[Wat82d] Donald C. Watt. Film. Marine Policy, 6(1):68–69, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900471.

Watt:1984:TTL

[Wat84a] Donald Cameron Watt. A tale of two legal systems: 17th Annual Conference of the Law of the Sea Institute cosponsored by the Fridtjof Nansen Institute, Oslo, Norway, 13–16 July 1983. Marine Policy, 8(1):74–75, January 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900654.

Watt:1984:BRC

[Wat84b] Donald Cameron Watt. Book review: A crime against the world: Memoirs of a Russian Sea Captain Vladil Lysenko, translated by Michael Glenny, Victor Gollancz, London, 1983. 254 pp. Marine Policy, 8(3):284–285, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X84900113.

Watt:1984:BRE

[Wat84c] Donald Cameron Watt. Book review: Eighth voyage of the dragon: a history of China’s quest for seapower: Bruce Swanson, US Naval Institute Press, Annapolis, MD, 1982, 348 pp. Marine Policy, 8 (1):69–71, January 1984. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X84900617.

Watt:1985:BRO

[Wat85a] Donald Cameron Watt. Book review: Ocean Yearbook 4: edited by Elisabeth Mann Borgese and Norton Ginsburg, Chicago University Press, $49. Marine Policy, 9(2):163–164, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X85900120.

Watt:1985:BRF

[Wat85b] Donald Cameron Watt. Book review: The fish feud: The US and Canadian Boundary Dispute: David VanderZwaag, Lexington, MA (UK: Gower, Aldershot, Hants), 1984, 112 pp, £19.50. Marine Policy, 9(1):86–87, January 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900910.

Watt:1985:BRI

[Wat85c] Donald Cameron Watt. Book review: The International Maritime Organization: edited by Samir Mankabady, Croom Helm, London and Sydney, £35. Marine Policy, 9(2):157–158, April 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X85900077.

Watt:1986:BRF

[Wat86a] D. Cameron Watt. Book review: The future of British sea power: edited by Geoffrey Till, Macmillan’s Press, London, 1984, 278 pp. Marine Policy, 10(1):79–80, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X86900436.

Watt:1986:BRO

[Wat86b] Donald Cameron Watt. Book review: Ocean yearbook 5: edited by Elisabeth Mann Borgese and Norton Ginsburg, University of Chicago Press, Chicago and London, 1985, 544 pp, $49.00. Marine Policy, 10(1):82–84, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X86900461.

Watt:1987:BRA

[Wat87a] Donald Cameron Watt. Book review: Advances in underwater technology, ocean science and underwater engineering, volume 8 exclusive economic zones, resource opportunities and the legal regime: Society for underwater technology Graham and Trotman, London, 1986, 187 pp, £30.00. Marine Policy, 11(4):324–325, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900273.

Watt:1987:BRD

[Wat87b] Donald Cameron Watt. Book review: Diplomacy at sea: James Cable, Macmillan, London, 1985, 191 pp, £25.00. Marine Policy, 11(1):82–83, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900480.

Watt:1987:BRL

[Wat87c] Donald Cameron Watt. Book review: Law, force and diplomacy at sea: Ken Booth, George Allen and Unwin, London, 1985, 231 pp, £18.50 hb, £7.95 pb. Marine Policy, 11(1):82–83, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900479.

Watt:1987:BRMb

[Wat87d] Donald Cameron Watt. Book review: Maritime affairs — a world handbook: edited by Henry W. Dagenhardt, Longman’s Group, London, 1986, 412 pp, £55.00. Marine Policy, 11(4):320–321, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900248.

Watt:1987:BRMa

[Wat87e] Donald Cameron Watt. Book review: Maritime strategy for medium powers: Rear Admiral J. R. Hill RN (Ret), Croome Helm, London, 1986, 247 pp, £19.95. Marine Policy, 11(1):82–83, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900492.

Watt:1987:BROb

[Wat87f] Donald Cameron Watt. Book review: Oil pollution control: by Sonia Zaide Pritchard, Croom Helm, London, 1987, x + 231 pp, £25.00. Marine Policy, 11(3):245–246, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900649.

Watt:1987:BROa

[Wat87g] Donald Cameron Watt. Book review: Ordering the oceans the making of the law of the sea: by Clyde Sanger, Zed Books Ltd, London, 1986, 272 pp, £18.95 hardback, £6.95 paperback. Marine Policy, 11(3):242–244, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900625.

Watt:1987:BRN

[Wat87h] Donald Cameron Watt. Book review: The new order of the ocean: the advent of a managed environment: edited by Giulio Pontecorvo, Columbia University Press, New York, 1986, xv + 277 pp, US$30. Marine Policy, 11(3):244–245, July 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X87900637.

Watt:1987:BRW

[Wat87i] Donald Cameron Watt. Book review: The Weidenfeld Atlas of Maritime History: Richard Natkiel and Antony Preston, Weidenfeld and Nicolson, London, 256pp, £16.95. Marine Policy, 11(1):81–82, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X87900467.

Wells:1988:PME

[WC88] P. G. Wells and R. P. Côté. Protecting marine environmental quality from land-based pollutants: the strategic role of ecotoxicology. Marine Policy, 12(1):9–21, January 1988. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X88900036.

Wecker:1987:TAC

[Wec87] Miranda Wecker. Twentieth Annual Conference of the Law of the Sea Institute: Miami, FL, USA, 21–24 July 1986. Marine Policy, 11(1):72–74, January 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900418.

Wecker:1989:TSA

[Wec89] Miranda Wecker. Twenty-second annual conference of the Law of the Sea Institute: University of Rhode Island RI, USA, 12–16 June 1988. Marine Policy, 13(1):74–77, January 1989. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X89900456.

Weiss:1984:BRL

[Wei84] Freidl Weiss. Book review: The law of the sea: R. R. Churchill and A. V. Lowe, Manchester University Press, Manchester, 1983, xxxiii and 321 pp. Marine Policy, 8(3):282–284, July 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900101.

Westermeyer:1982:RAA

[Wes82] William Westermeyer. Resource allocation in Antarctica: a review. Marine Policy, 6(4):303–325, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X82900057.

Warren:1982:RFS

[WGG82] John P. Warren, Wade L. Griffin, and William E. Grant. Regional fish stock management: a model for North-West Africa. Marine Policy, 6(2):121–130, April 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900665.

Wijkman:1980:ECR

[Wij80] Per Magnus Wijkman. Effects of cargo reservation: a review of UNCTAD’s code of conduct for liner conferences. Marine Policy, 4(4):271–289, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X80900020.

Wildsmith:1982:BRC

[Wil82] Bruce H. Wildsmith. Book review: Coastal aquaculture law and policy: a case study of California: by Gerald Bowen, Westview Press, Boulder, CO, 1981, 241 pp. Marine Policy, 6(4):339–340, October 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X82900124.

Whitmarsh:1985:MUM

[WY85] David J. Whitmarsh and James A. Young. Management of the UK mackerel fisheries: an economic perspective. Marine Policy, 9(3):220–236, July 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X85900211.

Yahaya:1981:CFP

[Yah81] Jahara Yahaya. Capture fisheries in Peninsular Malaysia: Lessons from MAJUIKAN’s experience. Marine Policy, 5(4):322–330, October 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X81901603.

Yates:1988:SUT

[Yat88] John Yates. The Society for Underwater Technology/The Challenger Society: UWIST, Cardiff, 12–13 April 1988. Marine Policy, 12(4):411–412, October 1988. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X88900280.

Young:1980:UMP

[You80] Wayland Young. UK maritime policy — coordination at government level. Marine Policy, 4(4):322–325, October 1980. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X80900068.

Young:1981:ANI

[You81] Elizabeth Young. The Arctic — in need of an international regime. Marine Policy, 5(2):154–156, April 1981. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X81900075.

Young:1982:UFM

[You82] James A. Young. UK fish marketing — a fresh approach. Marine Policy, 6(1):61–62, January 1982. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X82900446.

Young:1984:PCS

[You84] P. Lewis Young. Potential conflict in the South China Sea and environs. Marine Policy, 8(4):360–363, October 1984. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0308597X84900320.

Young:1985:BRN

[You85] Elizabeth Young. Book review: The North Sea: Sea use management and planning: H. D. Smith and C. S. Lalwani, North Sea Research Unit, Centre for Marine Law and Policy, University of Wales Institute of Science and Technology, Cardiff, UK, 1984, 363 pp, bibliography, no index. Marine Policy, 9(4):346–348, October 1985. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X8590051X.

Young:1986:HEM

[You86] Elizabeth Young. The hybrid emergency: multiple risk and hazard in marine transportation: the Greenwich forum, Royal United Services Institution, London, 7 June 1985. Marine Policy, 10(1):66–68, January 1986. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/0308597X86900369.

Young:1987:BRS

[You87a] Elizabeth Young. Book review: The Strait of Dover, volume 8, International Straits of the World: by Luc Cuyvers, Martinus Nijhoff Publishers, Dordrecht, 150 pp, £31.95. Marine Policy, 11 (2):152–153, April 1987. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0308597X8790011X.

Young:1987:CBS

[You87b] Elizabeth Young. Closing the British seas: a comment. Marine Policy, 11(4):326–327, October 1987. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0308597X87900285.